### v2

# Tensor products of bimodules and bimodule pairs over monoids

Abstract.
We modify the well-known tensor product of modules over a semiring, in order to treat modules over hyperrings, and, more generally, for bimodules (and bimagmas) over monoids. The tensor product of residue hypermonoids is functorial. Special attention is paid to different kinds of morphisms and the work of Nakamura and Reyes.
Key words and phrases:
hyperfield, hypergroup, Krasner, residue, semiring, subgroup, supertropical algebra, surpassing relation, pair, tropical.
Key words and phrases:
bimodule, hyperfield, hyperring, pair, magma, quotient hyperring, surpassing relations, tensor product, weak morphism
2020 Mathematics Subject Classification:
Primary 14T10, 16Y20, 16Y60, 20N20; Secondary 15A80; .
2020 Mathematics Subject Classification:
Primary 15A69, Secondary 15A80, 16D90, 17D99
The author was supported by the ISF grant 1994/20 and the Anshel Peffer Chair.
The author is indebted to Jaiung Jun for many enlightening conversations about the categorical aspects.
1. Introduction
This is part of an ongoing project in attempting to find a general algebraic framework which is suitable for tropical mathematics, and also to handle varied structures such as idempotent semirings, hyperfields, and so forth. In the process, we need to bypass negation as much as feasible. This idea can be found in [6, 7], and was implemented for blueprints in [15, 16], and put in a more general context in [24]. A minimalist set of axioms, for a “pair,” was introduced in [11], and made more precise in [2], as an associative structure $(\mathcal{A},+)$ with a distinguished substructure $\mathcal{A}_{0}$ taking the place of a zero element, which is acted on by a set $\mathcal{T}$ which often is a group. To avoid further repetition, we refer to the introductions of [1] and [11] for more historical background.
Pairs can be described in terms of universal algebra, and are closed under products and sums and polynomial extensions, and provide a concrete tensor category in the sense of [9, 21]. The main example in this paper is for hypermodules. The tensor construction for residue (quotient) hypermodules is functorial.
However, tensor products are not so malleable when one engages with the sorts of morphisms which arise in hyperfields, such as “weak morphisms” (Definition 2.26). [4] came up with a fully satisfactory theory for monoids, (also see [3]), but it seems that the presence of the extra operation of addition often hampers basic results such as an adjoint isomorphism theorem. One could attempt to define the tensor product of weak morphisms on simple tensors, but even this need not be well-defined without some further assumption. In this paper we do obtain weak morphisms for tensor extensions, and for tensor products when one of the tensor components is free.
Recently, Nakayama and Reyes [20] have tackled the issue of tensor products of hypermagmas (in which hyperaddition $\boxplus$ is replaced by a binary operation *), in an interesting paper which also provides intriguing examples of noncommutative hypergroups. Their solution was to declare that non-simple tensors are vacuous, which is needed to obtain a general categorical tensor product of hypermagmas. The downside is that one could lose associativity of addition, as seen in Example 5.3.
1.1. Shape of the paper
We start by reviewing the basics of bimagmas, also endowed with a two-sided action, and of pairs, possibly with surpassing relations. There are three classes of morphisms (each pertaining to its own category): homomorphisms, $\preceq$-morphisms, and weak morphisms. Our main example in this paper is “hyperpairs,” inspired by hyperfield theory (and more generally the hypermagmas of [20]), although many more examples can be found in [1] and [11].
Then we bring in the main notion of this paper, the tensor product. We take a standard semiring-theoretic approach which is appropriate for the categories involving homomorphisms, but to handle the tensor product of $\preceq$-morphisms or weak morphisms in general, we need either to restrict to free bimodules, or to turn more towards ideas in [20].
As in the familiar theory of tensor products of modules, and rings, we differentiate between bimodules (resp. bimagmas) and semirings (resp. pre-semirings). The tensor extension of a pair works surprisingly well.
2. Preliminaries
2.1. Underlying algebraic structures
$\mathbb{N}^{*}$ denotes the positive natural numbers, and we set $\mathbb{N}=\mathbb{N}^{*}\cup{\mathbb{0}}.$
Definition 2.1.
  1. (i)
    A magma is a set $\mathcal{A}$ with a binary operation $*:(\mathcal{A}\cup\{\infty\})\times(\mathcal{A}\cup\{\infty\})\to\mathcal{A}\cup\infty$, not necessarily associative, satisfying $\infty*b=b*\infty=\infty$ for all $b\in\mathcal{A}.$ (Thus $\infty$ is an absorbing element.) The magma is total if the operation is total, i.e., $*:\mathcal{A}\times\mathcal{A}\to\mathcal{A}$ [^1] In most applications the magma is total, but we shall utilize $\infty$ to describe tensor products. Given any total operation, one can formally adjoin the absorbing element $\infty$.
    In this paper a magma always has a neutral element, often denoted $\iota$, i.e., $\iota*b=b*\iota=b$ for all $b\in\mathcal{A}$222 In [20] these magmas are called unital. [20] calls this element $e,$ but here we use $e$ for another purpose, cf. §2.3.
  2. (ii)
    A monoid is a magma with an associative operation, not necessarily commutative. An additive semigroup is an abelian monoid, with the operation denoted by “$+$,” and with a zero element $\mathbb{0}.$
  3. (iii)
    A pre-semiring $\mathcal{A}$ is a monoid under two operations, multiplication, denoted as concatenation, and addition. We shall denote multiplication by concatenation. We can always adjoin an element  $\mathbb{0}$ to $\mathcal{A}$ that is additively neutral and also multiplicatively absorbing, and a unity element $\mathbb{1}$ that is multiplicatively neutral, so we only consider pre-semirings that have such a $\mathbb{0}$ and $\mathbb{1}$.
  4. (iv)
    A semiring [8] is a pre-semiring that satisfies all the properties of a ring (including associativity and distributivity of multiplication over addition), but without negation.
  5. (v)
    A semiring $\mathcal{A}$ is a semifield if $(\mathcal{A},\cdot)$ is a group.
2.1.1. $(\mathcal{T}_{1},\mathcal{T}_{2})$-magmas and bimagmas
We take some notions from [2], slightly more generally, taking [20] into account.
Definition 2.2.
Let $(\mathcal{T},\mathbb{1})$ be a monoid with a unit element $\mathbb{1}$.
  1. (i)
    A left $\mathcal{T}$-set is a set $\mathcal{A}$ together with a (left) $\mathcal{T}$-action $\mathcal{T}\times\mathcal{A}\to\mathcal{A}$ (denoted as concatenation), for which $\mathbb{1}b=b$ and $a_{1}(a_{2}b)=(a_{1}a_{2})b$ for all $b\in\mathcal{A}$ and $a_{i}\in\mathcal{T}$.
  2. (ii)
    A $(\mathcal{T}_{1},\mathcal{T}_{2})$-biset is a left $\mathcal{T}_{1}$-set and right $\mathcal{T}_{2}$-set $\mathcal{A}$, for which $(a_{1}b)a_{2}=a_{1}(ba_{2})$ for all $a_{i}\in\mathcal{T}_{i}$ and $b\in\mathcal{A}.$
  3. (iii)
    A left $\mathcal{T}$-magma is a magma $(\mathcal{A},*,\iota)$, together with a (left) $\mathcal{T}$-action $\mathcal{T}\times\mathcal{A}\to\mathcal{A}$ (also denoted as concatenation), for which
    1. (a)
      $\iota$ is absorbing, i.e. $a\iota=\iota,\ \text{for all}\;a\in\mathcal{T}.$
    2. (b)
      The action is distributive over $\mathcal{T}$, in the sense that
      $a(b_{1}*b_{2})=ab_{1}*ab_{2},\quad\text{for all}\;a\in\mathcal{T},\;b_{i}\in\mathcal{A}.$
      (In particular, $a(b_{1}*b_{2})=\infty$ if and only if $ab_{1}*ab_{2}=\infty.$)
    Then we adjoin $\iota$ to $\mathcal{T}$, and denote $\mathcal{T}_{\iota}=\mathcal{T}\cup{\iota}$, declaring $\iota b=\iota$ for each $b\in\mathcal{A}$. We call the elements of $\mathcal{T}$ tangible. $\mathcal{T}$ will normally be understood from the context; in case of ambiguity, we write $\mathcal{T}_{\mathcal{A}}$ for emphasis.
  4. (iv)
    A left $\mathcal{T}^{\prime}$-submagma of a left $\mathcal{T}$-magma $\mathcal{A}$ is a left submagma $\mathcal{A}^{\prime}$ over some set $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ containing $\mathbb{1}.$
  5. (v)
    A $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $\mathcal{A}$ is a $({\mathcal{T}_{1}}_{0},{\mathcal{T}_{2}}_{0})$-biset which is both a left $\mathcal{T}_{1}$-magma and a right $\mathcal{T}_{2}$-magma.
  6. (vi)
    A $\mathcal{T}$-bimagma is a $(\mathcal{T},\mathcal{T})$-bimagma.
  7. (vii)
    $\mathcal{T}$ is $\mathcal{A}$-central if $ab=ba$ for all $a\in\mathcal{T}$ and $b\in\mathcal{A}.$
Remark 2.3.
  1. (i)
    The classical case is when $R$ is an algebra over a commutative ring $C$, and $\mathcal{T}=C$, viewed as a multiplicative monoid.
  2. (ii)
    If $\mathcal{A}$ did not already contain a neutral element $\iota,$ we already could adjoin it formally by declaring its operation on all $b\in\mathcal{A}$ by $\iota*b=b*\iota=b$, and $a\iota=\iota$ for all $a\in\mathcal{T}.$
  3. (iii)
    More generally, $\mathcal{T}$ could be an arbitrary set, with changes mutatis mutandis, dropping Defininition 2.2(i). For example, $\mathcal{T}$ could be, say, a set of matrix units.
(So far these definitions are rather general, and could include for example the case where $\mathcal{A}$ is a monoid with unit element $\iota$, and $\mathcal{T}$ is its set of monoid endomorphisms.)
Definition 2.4.
  1. (i)
    A $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $\mathcal{A}$ is left weakly admissible if $\mathcal{T}_{1}\subseteq\mathcal{A}$.
  2. (ii)
    A left weakly admissible $(\mathcal{T}_{1},\mathcal{T}_{2})$-magma $\mathcal{A}$ is called left admissible if $\mathcal{A}$ is spanned by $\mathcal{T}_{1}$.
  3. (iii)
    A right weakly admissible, resp. right admissible $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma is defined analogously to (i) and (ii).
  4. (iv)
    A weakly admissible $\mathcal{T}$-bimagma $\mathcal{A}$ is a left and right weakly admissible $(\mathcal{T},\mathcal{T})$-bimagma (identifying both copies of $\mathcal{T}$ inside $\mathcal{A}$).
Lemma 2.5.
Let $\mathcal{A}$ be a weakly admissible $\mathcal{T}$-bimagma. Then $\langle\mathcal{T}\rangle$ (the submagma generated by $\mathcal{T}$) is an admissible $\mathcal{T}$-submagma of $\mathcal{A},$ which is $\mathcal{A}$ if and only if $\mathcal{A}$ is left admissible.
Proof.
Define the height function $h:\mathcal{A}\to\mathbb{N}\cup\{\infty\}$ inductively, as follows:
  • Every element of $\mathcal{T}$ has height 1.
  • $h(b)=m$, if $h(b)$ is not less than $m$ and $b=b_{1}*b_{2}$ where $h(b_{i})<m$ for $i=1,2.$
We show by induction on $h(b)$ that $ab\in\langle\mathcal{T}\rangle$ for $a\in\mathcal{T}$ and $b\in\langle\mathcal{T}\rangle.$ If $a,b\in\mathcal{T}$, then $ab\in\mathcal{T}$ by definition. We continue inductively, to see that if $b=b_{1}*b_{2}$ for $b_{1},b_{2}\in\langle\mathcal{T}\rangle$, then $ab=a(b_{1}*b_{2})=ab_{1}*ab_{2}\in\langle\mathcal{T}\rangle$. ∎
2.2. Prepairs and pairs
Definition 2.6.
A $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair $(\mathcal{A},\mathcal{A}_{0})$ (sometimes written $(\mathcal{A},\mathcal{A}_{0};*)$ if we want to specify the operation *) is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $(\mathcal{A},*,\iota)$ together with a $(\mathcal{T}_{1},\mathcal{T}_{2})$-sub-bimagma $\mathcal{A}_{0}$ i.e., $a_{1}b_{0}\in\mathcal{A}_{0}$ and $b_{0}a_{2}\in\mathcal{A}_{0}$ for all $a_{i}\in\mathcal{T}_{i}$ and $b_{0}\in\mathcal{A}_{0}.$
Definition 2.7.
We modify [2, 11] slightly.
  1. (i)
    A $\mathcal{T}$-pair $(\mathcal{A},\mathcal{A}_{0})$ is a weakly admissible $\mathcal{T}$-prepair, where $\mathcal{A}_{0}$ is closed under the operation *, i.e., $\mathcal{A}_{0}$ is a left $\mathcal{T}_{1}$-submagma of $\mathcal{A}$ [^3] In [1] $\mathcal{A}_{0}$ is not required to be closed under ()(*)( ∗ ). In previous work [1, 2, 11] we assumed that all pairs are proper, to dismiss the degenerate case of $\mathcal{T}=\mathcal{A}_{0}=\mathcal{A}.$ But this is precisely the case treated so successfully in [10], so we permit it here. We suppress $\mathcal{T}$ in the notation when it is understood.
  2. (ii)
    A $\mathcal{T}$-pair $(\mathcal{A},\mathcal{A}_{0})$ is said to be proper if $\mathcal{A}_{0}\cap\mathcal{T}=\emptyset$.
  3. (iii)
    A $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair $(\mathcal{A},\mathcal{A}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma prepair when $\mathcal{A}$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma.
  4. (iv)
    A gp-pair is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-pair for which $\mathcal{T}_{1},\mathcal{T}_{2}$ are groups.
2.3. Property N [1, §3.1]
For $b\in\mathcal{A}$, an element $b^{\dagger}$ is a pseudo-negative of $b$ if $b+b^{\dagger}=b^{\dagger}+b\in\mathcal{A}_{0}$, and $b^{\dagger}\in\mathcal{T}_{\iota}$ when $b\in\mathcal{T}_{\iota}$. Note that a pseudo-negative need not be unique.
Definition 2.8.
We say that a pair $(\mathcal{A},\mathcal{A}_{0})$ satisfies weak Property N if $\mathbb{1}$ has a pseudo-negative $\mathbb{1}^{\dagger}$, with $\mathbb{1}^{\dagger}\mathcal{T}\subseteq\mathcal{T}$ and $\mathbb{1}^{\dagger}b=b\mathbb{1}^{\dagger}$ for each $b\in\mathcal{A}$. In this case, fixing $\mathbb{1}^{\dagger}$, we define $e:=\mathbb{1}*\mathbb{1}^{\dagger}\in\mathcal{A}_{0}.$ Then, we denote $b^{\dagger}=b\mathbb{1}^{\dagger}$, and $b^{\circ}=b*b^{\dagger}=b^{\dagger}*b$, for all $b\in\mathcal{A}.$ Let $\mathcal{A}^{\circ}=\{b^{\circ}:b\in\mathcal{A}\},$ and $\mathcal{T}^{\circ}=\{a^{\circ}:a\in\mathcal{T}\}.$
Definition 2.9.
Let $(\mathcal{A},\mathcal{A}_{0})$ be a pair satisfying Property N.
  1. (i)
    A quasi-zero is an element of the form $a^{\circ},$ for $a\in\mathcal{T}.$
  2. (ii)
    $\mathcal{T}^{\circ}=\{a^{\circ}:a\in\mathcal{T}\}$.
  3. (iii)
    Define left and right actions of $\mathcal{T}\cup\mathcal{T}^{\circ}$ on $\mathcal{A},$ by defining $a^{\circ}b:=(ab)^{\circ}$ and $ba^{\circ}:=(ba)^{\circ}$ for $a\in\mathcal{T},$ $b\in\mathcal{A}$.
  4. (iv)
    $(\mathcal{A},\mathcal{A}_{0})$ is $\circ$-distributive if $(a_{1}*a_{2})e=a_{1}e*a_{2}e$ for all $a_{i}\in\mathcal{T}.$
  5. (v)
    $(\mathcal{A},\mathcal{A}_{0})$ is $\circ$-idempotent if $e*e=e.$
By [2, Lemma 3.9], if $\mathcal{A}$ is a $\mathcal{T}$-magma, then the action of (iii) makes $\mathcal{A}$ a $\mathcal{T}\cup\mathcal{T}^{\circ}$-magma, when $(\mathcal{A},\mathcal{A}_{0})$ is $\circ$-distributive and $\circ$-idempotent.
Example 2.10.
  1. (i)
    The pair $(R,\{\mathbb{0}\})$, where $R$ is as in Remark 2.3(i), is called the classical pair.
  2. (ii)
    The classical field pair is the classical gp-pair $(F,\{\mathbb{0}\})$, where $F$ is a field.
Important Note 2.11.
Philosophically, $\mathcal{A}_{0}$ takes the place of $\mathbb{0}$ (or, multiplicatively, $\mathbb{1}$) in classical mathematics. The significance is that since modules need not have negation (for example, $\mathbb{N}$), $\mathbb{0}$ has no significant role except as a place marker in linear algebra.
2.3.1. Homomorphisms of pairs
We consider $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas $\mathcal{A}$ and $(\mathcal{T}_{1}^{\prime},\mathcal{T}_{2}^{\prime})$-bimagmas $\mathcal{A}^{\prime}$, and functions $f:\mathcal{A}\to\mathcal{A}^{\prime}$ together with module homomorphisms $f_{\mathcal{T}_{i}}:\mathcal{T}_{i}\to\mathcal{T}_{i}^{\prime}.$ (When $\mathcal{A}$ is left weakly admissible then $f_{\mathcal{T}_{1}}$ is the restriction of $f$ to $\mathcal{T}_{1}.$ Likewise on the right.)
Definition 2.12.
  1. (i)
    A multiplicative map is a function $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying $f(\iota)=\iota,$ $f(a_{1}b)=f_{\mathcal{T}_{1}}(a_{1})f(b)$, and $f(ba_{2})=f(b)f_{\mathcal{T}_{2}}(a_{2})$ for all $a_{i}\in\mathcal{T}_{i},\ b\in\mathcal{A}.$
  2. (ii)
    The support $\operatorname{supp}(f)$ of a multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ is $\{b\in\mathcal{A}:f(b)\neq\iota^{\prime}\}.$
  3. (iii)
    A bimagma homomorphism (also just called homomorphism) $f:\mathcal{A}\to\mathcal{A}^{\prime}$ is a multiplicative map satisfying $f(b_{1}*b_{2})=f(b_{1})*f(b_{2}),$ $\forall b_{i}\in\mathcal{A}.$
  4. (iv)
    $\operatorname{Hom}(\mathcal{A},\mathcal{A}^{\prime})$ is the set of homomorphisms $f:\mathcal{A}\to\mathcal{A}^{\prime}$.
Remark 2.13.
In this paper $f_{\mathcal{T}_{i}}$ always is the identity map. Thus a multiplicative map satisfies $f(a_{1}b)=a_{1}f(b)$ and $f(ba_{2})=f(b)a_{2}$ for $a_{i}\in\mathcal{T}_{i}.$
Lemma 2.14.
For $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas, $\operatorname{Hom}(\mathcal{A},\mathcal{A}^{\prime})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma under the operation $(f_{1}*f_{2})(b)=f_{1}(b)*f_{2}(b)$ and the actions $af:b\mapsto af(b)$ and $fa:b\mapsto f(b)a$.
Proof.
Standard. ∎
Definition 2.15.
Suppose $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ are $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepairs.
  1. (i)
    A paired map $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying $f(\mathcal{A}_{0})\subseteq\mathcal{A}_{0}^{\prime}.$
  2. (ii)
    A paired homomorphism is a paired map which is a homomorphism.
Lemma 2.16.
  1. (i)
    Given a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair $(\mathcal{A},\mathcal{A}_{0})$ and any homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ , $(\mathcal{A}^{\prime},f(\mathcal{A}_{0}))$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair, and $f$ can be viewed as a paired homomorphism $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},f(\mathcal{A}_{0}))$.
  2. (ii)
    For any $(\mathcal{T}_{1},\mathcal{T}_{2})$-biomodule homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ where $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair, there is a prepair $(\mathcal{A},\mathcal{A}_{0})$ where $\mathcal{A}_{0}=\{b\in\mathcal{A}:f(b)\in\mathcal{A}_{0}^{\prime}\},$ and then $f$ is a paired homomorphism.
Proof.
(i) The image of $\mathcal{T}$ is disjoint from the image of $\mathcal{A}_{0}$.
(ii) If $a\in\mathcal{T}$ and $b\in\mathcal{A}_{0}$ then $f(ab)=f(a)f(b)\in\mathcal{A}_{0}^{\prime},$ so $ab\in\mathcal{A}_{0}.$ Then $f(\mathcal{A}_{0})\subseteq\mathcal{A}^{\prime}_{0},$ by definition. ∎
2.4. Pre-orders and surpassing relations on pairs
Definition 2.17.
  1. (i)
    A pre-order on a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $\mathcal{A}$, denoted $\preceq$, is a set-theoretic pre-order that respects the bimagma structure, i.e., for all $b,b_{i}\in\mathcal{A}$:
    1. (a)
      $b_{1}\preceq b_{2}$ implies $a_{1}b_{1}\preceq a_{1}b_{2}$ and $b_{1}a_{2}\preceq b_{2}a_{2}$ for $a_{i}\in\mathcal{T}_{i}.$
    2. (b)
      $b_{i}\preceq b_{i}^{\prime}$ implies $b_{1}*b_{2}\preceq b_{1}^{\prime}*b_{2}^{\prime}.$
  2. (ii)
    A surpassing relation on a $\mathcal{T}$-magma $\mathcal{A}$ is a pre-order satisfying the following:
    1. (a)
      $a_{1}\preceq a_{2}$ for $a_{1},a_{2}\in\mathcal{T}_{\iota}$ implies $a_{1}=a_{2}.$ (In other words, surpassing restricts to equality on $\mathcal{T}$.)
    2. (b)
      $b\preceq\iota$ for $b\in\mathcal{A}$ implies $b=\iota$.
  3. (iii)
    A pre-order (resp. surpassing relation) on a prepair $(\mathcal{A},\mathcal{A}_{0})$, is a pre-order (resp. surpassing relation) $\preceq$ satisfying the condition $\iota\preceq c$ for all $c\in\mathcal{A}_{0}$.
Lemma 2.18.
In a prepair $(\mathcal{A},\mathcal{A}_{0})$ with a pre-order, $b\preceq b*c$ and $b\preceq c*b$ for all $b\in\mathcal{A}$ and $c\in\mathcal{A}_{0}$.
Proof.
$b\preceq b$ and $\iota\preceq c,$ so $b=b*\iota\preceq b*c.$
Lemma 2.19 ([11, Lemma 2.11]).
If $b_{1}\preceq b_{2}$ holds in a pre-ordered prepair with a negation map, then $b_{2}(-)b_{1}\succeq\iota$ and $b_{1}(-)b_{2}\succeq\iota.$
Surpassing relations were introduced in [24], and in [2] for pairs, for the purposes of linear algebra.
2.4.1. $\preceq$-morphisms
We insert the surpassing relation into our categories.
Definition 2.20.
Let $(\mathcal{A},\mathcal{A}_{0})$, $(\mathcal{A}^{\prime},\mathcal{A}_{0}^{\prime})$ be prepairs.
  1. (i)
    When $\mathcal{A}$ also has a pre-order $\preceq$, a $\preceq$-morphism, (analogous to “colax morphism” in [20]) is a multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying the following conditions:
    • $f(b_{1})\preceq f(b_{2}),\quad\forall b_{1}\preceq b_{2}\in\mathcal{A}.$
    • $f(b_{1}*b_{2})\preceq f(b_{1})*f(b_{2}),\quad\forall b_{1},b_{2}\in\mathcal{A}.$
    ${\operatorname{Mor}_{\preceq}}(\mathcal{A},\mathcal{A}^{\prime})$ is the set of $\preceq$-morphisms from $\mathcal{A}$ to $\mathcal{A}^{\prime}.$
  2. (ii)
    A $\succeq$-morphism, (analogous to “lax morphism” in [20]) is a multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying the following conditions:
    • $f(b_{1})\preceq f(b_{2}),\quad\forall b_{1}\preceq b_{2}\in\mathcal{A}.$
    • $f(b_{1}*b_{2})\succeq f(b_{1})*f(b_{2}),\ \forall b_{1},b_{2}\in\mathcal{A}.$
Lemma 2.21.
Given an 1:1 multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ where $\mathcal{A}$ is a $\mathcal{T}$-magma and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a pair with surpassing relation $\preceq$, the pair $(\mathcal{A},\mathcal{A}_{0})$ of Lemma 2.16 has a surpassing relation given by $b_{1}\preceq b_{2}$ when $f(b_{1})\preceq f(b_{2}),$ and then $f$ becomes a $\preceq$-morphism.
Proof.
$a_{1}\preceq a_{2}$ implies $f(a_{1})\preceq f(a_{2}),$ so $f(a_{1})=f(a_{2}),$ so $a_{1}=a_{2}.$
2.5. Negation maps
At times we can define the negation map, the mainstay of [24]. A negation map ()(-)( - ) on a $\mathcal{T}$-module $(\mathcal{A},\mathcal{A}_{0})$ is an additive automorphism ()(-)( - ) of order $\leq 2$ also defined on $\mathcal{T}$, such that
$(-)(ab)=((-)a)b=a((-)b),\qquad\forall a\in\mathcal{T},\quad b\in\mathcal{A},$
and $(-)\mathcal{A}_{0}=\mathcal{A}_{0}.$
We write $b_{1}(-)b_{2}$ for $b_{1}+((-)b_{2}).$ Thus $b^{\circ}=b(-)b,$ and $\mathcal{A}_{0}$ contains the set $\mathcal{A}^{\circ}=\{b^{\circ}:b\in\mathcal{A}\}.$ Often $\mathcal{A}_{0}=\mathcal{A}^{\circ}.$
Lemma 2.22.
The negation map satisfies $(-)b=((-)\mathbb{1})b.$
Proof.
$(-)b=(-)(\mathbb{1}b)=((-)\mathbb{1})b.$
Hence, perhaps surprisingly, if $b_{1}\preceq b_{2},$ then $(-)b_{1}\preceq(-)b_{2}$.
2.6. Bimodules
We pinpoint the additively associative case.
Definition 2.23.
  1. (i)
    A left $\mathcal{T}$-module $\mathcal{M}$ is a left $\mathcal{T}_{\iota}$-magma where * is also abelian and associative, in which case we write $+$ in place of *, and $\mathbb{0}$ instead of $\iota$.
  2. (ii)
    A left $\mathcal{T}$-module $\mathcal{M}$ is free with base $\{b_{i}:i\in I\}$, if every element of $\mathcal{M}$ can be written uniquely in the form $\sum_{i\in I}a_{i}b_{i}$ for $a_{i}\in\mathcal{T}_{\iota}.$
  3. (iii)
    A $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimodule is a $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimagma $(\mathcal{M},+)$ which is also abelian and associative.
  4. (iv)
    Definitions 2.4 and 2.7 also apply to $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodules.
Remark 2.24.
For $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodules, $\operatorname{Hom}(\mathcal{A},\mathcal{A}^{\prime})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule under the action of Remark 2.14.
Proof.
Standard. ∎
Just as in classical ring theory and module theory, the category theory for semialgebras differs from the category theory for bimodules, we want to differentiate between $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimodules (resp. bimagmas) and $\mathcal{T}$-semialgebras (resp. pre-semialgebras), which we now define.
Definition 2.25.
  1. (i)
    A $\mathcal{T}$-pre-semialgebra is a $(\mathcal{T},\mathcal{T})$-bimagma (with respect to *) where $\mathcal{T}$ is $\mathcal{A}$-central, which also has a second operation which we denote as $\cdot$, with $a_{1}(b_{1}\cdot b_{2})=(a_{1}b_{1})\cdot b_{2}$ and $(b_{1}\cdot b_{2})a_{2}=b_{1}\cdot(b_{2}a_{2})$ for $a_{i}\in\mathcal{T},$ $b_{1},b_{2}\in\mathcal{A}.$ In particular $a\cdot b=ab$ and $b\cdot a=ba$ for all $a\in\mathcal{T}$ and $b\in\mathcal{A}.$
  2. (ii)
    A $\mathcal{T}$-semialgebra is a $\mathcal{T}$-pre-semialgebra which is a semiring, where * is the usual addition.
  3. (iii)
    A pre-semialgebra homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ of $\mathcal{T}$-pre-semialgebras is a homomorphism also satisfying $f(b_{1}b_{2})=f(b_{1})f(b_{2})$ for all $b_{i}\in\mathcal{A}.$
  4. (iv)
    A pre-semialgebra $\mathcal{A}$ with a surpassing relation $\preceq$ is $\preceq$-distributive if $b(b_{1}*b_{2})\preceq bb_{1}\,*\,b_{2}b_{2}$ and $(b_{1}*b_{2})b\preceq b_{1}b\,*\,b_{2}b$ for all $b,b_{i}\in\mathcal{A}.$
2.6.1. Bimodule-theoretic notions for prepairs
The following notions are difficult to notate for bimagmas (which need not be associative), so we introduce them only for bimodules.
2.6.2. Weak morphisms
Definition 2.26.
A weak morphism of bimodule prepairs is a paired multiplicative map $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$, satisfying $\sum b_{i}\in\mathcal{A}_{0}$ for $b_{i}\in\mathcal{A}$ implies $\sum f(b_{i})\in\mathcal{A}^{\prime}_{0}$.
$\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime})$ denotes the set of weak morphisms from $(\mathcal{A},\mathcal{A}_{0})$ to $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0}).$
Lemma 2.27.
For prepairs $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0}),$ define $\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime})_{0}$ to be the weak morphisms for which $f(\mathcal{A})\subseteq\mathcal{A}^{\prime}_{0}.$
  1. (i)
    $(\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime}),\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime})_{0})$ is a bimodule prepair.
  2. (ii)
    For pairs $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0}),$ $(\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime}),\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime})_{0})$ is a bimodule pair, where we define $(f_{1}+f_{2})(b)=f_{1}(b)+f_{2}(b),$ $(af)(b)=f(ab)$, $(fa)(b)=f(b)a$.
Proof.
The main verification for (ii) is that the sum $f_{1}+f_{2}$ of weak morphisms is a weak morphism. If $\sum b_{i}\in\mathcal{A}_{0},$ then $(f_{1}+f_{2})(\sum b_{i})=f_{1}(\sum b_{i})+f_{2}(\sum b_{i})\in\mathcal{A}_{0}+\mathcal{A}_{0}\subseteq\mathcal{A}_{0}.$
Lemma 2.28 (As in [1, Lemma 2.10]).
Every $\preceq$-morphism of prepairs is a weak morphism.
3. Hypermagmas and hyperpairs
We follow the treatment of Nakamura and Reyes [20], casting hypergroups into a broader context which has many more interesting examples.
Definition 3.1 ([20], generalizing [17]).
  1. (i)
    $\mathcal{P}(\mathcal{H})$ denotes the power set of a set $\mathcal{H},$ and ${\mathcal{P}}^{\sharp}(\mathcal{H})=\mathcal{P}(\mathcal{H})\setminus\emptyset.$
    A hypermagma $(\mathcal{H},*)$ is a set with a binary operation :×$*:\mathcal{H}\times\mathcal{H}\to\mathcal{P}(\mathcal{H})$, permitting the product of two elements to be the empty set [^4] Customarily one requires $a_{1}*a_{2}\in{\mathcal{P}}^{\sharp}(\mathcal{H})$ for all $a_{1},a_{2}\in\mathcal{H}.$ But one could formally adjoin $\emptyset$, so the Nakamura-Reyes definition subsumes the customary definition, and we shall need it in what follows. $\emptyset$ could be interpreted as “undefined.”, together with a left and right $\mathcal{H}$-action whose natural extension to $\mathcal{P}(\mathcal{H})$ makes $(\mathcal{P}(\mathcal{H}),*)$ an $\mathcal{H}$-bimagma when we define, for $S_{1},S_{2}\in\mathcal{P}(\mathcal{H}),$
    $S_{1}*S_{2}:=\cup_{s_{i}\in S_{i}}\ s_{1}*s_{2},\qquad\emptyset*S=S*\emptyset=\emptyset.$
    We view $\mathcal{H}\subseteq\mathcal{P}(\mathcal{H})$ by identifying $a$ with $\{a\}$.
    From now on we follow the customary use of $\boxplus$ and $\mathbb{0}$ instead of * and $\iota$, when the operation is associative. We call $\boxplus$ “hyperaddition.”
  2. (ii)
    A hyperzero $\mathbb{0}$ (if it exists) satisfies $\mathbb{0}\boxplus a=a=a\boxplus\mathbb{0},$ for every $a\in\mathcal{H}.$
  3. (iii)
    A hypersemigroup is a hypermagma $(\mathcal{H},\boxplus,\mathbb{0})$, where
    1. (a)
      The hyperaddition $\boxplus$ is associative in the sense that $(a_{1}\boxplus a_{2})\boxplus a_{3}=a_{1}\boxplus(a_{2}\boxplus a_{3})$ for all $a_{i}$ in $\mathcal{H}.$
    2. (b)
      $\mathbb{0}\in\mathcal{H}$ is the hyperzero.
  4. (iv)
    A hypernegative of an element $a$ in $(\mathcal{H},\boxplus,\mathbb{0})$ (if it exists) is an element “$-a$” for which $\mathbb{0}\in a\boxplus(-a)$ and $\mathbb{0}\in(-a)\boxplus a$.
  5. (v)
    If the hypernegative $-\mathbb{1}$ exists in $\mathcal{H},$ then we define $e=\mathbb{1}\boxplus(-\mathbb{1}).$
  6. (vi)
    A hypergroup is a hypersemigroup $(\mathcal{H},\boxplus,\mathbb{0})$ for which every element $a$ has a unique hypernegative denoted $-a$, whereby, for all $a_{i}\in\mathcal{H},$
    1. (a)
      $(-)(a_{1}\boxplus a_{2})=(-)a_{2}\boxplus(-)a_{1}.$
    2. (b)
      $-(-a_{1})=a_{1}.$
    3. (c)
      $\mathcal{H}$ is reversible in the following sense:
      $a_{3}\in a_{1}\boxplus a_{2}$ iff $a_{2}\in a_{3}\boxplus(-a_{1})$.p [^5] In [20], a mosaic is a reversible with a hyperzero.
  7. (vii)
    A hypersemiring (resp. hyperring) is an abelian [^6] In [20] this is called “canonical.” hypersemigroup (resp. hypergroup) $(\mathcal{H},\boxplus,\mathbb{0})$, providing $\mathcal{P}(\mathcal{H})$ with a natural elementwise multiplication, distributing over hyperaddition, making $\mathcal{P}(\mathcal{H})$ a pre-semialgebra pair over $\mathcal{T}=\mathcal{H}\setminus\{\mathbb{0}\}$ under the action $aS=\{as:s\in S\}$. [^7] In general $\mathcal{P}(\mathcal{H})$ is not distributive, but does satisfy $(\boxplus_{i}S_{i})(\boxplus_{j}S_{j}^{\prime})\subseteq(\boxplus_{i}S_{i}S_{j}^{\prime})$, cf. [18, Proposition 1.1].
  8. (viii)
    A hypergroup $\mathcal{H}$ is a hyperfield if $\mathcal{H}\setminus\{\mathbb{0}\}$ is a multiplicative group, with $\mathbb{0}a=\mathbb{0}$ for all $a\in\mathcal{H}.$
Remark 3.2 ([1, 24]).
  1. (i)
    If it exists, the hyperzero of a hypersemigroup is unique, by the familiar elementary argument. We write $\mathcal{H}$ for the $(\mathcal{H},\boxplus,\mathbb{0})$ in the rest of this remark.
  2. (ii)
    $\mathcal{H}$ gives rise to a $\mathcal{H}$-magma $\mathcal{P}(\mathcal{H}),$ with hyperaddition given by
    $S_{1}\boxplus S_{2}=\cup\{s_{1}\boxplus s_{2}:s_{i}\in S_{i}\}.$
    Thus $\emptyset$ is the $\infty$ element in the sense of Definition 2.1.
  3. (iii)
    $\mathcal{H}$ is a weakly admissible $\mathcal{T}$-bimagma in case
    1. (a)
      $\mathcal{T}=\{\mathbb{1}\}$.
    2. (b)
      $\mathcal{T}$ is a submonoid of $\mathcal{H}$, in particular when $\mathcal{H}\setminus\{\mathbb{0}\}$ is a monoid.
  4. (iv)
    Take any $\mathcal{T}$-submodule $S_{0}$ of $\mathcal{A}:=\mathcal{P}(\mathcal{H})$. Then we get a $\mathcal{T}$-pair $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$ where $\mathcal{P}(\mathcal{H})_{0}=\{S\subseteq\mathcal{H}:S_{0}\cap S\neq\emptyset\}.$
  5. (v)
    Take $S_{0}=\{\mathbb{0}\}$ in (iv). Thus $\mathcal{P}(\mathcal{H})_{0}=\{S\subseteq\mathcal{H}:\mathbb{0}\in S\},$ the standard definition.
Lemma 3.3.
Say a subset $S\subseteq\mathcal{P}(\mathcal{H})$ is weakly neutral if $a\in a+S$ for each $a\in\mathcal{H}$. The weakly neutral subsets comprise a submagma of $\mathcal{P}(\mathcal{H})$.
Proof.
If $S_{i}$ are weakly neutral elements, then for any $a\in\mathcal{H},$ $a\boxplus(S_{1}\boxplus S_{2})=(a\boxplus S_{1})\boxplus S_{2}$, which contains $a\boxplus S_{2},$ which in turn contains $a.$
Definition 3.4.
In Remark 3.2(v), the sub-pair of the pair $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$ generated by $\mathcal{H}$ is called the hyperpair of $\mathcal{H}.$ [^8] This is the definition in [2], but although it could lead to difficulties if $\mathcal{P}(\mathcal{H})$ is not distributive, because then the $\mathcal{H}$-submagma of the power set $\mathcal{P}(\mathcal{H})$ spanned by $\mathcal{H}$ need not be closed under multiplication. Alternatively, one could take $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$, but it need not be admissible.
Remark 3.5.
  1. (i)
    Any hyperpair has the important surpassing relation $\subseteq$,⊆ , i.e., $S_{1}\preceq S_{2}$ when $S_{1}\subseteq S_{2}.$ Of course $\mathcal{P}(\mathcal{H})$ is a lower semilattice with respect to ()($\subseteq$)( ⊆ ), but not ${\mathcal{P}}^{\sharp}(\mathcal{H}).$
  2. (ii)
    More generally, if $\mathcal{H}$ has a surpassing relation $\preceq$, then $\mathcal{P}(\mathcal{H})$ has the surpassing relation $S_{1}\preceq S_{2}$ if for each $s_{1}\in S_{1}$ there is $s_{2}\in S_{2}$ for which $s_{1}\preceq s_{2}.$
  3. (iii)
    For hypermagmas $\mathcal{H}$, the condition that $b_{1}\subseteq b_{2}$ implies $f(b_{1})\subseteq f(b_{2})$ is automatic on $\mathcal{P}(\mathcal{H})$.
  4. (iv)
    If each $a\in\mathcal{H}$ has a unique hypernegative, then $(\mathcal{A},\mathcal{A}_{0})$ has a negation map given by applying the hypernegative element-wise.
3.1. Examples of hypersemigroups and their pairs
Example 3.6.
Let us first recall some of the famous hyperfields (and their accompanying hyperpairs, cf. Definition 3.4), from [1]. In every case the sum of two distinct nonzero elements is not a singleton.
  1. (i)
    The tropical hyperfield consists of the set ${}\mathcal{H}=\mathbb{R}\cup\{-\infty\}$, with $\infty$- ∞ here as the zero element and 00 as the unit, equipped with the addition $a\boxplus b=\{a\}$ if $a>b$, $a\boxplus b=\{b\}$ if $a<b$, and $a\boxplus a=[-\infty,a]$.
  2. (ii)
[…]
  3. (iii)
    The phase hyperfield. Let $S^{1}$ denote the complex unit circle, and take $\mathcal{H}=S^{1}\cup\{0\}$. Nonzero points $a$ and $b$ are antipodes if $a=-b.$ Multiplication is defined as usual (so corresponds on $S^{1}$ to addition of angles). We denote an open arc of less than 180 degrees connecting two distinct points $a,b$ of the unit circle by $({a,b})$. The hypersum is given, for $a,b\neq 0$, by
    $a\boxplus b=\begin{cases}({a,b})\text{ if }a\neq\pm b;\\ \{-a,0,a\}\text{ if }a=-b\enspace,\\ \{a\}\text{ if }a=b\enspace.\end{cases}$
    The hypersystem $\overline{\mathcal{H}}$ contains only the points of $\mathcal{H}$, the elements of the form $a\boxplus b$ with $a,b\in S^{1}$, and subsets $C$ of $\mathcal{H}$, where either $C$ is an open half circle connecting an element $a\in S^{1}$ to $-a$ (which is obtained as the sum $a\boxplus b\boxplus(-a)$, where $b\in C$), or $C=\mathcal{H}$ (which is obtained as $a\boxplus b\boxplus(-a)\boxplus(-b)$).
    $e=\mathbb{1}\boxplus-\mathbb{1}=\{-\mathbb{1},0,\mathbb{1}\},$ so $e\boxplus e=e$. Thus the phase hyperfield is $\circ$-idempotent, but is not $\circ$-distributive since $(a,b)e=(a,b)\cup\{\mathbb{0}\}\cup(-a,-b),$ whereas $(a,b)\boxplus(-a,-b)$ is the whole circle.
  4. (iv)
    (The weak phase hyperfield [1]; Akian-Gaubert’s modification of the phase hyperfield). The non-zero elements can still be represented by elements of the unit circle $S^{1}$. However, the hyperaddition differs, for $a,b\neq 0$, we now have
    $a\boxplus b=\begin{cases}{[a,b]}&\text{ if }a\neq b\text{ and }a\neq-b,\\ S^{1}\cup\{0\}&\text{ if }a=-b,\\ \{a\}&\text{ if }a=b.\end{cases}$
    where ${[a,b]}$ denotes the closed arc of the angle inferior to $180$ degrees joining $a$ and $b$ (compare with the open arc $(a,b)$ in the phase hyperfield). As the phase hyperfield, the weak phase hyperfield is not distributive.
  5. (v)
    One can further modify the weak phase hyperfield, to get a hypersemigroup which is not a hypergroup by defining $a\boxplus a=\emptyset.$
Example 3.7.
Let $\mathcal{H}$ be any nonempty set. Stipulate a hyperzero $\mathbb{0}\in\mathcal{H}.$ In each case, its hyperpair $(\mathcal{A},\mathcal{A}_{0})$ will satisfy property N.
  1. (i)
    The following are uniquely negated, with $\mathcal{A}_{0}$ as in Remark 3.2(v).
    1. (a)
      Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\{\mathbb{0},s\}$, $\mathbb{0}\boxplus s=s\boxplus\mathbb{0}=s,$ and $s\boxplus s^{\prime}=\mathcal{H}$ for all nonzero $s\neq s^{\prime}$ in $\mathcal{H}.$ Here $\mathcal{A}=\mathcal{H}\cup\{\mathcal{H}\},$ and $-s=s.$
    2. (b)
      (As in [18]) Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\mathcal{H}\setminus s$ and $s\boxplus s^{\prime}=\{s,s^{\prime}\}$ for all nonzero $s\neq s^{\prime}$ in $\mathcal{H}.$ This is a hypergroup when $|\mathcal{H}|\geq 3.$ Here $\mathcal{A}=\mathcal{H}\cup_{s,s^{\prime}\in\mathcal{H}}\{s,s^{\prime}\}\cup_{s\in\mathcal{H}}(H\setminus\{s\})\cup\{\mathcal{H}\}.$ Again $-s=s.$
    3. (c)
      Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s^{\prime}=\mathcal{H}\setminus\{s,s^{\prime}\}$ for all nonzero $s,s^{\prime}\in\mathcal{H}.$ This is a hypergroup when $|\mathcal{H}|\geq 4.$ Here $\mathcal{A}=\mathcal{H}\cup_{s,s^{\prime}\in\mathcal{H}}\{s,s^{\prime}\}\cup_{s\in\mathcal{H}}(H\setminus\{s\})\cup_{s,s^{\prime}\in\mathcal{H}}(H\setminus\{s,s^{\prime}\})\cup\{\mathcal{H}\}.$ (Here $H\setminus\{s\}=(s^{\prime}\boxplus s^{\prime\prime})\boxplus s)$ where $s^{\prime},s^{\prime\prime}$ are distinct from $\{\mathbb{0},s\}.$)
  2. (ii)
    Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=s$ and $s\boxplus s^{\prime}=\mathcal{H}$ for all $s\neq s^{\prime}$ in $\mathcal{H}.$ This is idempotent, but is not uniquely negated.
  3. (iii)
    The following two examples are similar to those of [20, Proposition 3.1]. For all $a_{i}\in\mathcal{H},$
    1. (a)
      $a_{1}\boxplus a_{2}=\mathcal{H}$. This gives the admissible pair ({},$(\mathcal{H}\cup\{\mathcal{H}\},\{\mathbb{0},\{\mathcal{H}\})$ which satisfies Property N.
    2. (b)
      $a_{1}\boxplus a_{2}=\emptyset$. $(\mathcal{H}\cup\{\emptyset\},\emptyset)$ is the minimal pair containing $\mathcal{H}$, but is not a hyperpair.
  4. (iv)
    $a_{1}\boxplus a_{2}=\{a_{1},a_{2}\}$. Hence $S_{1}\boxplus S_{2}=S_{1}\cup S_{2}$. $(\mathcal{H},\mathcal{H}_{0})$ fails Property N.
  5. (v)
    Now suppose $\mathcal{H}$ is ordered, with $\mathbb{0}\in\mathcal{H}$ minimal. Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\mathcal{H}$ and $s\boxplus s^{\prime}=s^{\prime}$ for all $\mathbb{0}\neq s<s^{\prime}$ in $\mathcal{H}.$ This is an idempotent hypergroup giving rise to an admissible hyperpair $\mathcal{A}=\mathcal{H}\cup\{\mathcal{H}\},$ and $\mathcal{A}_{0}=\{\mathcal{H}\}$.
In each case, $\mathcal{H}$ becomes a hyper-semiring when $\mathcal{H}$ is a cancellative multiplicative monoid.
3.2. Residue (quotient) hypermodules and hyperpairs
The following definition was inspired by Krasner [14], a bit more general than [25, Definition 3.1], but with the same verification as [25, Lemma 3.2].
Definition 3.8.
Suppose that $\mathcal{M}$ is a $C$-bimodule where the ring $C$ is viewed as a multiplicative monoid $\mathcal{T}$ as in Remark 2.3(i), and $G$ is a normal multiplicative subgroup of $\mathcal{T}$, in the sense that $bG=Gb$ for all $a\in\mathcal{M}$. Define the residue hypermodule $\mathcal{H}=\mathcal{M}/G$ over $\mathcal{T}/G$ to have multiplication induced by the cosets, and hyperaddition :×$\boxplus:\mathcal{H}\times\mathcal{H}\to\mathcal{P}(\mathcal{M})$ by
$b_{1}G\boxplus b_{2}G=\{cG:c\in b_{1}G+b_{2}G\}.$
When $\mathcal{M}=C$ is a field, the residue hypermodule $\mathcal{H}$ is a hyperfield (under multiplication of cosets), called the quotient hyperfield in the literature.
As defined in Remark 3.2(iv), $(\mathcal{M},\mathcal{M}_{0})$ is a pair, by [1, Proposition 7.21]. For the applications in the hyperfield literature one would take the classical field $\mathcal{M}$ with a multiplicative subgroup $\mathcal{T}$.
Remark 3.9.
In the residue hypermodule $\mathcal{H}=\mathcal{M}/G$,
  1. (i)
    $e=\mathbb{1}\boxplus(-)\mathbb{1}=\{g_{1}-g_{2}:g_{i}=G\}.$
  2. (ii)
    According to Definition 2.8(iv), $Se=\{b_{1}g_{1}-b_{2}g_{2}:b_{i}\in S,g_{i}\in G\}$
  3. (iii)
    In particular, $ee=\{(g_{1}-g_{2})g_{3}\boxplus(g_{4}-g_{5})g_{6}:g_{i}=G\}.$
  4. (iv)
    $e\boxplus e=\{(g_{1}-g_{2})-(g_{4}-g_{5}):g_{i}=G\}.$
$\mathcal{M}/G$ need not be $\circ$-distributive, since the phase hyperfield is a counterexample, cf. Examples 3.6(iii), but $\mathcal{M}/G$ does satisfy $ee=e\boxplus e$ by [25, Lemma 3.5].
Lemma 3.10.
Any surpassing relation $\preceq$ on $\mathcal{M}$ induces a surpassing relation on $\mathcal{M}/G$.
Proof.
As in Remark 3.5, define $b_{1}G\preceq b_{2}G$ when for each $g\in G$ there is $g^{\prime}\in G$ such that $b_{1}g\preceq b_{2}g$. We claim that if $b_{i}G\preceq b_{i}^{\prime}G$ for $i=1,2,$ then $b_{1}G+b_{2}G\preceq b_{1}^{\prime}G+b_{2}^{\prime}G.$
Indeed, if $b=b_{1}g_{1}+b_{2}g_{2}\in b_{1}G+b_{2}G,$ then taking $b_{i}^{\prime}g_{i}^{\prime}\succeq b_{i}g_{i}$ in $b_{i}^{\prime}G,$ we have
$b=b_{1}g_{1}+b_{2}g_{2}\preceq=b_{1}^{\prime}g_{1}^{\prime}+b_{2}^{\prime}g_{2}^{\prime}\in b_{1}^{\prime}G+b_{2}^{\prime}G.$
Example 3.11.
A huge assortment of examples of quotient hyperfields is given in [19, §2] [^9] In [25] more examples are obtained when addition and multiplication are reversed.. Here are some of them. We shall take $\mathcal{H}=\mathcal{M}/G,$ and its hyperpair $(\mathcal{A},\mathcal{A}_{0}),$ as in Remark 3.2(iv).
  1. (i)
    $G=\{\pm 1\}$. Then $\mathbb{0}\in\mathcal{H}\boxplus\mathcal{H},$ so $(\mathcal{A},\mathcal{A}_{0})$ has characteristic 2 and is multiplicatively idempotent.
  2. (ii)
    The Krasner hyperfield is $F/F^{\times}$, for any field $F$.
  3. (iii)
    The sign hyperfield is $\mathbb{R}/\mathbb{R}^{+}.$
  4. (iv)
    The phase hyperfield can be identified with the quotient hyperfield $\mathbb{C}/\mathbb{R}_{>0}$.
  5. (v)
    The weak phase hyperfield can be obtained by taking the quotient $\mathcal{F}/G$, where =$\mathcal{F}=\mathbb{C}\{\{t^{\mathbb{R}}\}\}$, and $G$ is the group of (generalized) Puiseux series with positive real leading coefficient, where the leading coefficient is the coefficient $f_{\lambda}$ of the series $f=\sum_{\lambda\in\Lambda}f_{\lambda}t^{\lambda}$ such that $\lambda$ is the minimal element of $\{\lambda\in\Lambda:f_{\lambda}\neq 0\}$.
4. Tensor products of bimodules and bimagmas
Remark 4.1.
Let us start with the observation that the classical treatment of a right module and a left module over a ring $C$ only uses the fact that $C$ is a monoid. So we shall talk of tensor products over a multiplicative monoid.
In [12] tensor products were treated from the point of view of semirings, following [13, 26]. There is a delicate issue here, concerning which are the morphisms in our category. One may start with the more straightforward instance of homomorphisms. Then it is rather easy to construct a tensor product satisfying the theory of [9, 21], and their theorems are applicable.
When we bring in weak morphisms on both sides, things become much more complicated, and our approach must become more compatible with [4] and [20].
4.1. Construction of the tensor product
Example 4.2.
  1. (i)
    The free magma $(\mathcal{F}(X),*)$ on a set $X$ is the set of formal expressions with parentheses, defined by $X\subseteq\mathcal{F}(X)$ of length 1, and inductively, for $v,w\in\mathcal{F}(X)$ of lengths $m$ and $n$, $(v*w)\in\mathcal{F}(X)$ of length $m+n.$
  2. (ii)
    The free abelian semigroup $(\mathcal{S}(X),+)$ on a set $X$ is the set of formal sums without parentheses, defined by $X\subseteq\mathcal{S}(X)$ of length 1, and inductively, for $v,w\in\mathcal{S}(X)$ of lengths $m$ and $n$, $v+w=w+v\in\mathcal{F}(X)$ of length $m+n.$
We follow the exposition of the classical tensor product, as in [23, Chapter 18]. Throughout this section, $\mathcal{M}_{1}$ is a right $\mathcal{T}$-module and $\mathcal{M}_{2}$ is a left $\mathcal{T}$-module, for convenience.
Definition 4.3.
  1. (i)
    Define the $\mathcal{T}$-tensor product semigroup [^10] More generally, in parallel to [20], for $\mathcal{M}_{1}$ a right $\mathcal{T}$-magma and $\mathcal{M}_{1}$ a left $\mathcal{T}$-magma, we can define the $\mathcal{T}$-tensor product magma $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ to be the magma $\mathcal{F}(\mathcal{M}_{1}\times\mathcal{M}_{2})/\Phi,$ where $\Phi$ is the congruence generated by all $\bigg{(}\big{(}v_{1}*w_{1},x_{2}\big{)},\big{(}v_{1},x_{2})*(w_{1},x_{2}\big{)}\bigg{)},\quad\bigg{(}\big{(}x_{1},v_{2}*w_{2}\big{)},(x_{1},v_{2})*(x_{1},w_{2})\bigg{)},$ (4.1) $\bigg{(}(x_{1}a,x_{2}),(x_{1},ax_{2})\bigg{)}$ (4.2) $\forall x_{i},v_{i},w_{i}\in\mathcal{\mathcal{M}}_{i},$ $a\in\mathcal{T}$. The remainder of this section would go over, mutatis mutandis, but the notation is less familiar. $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ to be the semigroup $\mathcal{S}(\mathcal{M}_{1}\times\mathcal{M}_{2})/\Phi,$ where $\Phi$ is the congruence generated by all
    $\bigg{(}\big{(}v_{1}+w_{1},x_{2}\big{)},\big{(}v_{1},x_{2})+(w_{1},x_{2}\big{)}\bigg{)},\quad\bigg{(}\big{(}x_{1},v_{2}+w_{2}\big{)},(x_{1},v_{2})+(x_{1},w_{2})\bigg{)},$ (4.3)
    $\bigg{(}(x_{1}a,x_{2}),(x_{1},ax_{2})\bigg{)}$ (4.4)
    $\forall x_{i},v_{i},w_{i}\in\mathcal{\mathcal{M}}_{i},$ $a\in\mathcal{T}$.
  2. (ii)
    When $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ have negation maps, we incorporate the negation map into the definition of tensor product, by enlarging $\Phi$ to include also $(\big{(}(-)v_{1},v_{2}\big{)},\big{(}v_{1},(-)v_{2})\big{)}$ for all $v_{i}\in\mathcal{F}_{i}.$
  3. (iii)
    A simple tensor of $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is an element of the form $v\otimes w$ for $v\in\mathcal{M}_{1}$ and $w\in\mathcal{M}_{2}$.
Note that this matches the situation that $\mathcal{T}$ is a semiring, in which case we would just forget its addition.
Remark 4.4.
  1. (i)
    The sum of two non-simple tensors could be simple, modulo associativity. For example, if $v_{3}=v_{1}+v_{2}$ then $(v_{3}\otimes v_{2})+(v_{2}\otimes v_{1})$ and $(v_{2}\otimes v_{2})+(v_{1}\otimes v_{3})$ are non-simple whereas
    $(v_{3}\otimes v_{2})+(((v_{2}\otimes v_{1})+(v_{2}\otimes v_{2}))+(v_{1}\otimes v_{3}))$ $=(v_{3}\otimes v_{2})+((v_{2}\otimes v_{3}))+(v_{1}\otimes v_{3}))$ (4.5)
    $=(v_{3}\otimes v_{2})+(v_{3}\otimes v_{3})=v_{3}\otimes(v_{2}+v_{3}).$
  2. (ii)
    If $a_{1}w_{1}=a_{2}w_{2},$ then $v_{1}a_{1}\otimes w_{1}+v_{2}a_{2}\otimes w_{2}=(v_{1}+v_{2})\otimes a_{2}w_{2}$. This process of “recombining” is straightforward when $\mathcal{T}$ is a group, since then it is applicable for any sum
    $v_{1}\otimes w_{1}+v_{2}\otimes w_{2}=(v_{1}a_{1}^{-1}+v_{2}a_{2}^{-1})\otimes a_{2}w_{2}.$
Remark 4.5.
As in [4], one could define the tensor product $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$ of a right $\mathcal{T}$-monoid $\mathcal{T}_{1}$ and a left $\mathcal{T}$-monoid $\mathcal{T}_{2}$ to be $(\mathcal{T}_{1}\times\mathcal{T}_{2})/\Phi$ where where $\Phi$ is the congruence generated by all $\bigg{(}(x_{1}a,x_{2}),(x_{1},ax_{2})\bigg{)}$, $\forall x_{i},v_{i},w_{i}\in\mathcal{\mathcal{M}}_{i},$ $a\in\mathcal{T}$.
4.1.1. Tensor products of prepairs
In order to define the tensor product prepair, we need to define $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}$.
Definition 4.6.
  1. (i)
    Suppose $(\mathcal{M}_{1},{\mathcal{M}_{1}}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule prepair and $(\mathcal{M}_{2},{\mathcal{M}_{2}}_{0})$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule prepair. Define $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}:=({\mathcal{M}_{1}}_{0}\otimes_{\mathcal{T}}\mathcal{M}_{2})\cup(\mathcal{M}_{1}\otimes_{\mathcal{T}}{\mathcal{M}_{2}}_{0}).$
  2. (ii)
    Suppose $(\mathcal{M}_{1},{\mathcal{M}_{1}}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule pair and $(\mathcal{M}_{2},{\mathcal{M}_{2}}_{0})$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule pair. Define $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}:=({\mathcal{M}_{1}}_{0}\otimes_{\mathcal{T}}\mathcal{M}_{2})+(\mathcal{M}_{1}\otimes_{\mathcal{T}}{\mathcal{M}_{2}}_{0}).$
Lemma 4.7.
$((\mathcal{M}_{1}\otimes\mathcal{M}_{2}),(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair in (i), and a pair in (ii).
Proof.
$((\mathcal{M}_{1}\otimes\mathcal{M}_{2}),(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0})$ is closed under left multiplication by $\mathcal{T}_{1}$ and right multiplication by $\mathcal{T}_{2}.$ $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}$ is closed under addition in (ii). ∎
Write $\mathcal{T}^{\operatorname{op}}$ for the opposite monoid. Then any $\mathcal{T}_{1},\mathcal{T}_{2}$ bimodule is a $\mathcal{T}_{1}\times\mathcal{T}_{2}^{\operatorname{op}}$ module, under the operation $(a_{1},a_{2})b=a_{1}ba_{2},$ by the obvious verification.
Lemma 4.8.
Suppose $\mathcal{T}_{i}\supseteq\mathcal{T}$ are commutative monoids. Then $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})$ is a $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$-module, under the action $(a_{1}\otimes a_{2})b=a_{1}ba_{2}.$
Proof.
We have the $\mathcal{T}_{1}\times_{\mathcal{T}}\mathcal{T}_{2}$ action, which induces a $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$-action since $(a_{1}a\otimes a_{2})b=a_{1}aba_{2}=a_{1}baa_{2}=(a_{1}\otimes aa_{2})b.$
Corollary 4.9.
If $\mathcal{A}_{i}$ are $\mathcal{T}_{i}$-semialgebras with each $\mathcal{T}_{i}\subseteq\mathcal{T}$ and if $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ is a semialgebra as in Corollary 4.17(vi), then $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ is a $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$-semialgebra.
4.1.2. Pre-orders on the tensor product
Note that the multiplicative structure in the tensor product could be carried out along the lines of [4], but coping with addition makes the situation much more intricate.
Lemma 4.10.
If $\mathcal{M}_{i}$ each have a pre-order $\preceq$, then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ has the pre-order given by $x\preceq y$ if whenever we can write $x=\sum v_{i}\otimes w_{i}$ then we can write $y=\sum v^{\prime}_{i}\otimes w^{\prime}_{i}$ such that $v_{i}\preceq v^{\prime}_{i}$ and $w_{i}\preceq w_{i}^{\prime}$ for each $i.$
Proof.
Clearly this condition is transitive and passes to sums. ∎
Remark 4.11.
In general, for semialgebra pairs, one conceivably might have $\mathbb{1}\otimes\mathbb{1}\in(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}.$ In such a case the tensor product of surpassing relations cannot be a surpassing relation. But one must realize that even in the classical situation we could have $\mathbb{1}\otimes\mathbb{1}=\mathbb{0}.$
4.2. Properties of tensor products
Definition 4.12.
  1. (i)
    A 1-balanced map is a function $\psi:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{M}$ where $\mathcal{M}$ is a semigroup, satisfying
    $\psi(v_{1}a,v_{2})=\psi(v_{1},av_{2}),$ (4.6)
    for all $v_{i}\in\mathcal{M}_{i},$ $a\in\mathcal{T}.$
  2. (ii)
    A balanced map is a 1-balanced map satisfying
    $\psi(v_{1}+w_{1},x_{2})=\psi(v_{1},x_{2})+\psi(w_{1},x_{2}),$ (4.7)
    $\psi(x_{1},v_{2}+w_{2})=\psi(x_{1},v_{2})+\psi(x_{1},w_{2}),$ (4.8)
    for all $v_{i},w_{i},x_{i}\in\mathcal{M}_{i}.$
When $\mathcal{M}_{1},\mathcal{M}_{2}$ have negation maps, then a balanced map is required to satisfy $((-)v)\otimes w=v\otimes(-w),$ which we define to be $(-)(v\otimes w)$.
Proposition 4.13.
Suppose $\mathcal{M}_{1},\mathcal{N}_{1}$ are $(\mathcal{T}_{1},\mathcal{T})$-bimodules and $\mathcal{M}_{2},\mathcal{N}_{2}$ are $(\mathcal{T},\mathcal{T}_{2})$-bimodules. If $f_{1}:\mathcal{M}_{1}\to\mathcal{N}_{1}$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule homomorphism and $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule homomorphism, then there is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule homomorphism
$f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes_{\mathcal{T}}\mathcal{N}_{2}$
given by $(f_{1}\otimes f_{2})(v\otimes w)=f_{1}(v)\otimes f_{2}(w).$
Proof.
The balanced map $\Psi:(v,w)\mapsto f_{1}(v)\otimes f_{2}(w)$ sends $\Phi_{1}$ to $\Phi_{2},$ seen by applying $f_{1}$ and $f_{2}$ to the respective components. ∎
Corollary 4.14.
Assume throughout that $\mathcal{M}_{1}$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule and $\mathcal{M}_{2}$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule.
  1. (i)
    If $\mathcal{M}_{1}$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule, then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is a $\mathcal{T}_{1}$-module, with the $\mathcal{M}_{1}$-action $a(v\otimes w)=av\,\otimes w$. Analogously for $\mathcal{M}_{2}.$
  2. (ii)
    If $\mathcal{M}_{1}$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule and $\mathcal{M}_{2}$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule, then
    1. (a)
      $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule, with the actions $a_{1}(v\otimes w)a_{2}=a_{1}v\,\otimes\,wa_{2}$.
    2. (b)
      Any two bimodule homomorphisms $f_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}$ factor through a bimodule homomorphism $f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes\mathcal{N}_{2}.$ If $\mathcal{A}_{1},\mathcal{A}_{2}$ are weakly admissible with $\mathbb{1}\in\mathcal{T}$, then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ are weakly admissible over $\mathbb{1}\otimes\mathcal{T}=\mathcal{T}\otimes\mathbb{1}$.
  3. (iii)
    Any homomorphism $f:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}$ is determined by its action on simple tensors.
  4. (iv)
    If $\mathcal{T}$ is central in both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, then $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}\cong\mathcal{M}_{2}\otimes_{\mathcal{T}}\mathcal{M}_{1}$ via $(v_{1}\otimes v_{2})\mapsto v_{2}\otimes v_{1}.$
  5. (v)
    If both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are $\mathcal{T}$- semialgebras, then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is a semialgebra with multiplication given by $(v_{1}\otimes v_{2})(w_{1}\otimes w_{2})=(v_{1}w_{1}\otimes v_{2}w_{2}).$
Proof.
Repeated applications of Proposition 4.13.
(i) Take $f_{1}$ to be the identity, and $f_{2}$ to be right multiplication by $a,$ and symmetrically.
(ii) Take $f_{1}$ to be left multiplication by $a_{2}$, and $f_{2}$ to be right multiplication by $a_{2}$.
(iii) $f$ is determined by the balanced map $\psi:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}$ given by $\psi(v,w)=f(v\otimes w).$
(iv) Take the balanced map $(v_{1},v_{2})\mapsto(v_{2}\otimes v_{1})$.
(v) First define multiplication by $w_{1}\otimes\mathbb{1}$ via the balanced map $(v_{1},v_{2})\mapsto(v_{1}\otimes w_{1}v_{2}),$ and then left multiplication by $w_{2}\otimes\mathbb{1}$. ∎
Remark 4.15.
  1. (i)
    If one takes $\mathcal{T}=\{\mathbb{1}\}$ as in Remark 3.2(iii), then Equation (4.4) is redundant.
  2. (ii)
    Associativity of addition in $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ follows from associativity in $\mathcal{F}_{1}\times\mathcal{F}_{2}.$
Corollary 4.16.
The tensor product in this case satisfies the criteria of [21, Definitions 1,2].
Proof.
By Corollary 4.14. ∎
Corollary 4.17.
Suppose $\mathcal{M},\mathcal{M}_{i}$ are $\mathcal{T}_{1},\mathcal{T}$-bimodules and $\mathcal{N},\mathcal{N}_{j}$ are $\mathcal{T},\mathcal{T}_{2}$-bimodules.
  1. (i)
    $\mathcal{T}_{1}\otimes_{\mathcal{T}_{1}}\mathcal{M}\cong\mathcal{M}$ and $\mathcal{M}\otimes_{\mathcal{T}}{\mathcal{T}}\cong\mathcal{M}$,
  2. (ii)
    $(\oplus\mathcal{M}_{i})\otimes N\cong\oplus(\mathcal{M}_{i}\otimes N)$,
  3. (iii)
    $\mathcal{M}\otimes(\oplus\mathcal{N}_{i})\cong\oplus(M\otimes\mathcal{N}_{i})$,
  4. (iv)
    $(\oplus\mathcal{M}_{i})\otimes(\oplus\mathcal{N}_{j})\cong\oplus(\mathcal{M}_{i}\otimes\mathcal{N}_{j})$, as $\mathcal{T}_{1},\mathcal{T}_{2}$-bimodules.
  5. (v)
    If $\mathcal{N}$ is free over $\mathcal{T}$ with base $\{b_{i}:i\in I\}$ then every element of $\mathcal{M}\otimes\mathcal{N}$ can be written in the form $\sum_{i\in I}v_{i}\otimes b_{i}$, for uniquely determined $v_{i}\in\mathcal{M}$.
  6. (vi)
    If $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are semialgebras over $\mathcal{T}$ with $\mathcal{A}_{2}$ free, then $\mathcal{A}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{2}$ is a semialgebra. (Likewise if $\mathcal{A}_{1}$ is free.)
Proof.
(i) –(iv) These are results in [9, 21].
(v) The projection $\pi_{j}:\mathcal{N}\to\mathcal{N}$ onto the $j$ coordinate yields a projection $1\otimes\pi_{j}:\mathcal{M}\otimes\mathcal{N}\to\mathcal{M}\otimes\mathcal{N}$. If $\sum_{i\in I}v_{i}\otimes b_{i}=\sum_{i\in I}v^{\prime}_{i}\otimes b_{i}$ then applying $1\otimes\pi_{j}$ yields $v_{j}\otimes b_{j}=v^{\prime}_{j}\otimes b_{j}.$
(vi) Write tensors uniquely in terms of a base, by $\sum_{i\in I}v_{i}\otimes b_{i}$ as in (v), and then work with structure constants, writing $b_{i}b_{j}=\sum a_{ijk}b_{k}$, and putting $(\sum_{i\in I}v_{i}\otimes b_{i})(\sum_{j\in I}v_{j}\otimes b_{j})=(\sum_{i\in I}a_{ijk}v_{i}v_{j}\otimes b_{k})$. ∎
Remark 4.18.
We do not see how to remove the freeness restriction in (vi). We could follow [23, Theorem 18.21], to get $\mathcal{T}$-module isomorphisms
$(\mathcal{A}_{1}\otimes\mathcal{A}_{2})\otimes(\mathcal{A}_{1}\otimes\mathcal{A}_{2})\to(\mathcal{A}_{1}\otimes\mathcal{A}_{1})\otimes(\mathcal{A}_{2}\otimes\mathcal{A}_{2}),$
but now we encounter the difficulty that $\mu:\mathcal{A}_{1}\to\mathcal{A}_{1}$ might not induce a map $\bar{\mu}:\mathcal{A}_{1}\otimes\mathcal{A}_{1}\to\mathcal{A}_{1},$ since we don’t know that $(b_{1}+b_{2})\otimes(b_{1}+b_{2})=\sum_{i,j=1}^{2}b_{i}\otimes b_{j}$ and even if it does, what is $\bar{\mu}\otimes\bar{\mu}^{\prime}$?
Corollary 4.19.
Suppose $\mathcal{M}_{i}$ are $\mathcal{T}_{i},\mathcal{T}_{i+1}$-bimodules. Then $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})\otimes\mathcal{M}_{3}\cong\mathcal{M}_{1}\otimes(\mathcal{M}_{2}\otimes\mathcal{M}_{3})$ as $\mathcal{T}_{1},\mathcal{T}_{3}$-bimodules.
Proof.
The proof is from [21, Theorem 2]. ∎
Proposition 4.20.
If $\mathcal{M}_{i}=\mathcal{P}(\mathcal{H}_{i})$ where $\mathcal{M}_{1}$ is a right $\mathcal{T}$-module and $\mathcal{M}_{2}$ is a left $\mathcal{T}$-module. then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is $\subseteq$-distributive.
Proof.
We shall show that $S\otimes(S_{1}^{\prime}+S_{2}^{\prime})\subseteq SS_{1}^{\prime}\,+\,SS_{2}^{\prime}$ for all $S\subseteq\mathcal{H}_{1}$ and $S_{1}^{\prime},S_{2}^{\prime}\subseteq\mathcal{H}_{2}$. By definition, $S\otimes(S_{1}^{\prime}+S_{2}^{\prime})=\{a\otimes a^{\prime}:a\in S,\ a^{\prime}\in S_{1}^{\prime}+S_{2}^{\prime}\}=\{a(S_{1}^{\prime}+S_{2}^{\prime}):a\in S\}.$ Thus we need to show that $a\otimes(S_{1}^{\prime}+S_{2}^{\prime})\subseteq a\otimes S_{1}^{\prime}\,+\,a\otimes S_{2}^{\prime}$ for each $a\in S.$ But for each $a^{\prime}\in S_{1}^{\prime}+S_{2}^{\prime}$ if $a^{\prime}\in b_{1}^{\prime}+b_{2}^{\prime}$ for suitable nonzero $b_{i}^{\prime}\in S_{2}^{\prime}$, then $a\otimes a^{\prime}=(a\otimes b_{1}^{\prime})+(a\otimes b_{2}^{\prime}),$ and we are done by induction on height. Hence we may assume that $a^{\prime}\in\mathcal{H}_{2},$ and the assertion is immediate. ∎
4.3. Functorial properties of residue tensors
The notation for hyperaddition in residue hypermodules continues to be $\boxplus$.⊞ .
Lemma 4.21.
Suppose that $\mathcal{T}_{1},\mathcal{T}_{2}$ are abelian monoids, $\mathcal{M}_{i}$ are $\mathcal{T}_{i}$-modules for $i=1,2,$ and $G$ is a common subgroup of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}.$ The tensor product of the quotient hypermodules $\mathcal{H}_{1}=\mathcal{M}_{1}/G_{1}$ and $\mathcal{M}_{2}=\mathcal{M}_{2}/G_{2}$ is isomorphic to $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})/(G_{1}\otimes G_{2})$.
Proof.
The natural map $\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to(\mathcal{M}_{1}/G_{1})\otimes(\mathcal{M}_{2}/G_{2})$ given by $y_{1}\otimes y_{2}\mapsto y_{1}G_{1}\otimes y_{2}G_{2}$ preserves the defining congruence, thereby inducing a map of the cosets, which is 1:1 and onto. ∎
4.4. Tensor extensions of a pair
There are two flavors to extending the base monoid $\mathcal{T}$.
Definition 4.22.
Suppose $\mathcal{M}$ is a $\mathcal{T}$-module, and $\mathcal{T}^{\prime}$ is a monoid containing $\mathcal{T}$.
  1. (i)
    Define the $\mathcal{T}$-tensor extension $\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{M}$ to be the semigroup $(\mathcal{\mathcal{T}}^{\prime}\times F(\mathcal{M})/\Phi,$ where $\Phi$ is the congruence generated by all
    $\bigg{(}\big{(}a^{\prime},v+w\big{)},\bigg{(}(a^{\prime}a,w),(a^{\prime},aw)\bigg{)}$ (4.9)
    $\forall v,w\in\mathcal{\mathcal{M}}_{i},$ $a^{\prime}\in\mathcal{T}$.
  2. (ii)
    When $\mathcal{M}$ is left (weakly) admissible, define the left (weakly) admissible $\mathcal{T}$-tensor extension $\mathcal{T}^{\prime}\otimes_{\mathcal{T};\operatorname{adm}}\mathcal{M}$ to be the semigroup $(\mathcal{F}(\mathcal{T}^{\prime}\times\mathcal{M})/\Phi,$ where $\Phi$ is the congruence generated by all
    $\bigg{(}\big{(}a^{\prime},v+w\big{)},(a^{\prime},v)+(a^{\prime},w)\bigg{)},\qquad\bigg{(}(a^{\prime}a,w),(a^{\prime},aw)\bigg{)},\qquad\forall v,w\in\mathcal{\mathcal{M}}_{i},\ a^{\prime}\in\mathcal{T}.$ (4.10)
  3. (iii)
    When $\mathcal{M}$ has a negation map, we incorporate the negation map into the definition of tensor product, by enlarging $\Phi$ to include also $(\big{(}(-)a,w),\big{(}a,(-)w)\big{)}$ for $a\in\mathcal{T},$ $w\in\mathcal{F}_{i}.$
Remark 4.23.
One sees easily that if $\mathcal{M}$ is left admissible (resp. left weakly admissible), then so is $\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{M}$.
Now we modify §4.2, starting with Definition 4.12.
Definition 4.24.
  1. (i)
    A balanced map is a function $\Psi:\mathcal{T}\times\mathcal{M}\to\mathcal{N}$ where $\mathcal{N}$ is a semigroup, satisfying
    $\Psi(a,w_{1}+w_{2})=\Psi(a,w_{1})+\Psi(a,w_{2}),$ (4.11)
    $\Psi(v_{1}a,w_{1})=\Psi(v_{1},aw_{1}),$ (4.12)
    for all $a\in\mathcal{T},$ $v_{i},w_{i}\in\mathcal{M}_{i}.$
Lemma 4.25.
$\mathcal{M}^{\prime}:=\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{M}$ is a $\mathcal{T}^{\prime}$-module.
Proof.
The map $\mathcal{T}^{\prime}\times\mathcal{M}^{\prime}\to\mathcal{M}^{\prime}$ given by $(a,\sum a_{i}\otimes y_{i})\mapsto\sum aa_{i}\otimes y_{i}$ factors through the congruence $\Phi.$
As in Proposition 4.13, we have the key observation leading to the isomorphism theorems.
Theorem 4.26.
  1. (i)
    Suppose $\mathcal{T}_{1},\mathcal{T}_{2}$ are monoids containing $\mathcal{T}$, and $\mathcal{M}_{i}$, are $\mathcal{T}$-modules for $i=1,2$. If $f_{1}:\mathcal{T}_{1}\to\mathcal{T}_{2}$ is a monoid homomorphism fixing $\mathcal{T}$ and $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$ is a module homomorphism, then there is a $\mathcal{T}$-module homomorphism
    $f_{1}\otimes f_{2}:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{T}_{2}\otimes_{\mathcal{T}}\mathcal{N}_{2}$
    given by $(f_{1}\otimes f_{2})(a\otimes w)=f_{1}(a)\otimes f_{2}(w).$
  2. (ii)
    If $\mathcal{M}$ is a $\mathcal{T}$-semialgebra and $\mathcal{T}_{1}$ is commutative, then $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}$ is a $\mathcal{T}_{1}$-semialgebra.
  3. (iii)
    Suppose $\mathcal{T}_{1},\mathcal{T}_{2}$ are monoids containing $\mathcal{T}$, and $\mathcal{A}_{1},\mathcal{A}_{2}$ are $\mathcal{T}$-semialgebras. If $f_{1}:\mathcal{T}_{1}\to\mathcal{T}_{2}$ is a monoid homomorphism and $f_{2}:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a semialgebra homomorphism, then there is a $\mathcal{T}_{1}$-semialgebra homomorphism
    $f_{1}\otimes f_{2}:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{1}\to\mathcal{T}_{2}\otimes_{\mathcal{T}}\mathcal{N}_{2}$
    given by $(f_{1}\otimes f_{2})(a\otimes b)=f_{1}(a)\otimes f_{2}(b),$ for $a\in\mathcal{T}_{1},\ b\in\mathcal{A}_{1}$
Proof.
(i) The balanced map $\Psi:(a,w)\mapsto f_{1}(a)\otimes f_{2}(w)$ sends the congruence $\Phi_{1}$ of Definition 4.3 to $\Phi_{2},$ seen by applying $f_{1}$ and $f_{2}$ to the respective components.
(ii) Consider the balanced map $f_{a_{2};w_{2}}\mathcal{T}_{1}\times\mathcal{M}\to\mathcal{T}_{1}\otimes\mathcal{M}$ given by $(a,w)\mapsto(aa_{2})\otimes(aw_{2})$. This induces a map $\mathcal{T}_{1}\otimes\mathcal{M}\to\mathcal{T}_{1}\otimes\mathcal{M}$.
(iii) $f_{1}\otimes f_{2}$ respects multiplication, namely
$(f_{1}\otimes f_{2})((a_{1}\otimes w_{1})(a_{2}\otimes w_{2}))=((f_{1}\otimes f_{2})(a_{1}\otimes w_{1}))((f_{1}\otimes f_{2})(a_{2}\otimes w_{2})).$
The most important case is when $\mathcal{T}_{1}=\mathcal{T}$ and $f_{1}:\mathcal{T}\to\mathcal{T}_{2}$ is the inclusion map.
Corollary 4.27.
Suppose $\mathcal{T}_{2}$ is a monoid containing $\mathcal{T}$, and $\mathcal{M}_{1},\mathcal{M}_{2}$ are $\mathcal{T}$-modules. If $f:\mathcal{M}_{1}\to\mathcal{M}_{2}$ is a bimodule homomorphism, then there is a bimodule homomorphism
$1\otimes f:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{1}\to\mathcal{T}_{2}\otimes_{\mathcal{T}}\mathcal{M}_{2}$
given by $(1\otimes f)(a\otimes w)=a\otimes f(w).$
Corollary 4.28.
Suppose that $\mathcal{A}$ is a $\mathcal{T}$-module.
  1. (i)
    If $\mathcal{T}\subseteq\mathcal{T}_{1}\subseteq\mathcal{T}_{2}$ then $\mathcal{T}_{2}\otimes(\mathcal{T}_{1}\otimes\mathcal{A})\cong\mathcal{T}_{2}\otimes\mathcal{A}.$
  2. (ii)
    $\mathcal{T}\otimes\mathcal{A}\cong\mathcal{A}$ as semialgebras.
  3. (iii)
    If $\mathcal{A}_{2}\to\mathcal{A}_{1}$ is an epimorphism of $\mathcal{T}$-semialgebras, then $\mathcal{T}_{1}\otimes\mathcal{A}_{2}\to\mathcal{T}_{1}\otimes\mathcal{A}_{1}$ is an epimorphism.
Proof.
As in [21, Theorem 3]. ∎
5. Tensor products of morphisms which need not be homomorphisms
The proof of Proposition 4.13 fails for morphisms which are not homomorphisms, and thus its usefulness for hypermagmas is limited. Our goal nevertheless is to salvage whatever information that we can. We do not even have a functor for residue hypermagmas.
Remark 5.1.
Suppose $f_{i}\mathcal{M}_{i}\to\mathcal{N}_{i}$ are homomorphisms of $\mathcal{T}_{i}$-modules, for $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ abelian, and $G$ is a common subgroup of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}.$ Then writing $\overline{\mathcal{M}_{i}}=\mathcal{M}_{i}/G$, $\overline{\mathcal{N}_{i}}=\mathcal{N}_{i}/G,$ and $\bar{f}_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}/G$ for the weak morphism induced by $f_{i}$, we would want a balanced map $\Psi:=\bar{f}_{1}\times\bar{f}_{2}:{\mathcal{M}_{1}}\times{\mathcal{M}_{2}}\to\overline{\mathcal{N}_{1}}\otimes\overline{\mathcal{N}_{2}}$ given by
$(y_{1}\times y_{2})\mapsto f_{1}(y_{1})Gf_{2}(y_{2})=\{f_{1}(y_{1})gf_{2}(y_{2}):g\in G\},$
to induce a weak morphism $\overline{\mathcal{M}_{1}}\otimes\overline{\mathcal{M}_{2}}\to\overline{\mathcal{N}_{1}}\otimes\overline{\mathcal{N}_{2}}$.
But $\Psi(v_{1}+w_{1},x_{2})=f_{1}(v_{1}+w_{1})Gf_{2}(x_{2})$ whereas $\Psi(v_{1},x_{2})\boxplus\Psi(w_{1},x_{2})=f_{1}(v_{1})Gf_{2}(x_{2})\boxplus f_{1}(w_{1})Gf_{2}(x_{2}),$ which could be larger.
Remark 5.2.
In [20], to enable a categorical approach for hypermagmas, the tensor product of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ is defined as satisfying
$(v_{1}\otimes v_{2})\boxplus(v_{1}\otimes v_{2})=(v_{1}\boxplus v_{1})\otimes v_{2}\ \cup\ v_{1}\otimes(v_{2}\boxplus v_{2}),$ (5.1)
$(v_{1}\otimes v_{2})\boxplus(v_{1}^{\prime}\otimes v_{2})=(v_{1}\boxplus v_{1}^{\prime})\otimes v_{2},$ (5.2)
$(v_{1}\otimes v_{2})\boxplus(v_{1}\otimes v_{2}^{\prime})=v_{1}\otimes(v_{2}\boxplus v_{2}^{\prime}),$ (5.3)
$(v_{1}\otimes v_{2})\boxplus(v_{1}^{\prime}\otimes v_{2}^{\prime})=\emptyset,$ (5.4)
for $v_{1}\neq v_{1}^{\prime}\in\mathcal{M}_{1}$ and $v_{2}\neq v_{2}^{\prime}\in\mathcal{M}_{2}$. As they point out, this is the smallest hypermagma satisfying $(v_{1}\boxplus v_{1}^{\prime})\otimes v_{2}\subseteq(v_{1}\otimes v_{2})\boxplus(v_{1}^{\prime}\otimes v_{2})$ and $v_{1}\otimes(v_{2}\boxplus v_{2}^{\prime})\subseteq(v_{1}\otimes v_{2})\boxplus(v_{1}\otimes v_{2}^{\prime})$ for all $v_{1},v_{1}^{\prime}\in\mathcal{M}_{1}$ and $v_{2},v_{2}^{\prime}\in\mathcal{M}_{2}$, and has excellent functorial properties, but there is considerable collapsing, and associativity of addition even fails in the bipotent situation:
Example 5.3.
If $v_{1}+v_{2}=v_{1}$ and $w_{1}+w_{2}=w_{1}$ then $((v_{1}\otimes w_{1})+(v_{1}\otimes w_{2}))+((v_{2}\otimes w_{1})+(v_{2}\otimes w_{2}))=(v_{1}\otimes w_{1})+(v_{2}\otimes w_{1})=(v_{1}\otimes w_{1})$ whereas
$(v_{1}\otimes w_{1})+(((v_{1}\otimes w_{2}))+(v_{2}\otimes w_{1}))+(v_{2}\otimes w_{2}))=(v_{1}\otimes w_{1})+(\emptyset+(v_{2}\otimes w_{2}))=\emptyset.$
Thus, although [20] is the appropriate approach for $\preceq$-morphisms in the category of hypermagmas, we feel that the more traditional approach may be needed in the category of hypermodules.
5.1. Defining $f_{1}\otimes f_{2}$ via simple tensors
Given morphisms $f_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}$ for $i=1,2,$ we search for a well-defined morphism $f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes\mathcal{N}_{2}$.
Unfortunately, there does not seem to be a single approach that works for everything.
Example 5.4.
  1. (i)
    One might be tempted to define $\tilde{f}=f_{1}\otimes f_{2}$ by $\tilde{f}(v\otimes w)=f_{1}(v)\otimes f_{2}(w)$ and $f_{1}\otimes f_{2}$ to be zero on non-simple tensors. However, this would lead to the situation where $\tilde{f}(\lambda_{1}\otimes\lambda_{2}+\lambda_{2}\otimes\lambda_{1})=\tilde{f}(\lambda_{1}\otimes\lambda_{1}+\lambda_{2}\otimes\lambda_{2})=\mathbb{0}$ whereas $\tilde{f}((\lambda_{1}+\lambda_{2})\otimes(\lambda_{1}+\lambda_{2}))=f(\lambda_{1}+\lambda_{2})\otimes g(\lambda_{1}+\lambda_{2}).$
  2. (ii)
    Even worse, the decomposition of a tensor into the sum of two simple tensors need not be unique. For example
    $\lambda_{1}\otimes\lambda_{2}+(\lambda_{1}+\lambda_{2})\otimes\lambda_{3}=\lambda_{1}\otimes(\lambda_{2}+\lambda_{3})+\lambda_{2}\otimes\lambda_{3}.$
Here is the part that we can salvage in general.
Proposition 5.5.
Suppose $\mathcal{M}_{1},\mathcal{N}_{1}$ are $\mathcal{T}$-modules and $\mathcal{M}_{2},\mathcal{N}_{2}$ are left $\mathcal{T}$-modules. If $f_{1}:\mathcal{M}_{1}\to\mathcal{N}_{1}$ is a right multiplicative map and $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$ is a left multiplicative map, then there is a 1-balanced map denoted as $f_{1}\times_{\circ}f_{2}:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes\mathcal{N}_{2}$ given by $(b_{1},b_{2})\mapsto f_{1}(b_{1})\otimes f_{2}(b_{2}).$
Proof.
$(f_{1}\times_{\circ}f_{2})(b_{1}a,b_{2})=f_{1}(b_{1})a\otimes f_{2}(b_{2})=f_{1}(b_{1})\otimes af_{2}(b_{2})=(f_{1}\times_{\circ}f_{2})(b_{1},ab_{2}).$
Remark 5.6.
If $\mathcal{M}_{1},\mathcal{N}_{1}$ are $(\mathcal{T}_{1},\mathcal{T})$-bimodules and $\mathcal{M}_{2},\mathcal{N}_{2}$ are $(\mathcal{T},\mathcal{T}_{2})$-bimodules, then $\mathcal{M}_{1}\times\mathcal{M}_{2}$ is a $\mathcal{T}_{1},\mathcal{T}_{2}$-bimodule, where the $\mathcal{T}_{1}$ action is on the left on $\mathcal{M}_{1},$ and the $\mathcal{T}_{2}$ action is on the left on $\mathcal{M}_{2}.$ Then $f_{1}\times_{\circ}f_{2}$ of Proposition 5.5 preserves these left and right actions.
Given a $(\mathcal{T}_{1},\mathcal{T})$-bimodule prepair $({\mathcal{M}_{1}},{\mathcal{M}_{1}}_{0})$, a $\mathcal{T},\mathcal{T}_{2}$-bimodule prepair $({\mathcal{N}},{\mathcal{N}}_{0})$, and a $\mathcal{T}_{1},\mathcal{T}_{2}$-bimodule prepair $(\mathcal{N},\mathcal{N}_{0})$, define $\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{N})$ to be the set of $\mathcal{T}_{1},\mathcal{T}_{2}$ multiplicative maps $f:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}$ (as in Remark 5.6) satisfying the property that if $b_{j},b_{i,j}\in\mathcal{M}_{i}$ with $\sum_{j}b_{i,j}\in{\mathcal{M}_{i}}_{0},$ then
$f(\sum_{j}b_{1,j},b_{2}),f(b_{1},\sum_{j}b_{2,j})\in\mathcal{N}_{0}.$
Lemma 5.7.
If $f_{i}:(\mathcal{M}_{i},{\mathcal{M}_{i}}_{0})\to(\mathcal{N}_{i},{\mathcal{N}_{i}}_{0})$ are weak morphisms, then
$f_{1}\times_{\circ}f_{2}\in\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{N}_{1}\otimes\mathcal{N}_{2}).$
Proof.
$(f_{1}\times_{\circ}f_{2})(\sum_{j}b_{1,j},b_{2})=f_{1}(\sum_{j}b_{1,j})\otimes f_{2}(b_{2})\in{(\mathcal{N}_{1}\otimes\mathcal{N}_{2}})_{0}.$ Likewise for the other side. ∎
5.1.1. The tensor product of a weak morphism and a homomorphism when $\mathcal{M}_{2}$ is free
Proposition 5.5 raises the issue of how to define $(f\otimes g)(v\otimes w)$, with the hope of extending it to all of $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$.
When $\mathcal{M}_{2}$ is free as a $\mathcal{T}$-module, we can still define the tensor product of a weak morphism $f_{1}$ and a homomorphism $f_{2}$. Then we can write tensors uniquely in terms of a base, by taking an arbitrary element $\sum_{i\in I}v_{i}\otimes b_{i}$ as in Corollary 4.17(v), and put $(f\otimes g)(\sum_{i\in I}v_{i}\otimes b_{i})=f(v_{i})\otimes g(b_{i})$. We can phrase this as:
Proposition 5.8.
Notation as in Proposition 4.13, suppose $\mathcal{M}_{2}$ is free, with base $\{b_{i}:i\in I\}$. Then for any weak $(\mathcal{T}_{1},\mathcal{T})$-bimodule morphism (resp.  $\preceq$-morphism) $f_{1}:\mathcal{M}_{1}\to\mathcal{N}_{1}$ and any $(\mathcal{T},\mathcal{T}_{2})$-bimodule homomorphism $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$, there is a weak $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule morphism (resp.  $\preceq$-morphism)
$f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes_{\mathcal{T}}\mathcal{N}_{2}$
given by $(f_{1}\otimes f_{2})(\sum_{j}v_{j}\otimes b_{j})=\sum_{j}f_{1}(v_{j})\otimes f_{2}(b_{j}).$
Proof.
We do it for $\preceq$-morphisms. The proof for weak morphisms is parallel. Define the map $f_{1}\times_{\circ}f_{2}$ as in Proposition 5.5. Then $f_{1}\otimes f_{2}$ is well-defined when we write vectors of $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ uniquely in the form $\sum v_{i,j}\otimes b_{j}$, and
$(f_{1}\otimes f_{2})$ $(\sum v_{j}\otimes b_{j}+\sum v_{j}^{\prime}\otimes b_{j})=(f_{1}\otimes f_{2})\left((\sum v_{j}+\sum v_{j}^{\prime})\otimes b_{j}\right)$ (5.5)
$=f_{1}(\sum v_{j}+\sum v_{j}^{\prime})\otimes f_{2}(b_{j})\preceq\sum_{j}f_{1}(v_{j})\otimes f_{2}(b_{j})+\sum_{j}f_{1}(v_{j}^{\prime})\otimes f_{2}(b_{j})$
$=(f_{1}\otimes f_{2})(\sum v_{j}\otimes b_{j})+(f_{1}\otimes f_{2})(\sum v_{j}^{\prime}\otimes b_{j}).$
Remark 5.9.
Recall from [1, Theorem 5.36] that any hypermodule must satisfy the condition that any set of subsets of $\mathcal{P}(\mathcal{H})$ has an infinum, namely the intersection.
5.1.2. The tensor product of two weak morphisms or two $\preceq$-morphisms when $\mathcal{M}_{2}$ is free
In this case we can also modify the idea of [20], introducing the empty set and partial maps.
Proposition 5.10.
If $\mathcal{M}_{2}$ is free and $f_{1}$ and $f_{2}$ both are weak morphisms or $\preceq$-morphisms, we can define
$(f_{1}\otimes f_{2})(\sum_{i=1}^{m}v_{i}\otimes b_{i})=\begin{cases}f_{1}(v_{1})\otimes f_{2}(b_{1}),\qquad t=1,\\ \emptyset,\qquad t>1.\end{cases}$ (5.6)
Proof.
We do it for $\preceq$. The only instance where $f_{1}\otimes f_{2}$ is defined on a sum is for $v_{1}\otimes b+v_{1}^{\prime}\otimes b=(v_{1}+v_{1}^{\prime})\otimes b,$ where
$(f_{1}\otimes f_{2})((v_{1}+v_{1}^{\prime})\otimes b)=f_{1}(v_{1}+v_{1}^{\prime})\otimes f_{2}(b)\preceq f_{1}(v_{1})\otimes f_{2}(b)+f_{1}(v_{1}^{\prime})\otimes f_{2}(b)=(f_{1}\otimes f_{2})(v_{1}\otimes b)+(f_{1}\otimes f_{2})(v_{1}^{\prime}\otimes b).$
5.2. $\preceq$-Tensor products of $\preceq$-morphisms
Another way to obtain single-valued functions is by introducing an extra condition on the surpassing relation $\preceq$ (which holds for hyperrings, for example). This tensor product might not satisfy the adjoint isomorphism below, but is intriguing.
Theorem 5.11.
Suppose $(\mathcal{N},\preceq)$ is a semilattice closed under infina of arbitrary sets. Then given maps $f_{i}:\mathcal{M}_{i}\to\mathcal{N}$, define $f:=f_{1}\otimes_{\preceq}f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}$ by
$f(\sum v_{i}\otimes w_{i})=\bigwedge\left\{\sum f_{1}(v_{i}^{\prime})\otimes f_{2}(w_{i}^{\prime}):\sum v_{i}^{\prime}\otimes w_{i}^{\prime}=\sum v_{i}\otimes w_{i}\right\}.$ (5.7)
Then $f$ is a $\preceq$-morphism.
Proof.
We need to show that $f(\sum v_{i,1}\otimes w_{i,1}+\sum v_{i,2}\otimes w_{i,2})\preceq f(\sum v_{i,1}\otimes w_{i,1})+f(\sum v_{i,2}\otimes w_{i,2})$. But if $\sum v_{i,1}^{\prime}\otimes w_{i^{\prime},1}=\sum v_{i,1}\otimes w_{i,1}$ and $\sum v_{i^{\prime},2}\otimes w_{i^{\prime},2}=\sum v_{i,2}\otimes w_{i,2}$ then
$\sum v_{i,1}^{\prime}\otimes w_{i^{\prime},1}+\sum v_{i^{\prime},2}\otimes w_{i^{\prime},2}=\sum v_{i,1}\otimes w_{i,1}+\sum v_{i,2}\otimes w_{i,2},$
so
$f\left(\sum v_{i,1}\otimes w_{i,1}+\sum v_{i,2}\otimes w_{i,2}\right)$ $\preceq\bigwedge\left(\sum f_{1}(v_{i,1}^{\prime})\otimes f_{2}(w_{i,1}^{\prime})\right)+\bigwedge\left(\sum f_{1}(v_{i,2}^{\prime})\otimes f_{2}\left(w_{i,2}^{\prime}\right)\right)$ (5.8)
$=f\left(\sum v_{i,1}\otimes w_{i,1}\right)+f\left(\sum v_{i,2}\otimes w_{i,2}\right),$
as desired. ∎
Theorem 5.11 can be applied to hypermagmas in view of Remark 3.5, but at cost of introducing the empty set (the $\infty$ element of $\mathcal{P}(\mathcal{H}),$ cf. footnote 3 of Definition 3.1). In other words, $f$ could be undefined on tensors which can be rewritten in inequivalent ways.
5.2.1. Digression: Hypertensor products of maps
An intriguing (but problematic) way to obtain a theory that runs smoothly is by introducing a hyper-structure.
Remark 5.12.
Given maps $f_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}$, $i=1,2,$ one could try defining $f:=f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{P}(\mathcal{N}_{1}\otimes\mathcal{N}_{2})$ by $f(\sum v_{i}\otimes w_{i})=\{\sum f_{1}(v_{i}^{\prime})\otimes f_{2}(w_{i}^{\prime}):\sum v_{i}^{\prime}\otimes w_{i}^{\prime}=\sum v_{i}\otimes w_{i}\}.$
It is easy to see that if $f_{1},f_{2}$ are $\supseteq$-morphisms on hyperpairs then $f_{1}\otimes f_{2}$ is a $\supseteq$-morphism. The difficulty with this observation is that for the hyperpairs of greatest interest (for quotient hyperfields) the natural morphisms are $\subseteq$-morphisms, for which $f_{1}\otimes f_{2}$ is not a $\subseteq$-morphism.
5.3. Tensor extensions of $\preceq$-morphisms and weak morphisms of pre-semialgebra pairs
One can push this further, defining tensor extensions of pre-semialgebras as in Definition 4.22.
Theorem 5.13.
Suppose that $\mathcal{A}$ is a weakly admissible $\mathcal{T}$-algebra, $\mathcal{T}^{\prime}$ is a monoid containing $\mathcal{T}$, and $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a pre-semialgebra homomorphism.
  1. (i)
    If $\mathcal{T}^{\prime}$ is the disjoint union $\cup_{i\in I}c_{i}\mathcal{T}$ for $c_{i}\in\mathcal{T}^{\prime}$ then $f$ extends naturally to a weak pre-semialgebra morphism (resp.  $\preceq$-morphism) $\tilde{f}:{\mathcal{T}^{\prime}}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{N}_{2}$ given by $\tilde{f}(c_{i}a_{i}\otimes y_{i})=\sum c_{i}\otimes f(a_{i}y_{i}).$
  2. (ii)
    If $\mathcal{A}_{1}$ is free over $\mathcal{T}$ with base $\{b_{i}:i\in I\},$ then $f$ extends naturally to a weak pre-semialgebra morphism (resp.  $\preceq$-morphism) $\tilde{f}:{\mathcal{T}^{\prime}}\otimes_{\mathcal{T};\operatorname{adm}}\mathcal{M}_{2}\to\mathcal{T}^{\prime}\otimes_{\mathcal{T};\operatorname{adm}}\mathcal{N}_{2}$ given by $\tilde{f}(\sum_{i}a^{\prime}_{i}\otimes b_{i})=\sum_{i}a^{\prime}_{i}\otimes f(b_{i}).$
Proof.
We do it for $\preceq$-morphisms.
[…]
$\preceq c_{i}a_{i}\otimes f(y_{i})+c_{i}a_{i}\otimes f(y^{\prime}_{i})=\tilde{f}(c_{i}a_{i}\otimes y_{i})+f(c_{i}a_{i}\otimes y_{i}^{\prime}).$
[…]
$\preceq\sum_{i}a_{i}\otimes f(b_{i})+\sum_{i}a_{i}^{\prime}\otimes f(b_{i})=$ $\tilde{f}(\sum_{i}a_{i}\otimes b_{i})+\tilde{f}(\sum_{i}a_{i}^{\prime}\otimes b_{i}),\quad\forall a_{i},a_{i}^{\prime}\in\mathcal{T}^{\prime}.$
5.4. The adjoint correspondence
Remark 5.14.
Suppose that $\mathcal{M}_{i}$ are $(\mathcal{T}_{i},\mathcal{T}_{i+1})$-bimodules. The celebrated adjoint isomorphism theorem says
$\operatorname{Hom}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})\approx\operatorname{Hom}(\mathcal{M}_{2},\operatorname{Hom}(\mathcal{M}_{1},\mathcal{M}_{3})).$
The usual proof for modules over rings, say in [22, Proposition 2.10.9] matches values on simple tensors, and then extends the $\subseteq$-morphism $f=f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}$ from its action on simple tensors. This still works for homomorphisms of modules over semirings.
But what about different sorts of morphisms, i.e., $\preceq$-morphisms or weak morphisms? First of all, we have seen for $\preceq$-morphisms that $f\otimes g$ need not be well-defined on simple tensors. Even if it is well-defined, extending it to all tensors is a challenge, which we take on in this section. First we lower our sights in general.
Theorem 5.15.
$\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{M}_{3})\approx\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3})).$
Proof.
As in the proof of [22, Proposition 2.10.9]. Given $f\in\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{M}_{3})$ and $w\in\mathcal{M}_{2},$ we define $f_{w}:\mathcal{M}_{1}\to\mathcal{M}_{3}$ by $v\mapsto f(v\otimes w).$ Then for $\sum v_{i}\in{\mathcal{M}_{1}}_{0}$,
$f_{w}(v_{1}+v_{2})=f\left(\left(\sum v_{i}\right)\otimes w\right)=f\left(\sum v_{i}\otimes w\right)\preceq\sum f\left(v_{i}\otimes w\right),$
and likewise for the other side, so $f\in\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}).$ We get $\Phi\in\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}))$ by declaring $\Phi(f)(w)=f_{w}.$
Conversely, given $g:\mathcal{M}_{2}\to\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}),$ one defines the map $\psi_{g}:(v,w)\mapsto g(w)(v)$.
These correspondences clearly are inverses. ∎
However, our maps in $\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{M}_{3})$ need not be balanced, so might not produce maps of tensor products. In case $\mathcal{M}_{2}$ is free, with base $\mathcal{B}$ over $\mathcal{T},$ we could utilize Proposition 5.10. Define ${\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\mathcal{N})$ to be the multiplicative maps which are 0 except on multiples of base elements.
Theorem 5.16.
${\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}))$ is a section of ${\operatorname{Mor}_{\preceq}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ when $\mathcal{M}_{2}$ is free over $\mathcal{T}.$
Proof.
We adapt the proof of Theorem 5.15. Given $f\in{\operatorname{Mor}_{\preceq}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ and $w=ab$ for $b\in{\mathcal{B}},$ we define $f_{w}:\mathcal{M}_{1}\to\mathcal{M}_{3}$ by $v\mapsto f(v\otimes w).$ Then $f_{w}(v+v^{\prime})=f((v+v^{\prime})\otimes w)\preceq f(v\otimes w)+f(v^{\prime}\otimes w).$ We define $\Phi\in{\operatorname{Mor}_{\preceq}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ by $\Phi(f)(w)=f_{w}.$
In the other direction, given $g\in{\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3})),$ one defines the map $\psi_{g}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{M}_{3}$ by
$\psi_{g}(\sum_{j=1}^{t}v_{j}\otimes b_{j})=\begin{cases}g(b_{1})v_{1},\qquad t=1,\\ 0,\qquad\qquad t>1.\end{cases}$
$\psi_{g}$ is a $\preceq$-morphism since $\psi_{g}((\sum_{j=1}^{t}v_{j}\otimes b_{j}))+(\sum_{j=1}^{t^{\prime}}v^{\prime}_{j}\otimes b^{\prime}_{j})=0$ unless $t=t^{\prime}=1$ and $b_{1}^{\prime}=b_{1},$ in which case $\psi_{g}((v_{1}\otimes b_{1}))+(v^{\prime}_{1}\otimes b_{1}))=\psi_{g}((v_{1}+v^{\prime}_{1})\otimes b_{1})=g(b_{1})(v_{1}+v^{\prime}_{1})\preceq g(b_{1})(v_{1})+g(b_{1})(v_{1^{\prime}})$.
Clearly $\Phi(\psi_{g}(w))(v)=0$ unless $w=ab$ for $b\in{\mathcal{B}},$ in which case
$\Phi(\psi_{g}(w))(v)={\psi_{g}}_{w}(v)=ag(b)(v),$
implying $\Phi\psi_{g}=g$. ∎
The analogous result holds for $\mathcal{M}_{1}$ free. The following general observation specializes to the usual adjoint isomorphism in module theory.
Lemma 5.17.
$\mathcal{M}_{i}$ are $(\mathcal{T}_{i},\mathcal{T}_{i+1})$-bimodules. Define ${\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ to be the $\preceq$-morphisms. Then there are canonical functors
$\Phi:{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})\to{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{2}}(\mathcal{M}_{2},{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1},\mathcal{M}_{3})),$
and
$\Psi:{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{2}}(\mathcal{M}_{2},{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1},\mathcal{M}_{3}))\to{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3}),$
given in the proof.
Proof.
As in the proof of [22, Proposition 2.10.9]. Given $f\in{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ and an $w\in\mathcal{M}_{2},$ we define $f_{w}:\mathcal{M}_{1}\to\mathcal{M}_{3}$ by $v\mapsto f(v\otimes w).$ Then
$f_{w}(v_{1}+v_{2})=f((v_{1}+v_{2})\otimes w)=f(v_{1}\otimes w+v_{2}\otimes w)\preceq f(v_{1}\otimes w)+f(v_{2}\otimes w),$
so $f\in{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1},\mathcal{M}_{3}).$ We get $\Phi\in{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{2}}(\mathcal{M}_{2},{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1},\mathcal{M}_{3}))$ by declaring $\Phi(f)(w)=f_{w}.$
Conversely, given $g:\mathcal{M}_{2}\to{\operatorname{Mor}_{\preceq}}(\mathcal{M}_{1},\mathcal{M}_{3}),$ one can define the map $\psi_{g}:(v,w)\mapsto g(w)(v)$, which yields a right $\preceq$-defined morphism $\bar{\psi}_{g}$ using Theorem 5.11, namely
$\bar{\psi}_{g}(\sum_{i}v_{i}\otimes w_{i})=\bigwedge\left\{\sum g(w^{\prime}_{i})(v^{\prime}_{i}):\sum v_{i}\otimes w_{i}=\sum v^{\prime}_{i}\otimes w^{\prime}_{i}\right\}.$