### v4

# Tensor products of bimodules and bimodule pairs over monoids

Abstract.
We modify the well-known tensor product of modules over a semiring, in order to treat modules over hyperrings, and, more generally, for bimodules (and semi-algebras) over monoids. The tensor product of residue hypermodules is functorial with respect to this construction. Special attention is paid to different kinds of morphisms and the work of Nakamura and Reyes.
Key words and phrases:
bimodule, hyperfield, hyperring, pair, magma, quotient hyperring, residue hypermodule, surpassing relations, tensor product, weak morphism
2020 Mathematics Subject Classification:
Primary 14T10, 16Y20, 16Y60, 20N20; Secondary 15A80; .
2020 Mathematics Subject Classification:
Primary 15A69, Secondary 15A80, 16D90, 17D99
The author was supported by the ISF grant 1994/20 and the Anshel Peffer Chair.
The author is indebted to Jaiung Jun for many enlightening conversations about the categorical aspects.
1. Introduction
This is part of an ongoing project to find a general algebraic framework that is suitable for tropical mathematics, and also to handle varied structures such as idempotent semirings, F1-geometry/algebra, hyperfields, and so forth. In the process, we bypass negation as much as feasible. This idea originally is found in [6, 7] (also see [3]), and was implemented for blueprints in [15, 16], and put in a more general context in [24]. A minimalist set of axioms, for a “pair,” was introduced in [11], and made more precise in [2], as an additive semigroup $(\mathcal{A},+)$ together a distinguished null subsemigroup $\mathcal{A}_{0}$, taking the place of a zero element, both of which are acted on by a set $\mathcal{T}$ which often is a group. To avoid further repetition, we refer to the introductions of [1] and [11] for more background. The main example in this paper comes from the class of hypermodules.
Pairs can be studied by means of universal algebra, and are closed under products, sums, and polynomial extensions, and provide a concrete tensor category with respect to homomorphisms, in the sense of [9, 21]. An early reference for tensor products is [13].
However, tensor products are not so malleable when one engages with the sorts of morphisms that arise in hyperfields, such as “weak morphisms” (Definition 2.24) and $\preceq$-morphisms” (Definition 2.18). To begin with, it is not even clear what we want to take the tensor product over: the hyperfield or its power set? Since the power set of a hyperfield need not be a semiring, the latter seems an impossible task.
Remark 1.1.
Let us start with the observation that the classical treatment of the tensor product of a right module and a left module over a ring $C$ (e.g., see [23, Chapter 18]) only uses the fact that $C$ is a monoid. So we shall talk of tensor products over a multiplicative monoid.
A fully satisfactory tensor theory for tensor products of monoids over a monoid is provided in [4], but it seems that the presence of the extra operation of addition often gets in the way. One could attempt to define the tensor product of weak morphisms via simple tensors, but even this need not be well-defined without further assumptions.
Recently, Nakayama and Reyes [20] have tackled the issue of tensor products in full generality for hypermagmas (in which the hyperaddition $\boxplus$ is replaced by a binary operation *), in an interesting paper which also provides intriguing examples of noncommutative hypergroups. Their solution was to declare that non-simple tensors are vacuous, which is needed to obtain a general categorical tensor product of hypermagmas. The downside is that the classical theory of tensor products of algebras is not a special case of their results. Furthermore, one may lose associativity of addition, as seen in Example 5.3, and basic constructions such as polynomials cannot be viewed in terms of tensor products.
Our goal in this paper is to see how far one can carry a theory that includes the familiar tensor product of modules, which is compatible with the residue (quotient) hypermodule construction originated by Krasner [14].
When using homomorphisms for our morphisms, we get the analog of the familiar module-theoretic results in Theorem 4.24 and its corollaries; also see Remark 5.17 for the adjoint correspondence.
But weak morphisms are more relevant for hypermodules We do obtain weak morphisms of tensor extensions, and for tensor products when one of the tensor components is free, such as for polynomials, and get a chunk of the adjoint isomorphism theorem in Lemma 5.20.
In order to be applicable to hyperfields, the theory must not require distributivity of multiplication over addition, since the power sets of certain hyperfields are not distributive, and then taking polynomials ruins associativity of multiplication. Thus of necessity, as in [20], for a general theory one is led to magmas, dropping associativity in multiplication, and [20] frames the theory over magmas. Nevertheless, after a brief excursion to magmas, we focus on modules over monoids.
1.1. Shape of the paper
We start by reviewing the basics of magmas, also endowed with a two-sided action by a set, and of “pairs,” possibly with a surpassing relation denoted as $\preceq$. There are three major classes of morphisms (each pertaining to its own category): homomorphisms (the morphisms in universal algebra), $\preceq$-morphisms (which preserve the surpassing relation), and weak morphisms (which preserve the connection to the null subsemigroup). Our main motivation in this paper is from “hyperpairs,” inspired by hyperfield theory (and more generally the hypermagmas of [20]), which include supertropical algebra, although many more examples of pairs can be found in [2] and [11].
Then we bring in the main notion of this paper, the tensor product. We start with a standard module-theoretic approach which is appropriate for categories involving homomorphisms. The definition of tensor product in Definition 4.2 is standard. Some care needs to be taken for defining tensor products of homomorphisms, in Theorems 4.12 and 4.24].
However, more care is needed when dealing with $\preceq$-morphisms or weak morphisms:
  • We can restrict to free modules, as done in Theorem 4.24 when one of the morphisms is a homomorphism; this permits us to extend $\preceq$-morphisms or weak morphisms on tensor extensions.
  • We may modify ideas espoused in [20], in Proposition 5.11.
  • There is a natural tensor product map for hypermodules (Theorem 5.13), but which is not resonant with category theory.
We obtain tensor products of $\preceq$-morphisms in Theorem 5.13.
As in the familiar theory of tensor products of modules and algebras, we also obtain the $\mathcal{T}$-semialgebra version of tensor extensions in Theorem 5.15.
Finally we turn to a pillar of the category theory of modules, the adjoint isomorphism theorem. This seems impossible to obtain in full, but a partial result is given in Theorem 5.18. The closest we get to an adjoint isomorphism theorem involving weak morphisms is Theorem 5.19.
2. Preliminaries
2.1. Underlying algebraic structures
$\mathbb{N}^{*}$ denotes the positive natural numbers, and we set $\mathbb{N}=\mathbb{N}^{*}\cup{\mathbb{0}}.$
Definition 2.1.
  1. (i)
    A magma is a set $\mathcal{A}$ with a binary operation $(\mathcal{A}\cup\{\infty\})\times(\mathcal{A}\cup\{\infty\})\to\mathcal{A}\cup\{\infty\}$, not necessarily associative, satisfying $\infty*b=b*\infty=\infty$ for all $b\in\mathcal{A}.$ (Thus $\infty$ is an absorbing element.) The magma is total if the operation is total, i.e., $\mathcal{A}\times\mathcal{A}\to\mathcal{A}$. [^1] In most applications the magma is total, but we shall utilize $\infty$ to describe tensor products. Given any total operation, one can formally adjoin the absorbing element $\infty$.
    In this paper a magma always has a neutral element, often denoted $\iota$, i.e., $\iota*b=b*\iota=b$ for all $b\in\mathcal{A}$. [^2] In [20] these magmas are called unital, with the neutral element designated as $e,$ but we have used $e$ for another purpose, cf. §2.2.1.
  2. (ii)
    A monoid is a magma with an associative operation, not necessarily commutative. An additive semigroup is a commutative monoid, with the operation denoted by “++,” endowed with a zero element $\mathbb{0}.$
  3. (iii)
    A pre-semiring $\mathcal{A}$ is a monoid under two operations, multiplication, denoted as concatenation, and addition. We shall denote multiplication by concatenation. We can always adjoin an element  $\mathbb{0}$ to $\mathcal{A}$ that is additively neutral and also multiplicatively absorbing, and a unity element $\mathbb{1}$ that is multiplicatively neutral, so we only consider pre-semirings that have such a $\mathbb{0}$ and $\mathbb{1}$.
  4. (iv)
    A semiring [8] is a pre-semiring that satisfies all the properties of a ring (including associativity and distributivity of multiplication over addition), but without negation.
  5. (v)
    A semiring $\mathcal{A}$ is a semifield if $(\mathcal{A}\setminus\{\mathbb{0}\},\cdot)$ is a group.
2.1.1. $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas
We recall some notions from [2], slightly more generally, taking [20] into account.
Definition 2.2.
Let $(\mathcal{T},\mathbb{1})$ be a monoid with a unit element $\mathbb{1}$.
  1. (i)
    A left $\mathcal{T}$-magma is a magma $(\mathcal{A},*,\iota)$, together with a (left) $\mathcal{T}$-action $\mathcal{T}\times\mathcal{A}\to\mathcal{A}$ (also denoted as concatenation), for which, for all $a,a_{i}\in\mathcal{T}$, $b,b_{i}\in\mathcal{A}$,
    1. (a)
      $\mathbb{1}b=b$.
    2. (b)
      $a_{1}(a_{2}b)=(a_{1}a_{2})b$.
    3. (c)
      $\iota$ is absorbing, i.e. $a\iota=\iota.$
    4. (d)
      The action is distributive over $\mathcal{T}$, in the sense that $a(b_{1}*b_{2})=ab_{1}*ab_{2}.$ (In particular, $a(b_{1}*b_{2})=\infty$ if and only if $ab_{1}=\infty$ or $ab_{2}=\infty.$)
    Then we adjoin an absorbing element $\iota_{\mathcal{T}}$ to $\mathcal{T}$, and denote $\mathcal{T}_{\iota}=\mathcal{T}\cup\{\iota_{\mathcal{T}}\}$, declaring $\iota_{\mathcal{T}}b=\iota$ for each $b\in\mathcal{A}$. We call the elements of $\mathcal{T}$ tangible. $\mathcal{T}$ will normally be understood from the context; in case of ambiguity, we write $\mathcal{T}_{\mathcal{A}}$ for emphasis.
  2. (ii)
    A left $\mathcal{T}$-submagma of a left $\mathcal{T}$-magma $\mathcal{A}$ is a left submagma closed under the $\mathcal{T}$-action.
  3. (iii)
    A $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $\mathcal{A}$ is a left $\mathcal{T}_{1}$-magma and a right $\mathcal{T}_{2}$-magma satisfying $(a_{1}b)a_{2}=a_{1}(ba_{2})$ for all $a_{i}\in\mathcal{T}_{i}$ and $b\in\mathcal{A}.$
  4. (iv)
    A $\mathcal{T}$-bimagma is a $(\mathcal{T},\mathcal{T})$-bimagma.
  5. (v)
    $\mathcal{T}$ is central in a $\mathcal{T}$-bimagma $\mathcal{A}$ if $ab=ba$ for all $a\in\mathcal{T}$ and $b\in\mathcal{A}.$
Remark 2.3.
  1. (i)
    The classical case is when $R$ is an algebra over a commutative ring $C$, and $\mathcal{T}=C$, viewed as a multiplicative monoid.
  2. (ii)
    More generally, $\mathcal{T}$ could be a partial monoid, i.e., with multiplication only defined on certain pairs of elements, with changes mutatis mutandis, using Defininition 2.2(i)(b) only when the right side is defined. For example, $\mathcal{T}$ could be, say, a set of matrix units with multiplication defined only on compatible matrix units.
  3. (iii)
    If a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $\mathcal{A}$ did not already contain a neutral element $\iota,$ we could adjoin it formally by declaring its operation on all $b\in\mathcal{A}$ by $\iota*b=b*\iota=b$, and $a_{i}\iota=\iota a_{i}=\iota$ for all $a\in\mathcal{T}.$
(So far these definitions are rather general, and could include for example the case where $\mathcal{A}$ is a monoid with unit element $\iota$, and $\mathcal{T}$ is its semigroup of monoid endomorphisms.)
Definition 2.4.
  1. (i)
    A left $\mathcal{T}$-magma $\mathcal{A}$ is weakly admissible if $\mathcal{T}\subseteq\mathcal{A}$.
  2. (ii)
    A weakly admissible, resp. admissible, right $\mathcal{T}$-bimagma is defined analogously to (i) and (ii).
  3. (iii)
    A weakly admissible $\mathcal{T}$-bimagma $\mathcal{A}$ is a left and right weakly admissible $\mathcal{T}$-magma (identifying both copies of $\mathcal{T}$ inside $\mathcal{A}$).
Lemma 2.5.
Let $\mathcal{A}$ be a weakly admissible $\mathcal{T}$-magma. Then $\langle\mathcal{T}\rangle$ (the submagma generated by $\mathcal{T}$) is an admissible $\mathcal{T}$-submagma of $\mathcal{A},$ which is $\mathcal{A}$ if and only if $\mathcal{A}$ is admissible.
Proof.
Define the height function $h:\mathcal{A}\to\mathbb{N}\cup\{\infty\}$ inductively, as follows:
  • Every element of $\mathcal{T}$ has height 1.
  • $h(b)=m$, if $h(b)$ is not less than $m$ and $b=b_{1}*b_{2}$ where $h(b_{i})<m$ for $i=1,2.$
We show by induction on $h(b)$ that $ab\in\langle\mathcal{T}\rangle$ for $a\in\mathcal{T}$ and $b\in\langle\mathcal{T}\rangle.$ If $a,b\in\mathcal{T}$, then $ab\in\mathcal{T}$ by definition. We continue inductively, to see that if $b=b_{1}*b_{2}$ for $b_{1},b_{2}\in\langle\mathcal{T}\rangle$, then $ab=a(b_{1}*b_{2})=ab_{1}*ab_{2}\in\langle\mathcal{T}\rangle$. ∎
2.2. Prepairs and pairs
Definition 2.6.
A $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair $(\mathcal{A},\mathcal{A}_{0})$ (sometimes written $(\mathcal{A},\mathcal{A}_{0};*)$ if we want to specify the operation *) is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $(\mathcal{A},*,\iota)$ together with a $(\mathcal{T}_{1},\mathcal{T}_{2})$-sub-bimagma $\mathcal{A}_{0}$ (i.e., $a_{1}b_{0}\in\mathcal{A}_{0}$ and $b_{0}a_{2}\in\mathcal{A}_{0}$ for all $a_{i}\in\mathcal{T}_{i}$ and $b_{0}\in\mathcal{A}_{0}).$
A $(\mathcal{T},\mathcal{T})$-prepair is also called a $\mathcal{T}$-prepair. We suppress $\mathcal{T}$ in the notation when it is understood.
Definition 2.7.
We modify [2, 11] slightly.
  1. (i)
    A pair $(\mathcal{A},\mathcal{A}_{0})$ is a weakly admissible prepair with $ab=ba$ for each $a\in\mathcal{T},$ $b\in\mathcal{A}$. [^3] In [1] $\mathcal{A}_{0}$ was not required to be closed under ()(*). In previous work [1, 2, 11] we assumed that all pairs are proper, to dismiss the degenerate case of $\mathcal{T}=\mathcal{A}_{0}=\mathcal{A}.$ But this is precisely the case treated so successfully in [10], so we permit it here.
  2. (ii)
    A pair $(\mathcal{A},\mathcal{A}_{0})$ is said to be proper if $\mathcal{A}_{0}\cap\mathcal{T}=\emptyset$.
Important Note 2.8.
Philosophically, $\mathcal{A}_{0}$ takes the place of $\mathbb{0}$ in classical mathematics. The reason is that since modules need not have negation (for example, $\mathbb{N}$), $\mathbb{0}$ has no significant role except as a place marker in linear algebra.
2.2.1. Property N [1, §3.1]
Definition 2.9.
We say that a pair $(\mathcal{A},\mathcal{A}_{0})$ satisfies weak Property N if there is an element $\mathbb{1}^{\dagger}\in\mathcal{T}$ with $\mathbb{1}*\mathbb{1}^{\dagger}=\mathbb{1}^{\dagger}*\mathbb{1}\in\mathcal{A}_{0}$. (But $\mathbb{1}^{\dagger}$ need not be unique.) In this case, fixing $\mathbb{1}^{\dagger}$, we define $e:=\mathbb{1}*\mathbb{1}^{\dagger}\in\mathcal{A}_{0}.$ Then we denote $b^{\dagger}=b\mathbb{1}^{\dagger}$, and $b^{\circ}=b*b^{\dagger}=b^{\dagger}*b$, for all $b\in\mathcal{A}.$
Note for $a\in\mathcal{T}$ that $a^{\circ}=a(\mathbb{1}*\mathbb{1}^{\dagger})=ae\in\mathcal{A}_{0}.$
Definition 2.10.
Let $(\mathcal{A},\mathcal{A}_{0})$ be a pair satisfying Property N.
  1. (i)
    A quasi-zero is an element of the form $a^{\circ},$ for $a\in\mathcal{T}.$
  2. (ii)
    $\mathcal{T}^{\circ}:=\{a^{\circ}:a\in\mathcal{T}\}\subseteq\mathcal{A}_{0}$.
  3. (iii)
    $(\mathcal{A},\mathcal{A}_{0})$ is $\circ$-distributive if $(b_{1}*b_{2})e=b_{1}e*b_{2}e$ for all $b_{i}\in\mathcal{A}.$
  4. (iv)
    $(\mathcal{A},\mathcal{A}_{0})$ is $\mathcal{A}_{0}$-idempotent if bb=bb*b=b for all $b\in\mathcal{A}_{0}.$
  5. (v)
    Define left and right actions of $\mathcal{T}\cup\mathcal{T}^{\circ}$ on $\mathcal{A},$ by defining $a^{\circ}b:=(ab)^{\circ}$ and $ba^{\circ}:=(ba)^{\circ}$ for $a\in\mathcal{T},$ $b\in\mathcal{A}$.
By [2, Lemma 3.9], the action of (v) makes $\mathcal{A}$ a $\mathcal{T}\cup\mathcal{T}^{\circ}$-bimagma.
2.2.2. Homomorphisms of pairs
We consider $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas $\mathcal{A}$ and $(\mathcal{T}_{1}^{\prime},\mathcal{T}_{2}^{\prime})$-bimagmas $\mathcal{A}^{\prime}$, and functions $f:\mathcal{A}\to\mathcal{A}^{\prime}$ together with monoid homomorphisms $f_{\mathcal{T}_{i}}:\mathcal{T}_{i}\to\mathcal{T}_{i}^{\prime}.$
Definition 2.11.
  1. (i)
    A multiplicative map is a function $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying $f(\iota)=\iota,$ $f(a_{1}b)=f_{\mathcal{T}_{1}}(a_{1})f(b)$, and $f(ba_{2})=f(b)f_{\mathcal{T}_{2}}(a_{2})$ for all $a_{i}\in\mathcal{T}_{i},\ b\in\mathcal{A}.$
  2. (ii)
    A bimagma homomorphism (also just called a homomorphism) $f:\mathcal{A}\to\mathcal{A}^{\prime}$ is a multiplicative map satisfying $f(b_{1}*b_{2})=f(b_{1})*f(b_{2}),$ $\forall b_{i}\in\mathcal{A}.$
  3. (iii)
    $\operatorname{Hom}(\mathcal{A},\mathcal{A}^{\prime})$ is the set of homomorphisms $f:\mathcal{A}\to\mathcal{A}^{\prime}$.
Remark 2.12.
In this paper $f_{\mathcal{T}_{i}}$ always is the identity map. Thus a multiplicative map satisfies $f(a_{1}b)=a_{1}f(b)$ and $f(ba_{2})=f(b)a_{2}$ for $a_{i}\in\mathcal{T}_{i}.$
Lemma 2.13.
For $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas $\mathcal{A}$ and $\mathcal{A}^{\prime}$, $\operatorname{Hom}(\mathcal{A},\mathcal{A}^{\prime})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma under the operation $(f_{1}*f_{2})(b)=f_{1}(b)*f_{2}(b)$ and the actions $af:b\mapsto af(b)$ and $fa:b\mapsto f(b)a$.
Proof.
Standard. ∎
Definition 2.14.
Suppose $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ are prepairs.
  1. (i)
    A paired map $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying $f(\mathcal{A}_{0})\subseteq\mathcal{A}_{0}^{\prime}.$
  2. (ii)
    A paired homomorphism is a paired map which is a homomorphism.
Lemma 2.15.
  1. (i)
    Given a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair $(\mathcal{A},\mathcal{A}_{0})$ and any homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ , $(\mathcal{A}^{\prime},f(\mathcal{A}_{0}))$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair, and then $f$ is a paired homomorphism $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},f(\mathcal{A}_{0}))$.
  2. (ii)
    For any $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ where $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair, there is a prepair $(\mathcal{A},\mathcal{A}_{0})$ where $\mathcal{A}_{0}=\{b\in\mathcal{A}:f(b)\in\mathcal{A}_{0}^{\prime}\},$ and then $f$ is a paired homomorphism.
Proof.
(i) The image of $\mathcal{T}$ is disjoint from the image of $\mathcal{A}_{0}$.
(ii) If $a\in\mathcal{T}$ and $b\in\mathcal{A}_{0}$ then $f(ab)=f(a)f(b)\in\mathcal{A}_{0}^{\prime},$ so $ab\in\mathcal{A}_{0}.$ Then $f(\mathcal{A}_{0})\subseteq\mathcal{A}^{\prime}_{0},$ by definition. ∎
2.2.3. Pre-orders and surpassing relations
Definition 2.16.
  1. (i)
    A pre-order on a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagma $\mathcal{A}$, denoted $\preceq$, is a set-theoretic pre-order that respects the bimagma structure, i.e., for all $b,b_{i}\in\mathcal{A}$:
    1. (a)
      $b_{1}\preceq b_{2}$ implies $a_{1}b_{1}\preceq a_{1}b_{2}$ and $b_{1}a_{2}\preceq b_{2}a_{2}$ for $a_{i}\in\mathcal{T}_{i}.$
    2. (b)
      $b_{i}\preceq b_{i}^{\prime}$ implies $b_{1}*b_{2}\preceq b_{1}^{\prime}*b_{2}^{\prime}.$
  2. (ii)
    A surpassing relation on a weakly admissible $\mathcal{T}$-bimagma $\mathcal{A}$ is a pre-order satisfying the following:
    1. (a)
      $a_{1}\preceq a_{2}$ for $a_{1},a_{2}\in\mathcal{T}_{\iota}$ implies $a_{1}=a_{2}.$ (In other words, surpassing restricts to equality on $\mathcal{T}$.)
    2. (b)
      $b\preceq\iota$ for $b\in\mathcal{A}$ implies $b=\iota$.
  3. (iii)
    A pre-order (resp. surpassing relation) on a prepair $(\mathcal{A},\mathcal{A}_{0})$, is a pre-order (resp. surpassing relation) $\preceq$ satisfying the condition $\iota\preceq c$ for all $c\in\mathcal{A}_{0}$.
Lemma 2.17.
In a prepair $(\mathcal{A},\mathcal{A}_{0})$ with a surpassing relation, $b\preceq b*c$ and $b\preceq c*b$ for all $b\in\mathcal{A}$ and $c\in\mathcal{A}_{0}$.
Proof.
$b\preceq b$ and $\iota\preceq c,$ so $b=b*\iota\preceq b*c.$
Surpassing relations were introduced in [24], and in [2] for pairs, for the purposes of linear algebra.
We insert the preorder into our categories.
Definition 2.18.
Let $\mathcal{A},\mathcal{A}^{\prime}$ be pre-ordered $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimagmas.
  1. (i)
    A $\preceq$-morphism, (analogous to “colax morphism” in [20]) is a multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying the following conditions:
    • $f(b_{1})\preceq f(b_{2}),\quad\forall b_{1}\preceq b_{2}\in\mathcal{A}.$
    • $f(b_{1}*b_{2})\preceq f(b_{1})*f(b_{2}),\quad\forall b_{1},b_{2}\in\mathcal{A}.$
    ${\operatorname{Mor}_{\preceq}}(\mathcal{A},\mathcal{A}^{\prime})$ is the set of $\preceq$-morphisms from $\mathcal{A}$ to $\mathcal{A}^{\prime}.$
  2. (ii)
    A $\succeq$-morphism, (analogous to “lax morphism” in [20]) is a multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying the following conditions:
    • $f(b_{1})\preceq f(b_{2}),\quad\forall b_{1}\preceq b_{2}\in\mathcal{A}.$
    • $f(b_{1}*b_{2})\succeq f(b_{1})*f(b_{2}),\ \forall b_{1},b_{2}\in\mathcal{A}.$
Lemma 2.19.
Given an 1:1 multiplicative map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ where $\mathcal{A}$ is a $\mathcal{T}$-bimagma and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a pair with surpassing relation $\preceq$, the pair $(\mathcal{A},\mathcal{A}_{0})$ of Lemma 2.15(ii) has a surpassing relation given by $b_{1}\preceq b_{2}$ when $f(b_{1})\preceq f(b_{2}),$ and then $f$ becomes a $\preceq$-morphism.
Proof.
$a_{1}\preceq a_{2}$ implies $f(a_{1})\preceq f(a_{2}),$ so $f(a_{1})=f(a_{2}),$ so $a_{1}=a_{2}.$
2.2.4. Negation maps
At times we can define the negation map, the mainstay of [24]. A negation map ()(-) on a $\mathcal{T}$-module $(\mathcal{A},\mathcal{A}_{0})$ is an additive automorphism ()(-) of order $\leq 2$ also defined on $\mathcal{T}$, such that
$(-)(ab)=((-)a)b=a((-)b),\qquad\forall a\in\mathcal{T},\quad b\in\mathcal{A},$
and $(-)\mathcal{A}_{0}=\mathcal{A}_{0}.$
We write $b_{1}(-)b_{2}$ for $b_{1}*((-)b_{2}).$ Thus $b^{\circ}=b(-)b,$ and $\mathcal{A}_{0}$ contains the set $\mathcal{A}^{\circ}=\{b^{\circ}:b\in\mathcal{A}\}.$ Often $\mathcal{A}_{0}=\mathcal{A}^{\circ}.$
Lemma 2.20.
The negation map satisfies $(-)b=((-)\mathbb{1})b.$
Proof.
$(-)b=(-)(\mathbb{1}b)=((-)\mathbb{1})b.$
Hence, perhaps surprisingly, if $b_{1}\preceq b_{2},$ then $(-)b_{1}\preceq(-)b_{2}$. Hence, if $b_{1}\preceq b_{2}$ then $b_{2}(-)b_{1}\succeq\iota$ and $b_{1}(-)b_{2}\succeq\iota,$ cf. [11, Lemma 2.11].
2.3. Bimodules
We pinpoint the case of greatest interest.
Definition 2.21.
  1. (i)
    A left $\mathcal{T}$-module $\mathcal{M}$ is a left $\mathcal{T}_{\iota}$-magma where * is also commutative and associative, in which case we always write ++ in place of *, and $\mathbb{0}$ instead of $\iota$.
  2. (ii)
    A left $\mathcal{T}$-module $\mathcal{M}$ is free with base $\{b_{i}:i\in I\}$, if every element of $\mathcal{M}$ can be written uniquely in the form $\sum_{i\in I}a_{i}b_{i}$ for $a_{i}\in\mathcal{T}_{\iota}.$
  3. (iii)
    A $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimodule is a $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimagma which is a left $\mathcal{T}_{1}$-module and a right $\mathcal{T}_{2}$-module.
  4. (iv)
    Definitions 2.4 and 2.7 also apply to $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodules.
Remark 2.22.
For $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodules, $\operatorname{Hom}(\mathcal{A},\mathcal{A}^{\prime})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule under the action of Remark 2.13.
Just as the category theory for (semi)algebras differs from the category theory for bimodules, we want to differentiate between $({\mathcal{T}_{1}},{\mathcal{T}_{2}})$-bimodules (resp. bimagmas) and $\mathcal{T}$-semialgebras (resp. pre-semialgebras), which we now define.
Definition 2.23.
  1. (i)
    A $\mathcal{T}$-pre-semialgebra is a $\mathcal{T}$-bimodule $(\mathcal{A},+)$, which also has a second operation which we denote as $\cdot$, with $a_{1}(b_{1}\cdot b_{2})=(a_{1}b_{1})\cdot b_{2}$ and $(b_{1}\cdot b_{2})a_{2}=b_{1}\cdot(b_{2}a_{2})$ for $a_{i}\in\mathcal{T},$ $b_{1},b_{2}\in\mathcal{A}.$ In particular, when $\mathcal{A}$ is weakly admissible, $a\cdot b=ab$ and $b\cdot a=ba$ for all $a\in\mathcal{T}$ and $b\in\mathcal{A}.$
  2. (ii)
    A $\mathcal{T}$-semialgebra is a $\mathcal{T}$-pre-semialgebra which is a semiring.
  3. (iii)
    A pre-semialgebra homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ of $\mathcal{T}$-pre-semialgebras is a homomorphism also satisfying $f(b_{1}b_{2})=f(b_{1})f(b_{2})$ for all $b_{i}\in\mathcal{A}.$
  4. (iv)
    A pre-semialgebra $\mathcal{A}$ with a surpassing relation $\preceq$ is $\preceq$-distributive if $b(b_{1}+b_{2})\preceq bb_{1}\,+\,b_{2}b_{2}$ and $(b_{1}+b_{2})b\preceq b_{1}b\,+\,b_{2}b$ for all $b,b_{i}\in\mathcal{A}.$
2.3.1. Bimodule-theoretic notions for prepairs
The following notions are difficult to notate for bimagmas (which need not be associative under *), so we introduce them only for bimodules.
2.3.2. Weak morphisms
Definition 2.24.
A weak morphism of bimodule prepairs is a paired multiplicative map $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$, satisfying $\sum b_{i}\in\mathcal{A}_{0}$ for $b_{i}\in\mathcal{A}$ implies $\sum f(b_{i})\in\mathcal{A}^{\prime}_{0}$.
$\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime})$ denotes the set of weak morphisms from $(\mathcal{A},\mathcal{A}_{0})$ to $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0}).$
Lemma 2.25.
For bimodule prepairs $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0}),$ define $\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime})_{0}$ to be the weak morphisms for which $f(\mathcal{A})\subseteq\mathcal{A}^{\prime}_{0}.$
  1. (i)
    $(\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime}),\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime})_{0})$ is a bimodule prepair.
  2. (ii)
    For pairs $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0}),$ $(\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime}),\operatorname{Mor}_{\operatorname{wk}}(\mathcal{A},\mathcal{A}^{\prime})_{0})$ is a bimodule pair, where we define $(f_{1}+f_{2})(b)=f_{1}(b)+f_{2}(b),$ $(af)(b)=f(ab)$, $(fa)(b)=f(b)a$.
Proof.
(i) is as in Lemma 2.13.
The main verification for (ii) is that the sum $f_{1}+f_{2}$ of two weak morphisms is a weak morphism. If $\sum b_{i}\in\mathcal{A}_{0},$ then $(f_{1}+f_{2})(\sum b_{i})=f_{1}(\sum b_{i})+f_{2}(\sum b_{i})\in\mathcal{A}_{0}+\mathcal{A}_{0}\subseteq\mathcal{A}_{0}.$
Lemma 2.26 (As in [1, Lemma 2.10]).
Every $\preceq$-morphism of bimodule prepairs is a weak morphism.
3. Hypermagmas and hyperpairs
We follow the treatment of Nakamura and Reyes [20], casting hypergroups into a broader context which has many more interesting examples.
Definition 3.1 ([20], generalizing [17]).
  1. (i)
    $\mathcal{P}(\mathcal{H})$ denotes the power set of a set $\mathcal{H},$ and ${\mathcal{P}}^{\sharp}(\mathcal{H})=\mathcal{P}(\mathcal{H})\setminus\emptyset.$
    A hypermagma $(\mathcal{H},*)$ is a set with a binary operation $\mathcal{H}\times\mathcal{H}\to\mathcal{P}(\mathcal{H})$, permitting the product of two elements to be the empty set [^4] Customarily one requires $a_{1}*a_{2}\in{\mathcal{P}}^{\sharp}(\mathcal{H})$ for all $a_{1},a_{2}\in\mathcal{H}.$ But one could formally adjoin $\emptyset$, so the Nakamura-Reyes definition subsumes the customary definition, and we shall need it in what follows. $\emptyset$ could be interpreted as “undefined.”, together with a left and right $\mathcal{H}$-action whose natural extension to $\mathcal{P}(\mathcal{H})$ makes $(\mathcal{P}(\mathcal{H}),*)$ an $\mathcal{H}$-bimagma when we define, for $S_{1},S_{2}\in\mathcal{P}(\mathcal{H}),$
    $S_{1}*S_{2}:=\cup_{s_{i}\in S_{i}}\ s_{1}*s_{2},\qquad\emptyset*S=S*\emptyset=\emptyset.$
    We view $\mathcal{H}\subseteq\mathcal{P}(\mathcal{H})$ by identifying $a$ with $\{a\}$.
    From now on we follow the customary use of $\boxplus$ and $\mathbb{0}$ instead of * and $\iota$, to denote that the operation is associative. We call $\boxplus$ “hyperaddition.”
  2. (ii)
    A hyperzero $\mathbb{0}$ (if it exists) satisfies $\mathbb{0}\boxplus a=a=a\boxplus\mathbb{0},$ for every $a\in\mathcal{H}.$
  3. (iii)
    A hypersemigroup is a hypermagma $(\mathcal{H},\boxplus,\mathbb{0})$, where
    1. (a)
      The hyperaddition $\boxplus$ is associative in the sense that $(a_{1}\boxplus a_{2})\boxplus a_{3}=a_{1}\boxplus(a_{2}\boxplus a_{3})$ for all $a_{i}$ in $\mathcal{H}.$
    2. (b)
      $\mathbb{0}\in\mathcal{H}$ is the hyperzero.
  4. (iv)
    A hypernegative of an element $a$ in $(\mathcal{H},\boxplus,\mathbb{0})$ (if it exists) is an element “a-a” for which $\mathbb{0}\in a\boxplus(-a)$ and $\mathbb{0}\in(-a)\boxplus a$.
  5. (v)
    If the hypernegative $\mathbb{1}$ exists in $\mathcal{H},$ then we define $e=\mathbb{1}\boxplus(-\mathbb{1}).$
  6. (vi)
    A hypergroup is a hypersemigroup $(\mathcal{H},\boxplus,\mathbb{0})$ for which every element $a$ has a unique hypernegative denoted a-a, whereby, for all $a_{i}\in\mathcal{H},$
    1. (a)
      $(-)(a_{1}\boxplus a_{2})=(-)a_{2}\boxplus(-)a_{1}.$
    2. (b)
      $(-a_{1})=a_{1}.$
    3. (c)
      $\mathcal{H}$ is reversible in the following sense:
      $a_{3}\in a_{1}\boxplus a_{2}$ iff $a_{2}\in a_{3}\boxplus(-a_{1})$.p [^5] In [20], a mosaic is a reversible with a hyperzero.
  7. (vii)
    A hypersemiring (resp. hyperring) is an abelian [^6] In [20] this is called “canonical.” hypersemigroup (resp. hypergroup) $(\mathcal{H},\boxplus,\mathbb{0})$, providing $\mathcal{P}(\mathcal{H})$ with a natural elementwise multiplication, distributing over hyperaddition, making $\mathcal{P}(\mathcal{H})$ a pre-semialgebra pair over $\mathcal{T}_{\iota}=\mathcal{H}$ under the action $aS=\{as:s\in S\}$. [^7] Thus $\mathbb{0}a=\mathbb{0}$ for all $a\in\mathcal{H}.$ Although not necessarily distributive, $\mathcal{P}(\mathcal{H})$ but satisfies $(\boxplus_{i}S_{i})(\boxplus_{j}S_{j}^{\prime})\subseteq(\boxplus_{i}S_{i}S_{j}^{\prime})$, cf. [18, Proposition 1.1].
  8. (viii)
    A hypergroup $\mathcal{H}$ is a hyperfield if $\mathcal{H}\setminus\{\mathbb{0}\}$ is a multiplicative group.
Remark 3.2 ([1, 24]).
Although [20] uses hypermagmas, we work mostly with hypersemigroups and write $\mathcal{H}$ for $(\mathcal{H},\boxplus,\mathbb{0})$.
  1. (i)
    If it exists, the hyperzero of a hypersemigroup is unique, by the familiar elementary argument.
  2. (ii)
    $\mathcal{H}$ gives rise to a $\mathcal{H}$-module $\mathcal{P}(\mathcal{H}),$ with hyperaddition given by
    $S_{1}\boxplus S_{2}=\cup\{s_{1}\boxplus s_{2}:s_{i}\in S_{i}\}.$
  3. (iii)
    $\mathcal{P}(\mathcal{H})$ is a weakly admissible $\mathcal{T}$-bimodule in case
    1. (a)
      $\mathcal{T}=\{\mathbb{1}\}$.
    2. (b)
      $\mathcal{T}$ is a submonoid of $\mathcal{H}$, in particular when $\mathcal{H}\setminus\{\mathbb{0}\}$ is a monoid.
  4. (iv)
    Take any $\mathcal{T}$-submodule $S_{0}$ of $\mathcal{A}:=\mathcal{P}(\mathcal{H})$. Then we get a $\mathcal{T}$-pair $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$ where $\mathcal{P}(\mathcal{H})_{0}=\{S\subseteq\mathcal{H}:S_{0}\cap S\neq\emptyset\}.$
  5. (v)
    Take $S_{0}=\{\mathbb{0}\}$ in (iv). Thus $\mathcal{P}(\mathcal{H})_{0}=\{S\subseteq\mathcal{H}:\mathbb{0}\in S\},$ the standard definition.
Definition 3.3.
In Remark 3.2(v), the sub-pair of the pair $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$ generated by $\mathcal{H}$ is called the hyperpair of $\mathcal{H}.$  [^8] This is the definition in [2], because $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$. It could lead to complications if $\mathcal{P}(\mathcal{H})$ is not distributive, because then the submagma of the power set $\mathcal{P}(\mathcal{H})$ spanned by $\mathcal{H}$ need not be closed under multiplication.
Remark 3.4.
  1. (i)
    Any hyperpair has the important surpassing relation $\subseteq,$ i.e., $S_{1}\preceq S_{2}$ when $S_{1}\subseteq S_{2}.$ Of course $\mathcal{P}(\mathcal{H})$ is a lower semilattice with respect to $(\subseteq)$, but not ${\mathcal{P}}^{\sharp}(\mathcal{H}).$
  2. (ii)
    More generally, if $\mathcal{H}$ has a surpassing relation $\preceq$, then $\mathcal{P}(\mathcal{H})$ has the surpassing relation given by $S_{1}\preceq S_{2}$ if for each $s_{1}\in S_{1}$ there is $s_{2}\in S_{2}$ for which $s_{1}\preceq s_{2}.$
  3. (iii)
    For hypersemigroups $\mathcal{H}$, the condition that $b_{1}\subseteq b_{2}$ implies $f(b_{1})\subseteq f(b_{2})$ is automatic on $\mathcal{P}(\mathcal{H})$.
  4. (iv)
    If each $a\in\mathcal{H}$ has a unique hypernegative, then $(\mathcal{A},\mathcal{A}_{0})$ has a negation map given by applying the hypernegative element-wise.
Lemma 3.5.
Say a subset $S\subseteq\mathcal{P}(\mathcal{H})$ is weakly neutral if $a\in a\boxplus S$ for each $a\in\mathcal{H}$. The weakly neutral subsets comprise an $\mathcal{H}$-submagma of $\mathcal{P}(\mathcal{H})$.
Proof.
If $S_{i}$ are weakly neutral elements, then for any $a\in\mathcal{H},$ $a\boxplus(S_{1}\boxplus S_{2})=(a\boxplus S_{1})\boxplus S_{2}$, which contains $a\boxplus S_{2},$ which in turn contains $a.$
3.1. Examples of hypersemigroups and their pairs
Example 3.6.
Let us first recall some of the celebrated hyperfields (and their accompanying hyperpairs, cf. Definition 3.3), from [1]. In every case we take $\mathcal{T}=\mathcal{H},$ and the hypersum of two distinct nonzero elements is not a singleton.
  1. (i)
    The tropical hyperfield consists of the set $\mathcal{H}=\mathbb{R}\cup\{-\infty\}$, with -$\infty$ here as the zero element and 0 as the unit, equipped with the addition $a\boxplus b=\{a\}$ if $a>b$, $a\boxplus b=\{b\}$ if $a<b$, and $a\boxplus a=[-\infty,a]$.
  2. (ii)
    The hyperfield of signs $L:=\{0,1,-1\}$ has the intuitive multiplication law, and hyperaddition defined by $\boxplus 1=1, 11$$\boxplus-1=-1$, $x\boxplus 0=0\boxplus x=x$ for all $x,$ and $1\boxplus-1=-1\boxplus 1=\{0,1,-1\}.$
  3. (iii)
    The phase hyperfield. Let $S^{1}$ denote the complex unit circle, and take $\mathcal{H}=S^{1}\cup\{0\}$. Nonzero points $a$ and $b$ are antipodes if a=b.a=-b. Multiplication is defined as usual (so corresponds on $S^{1}$ to addition of angles). In this example, we denote an open arc of less than 180 degrees connecting two distinct points $a,b$ of the unit circle by (a,b)({a,b}). The hypersum is given, for $a,b\neq 0$, by
[…]
    $\langle\mathcal{T}\rangle$ contains only the points of $\mathcal{H}$, the elements of the form $a\boxplus b$ with $a,b\in S^{1}$, and subsets $C$ of $\mathcal{H}$, where either $C$ is an open half circle connecting an element $a\in S^{1}$ to a-a (which is obtained as the sum $a\boxplus b\boxplus(-a)$, where $b\in C$), or $C=\mathcal{H}$ (which is obtained as $a\boxplus b\boxplus(-a)\boxplus(-b)$).
    $e=\mathbb{1}\boxplus-\mathbb{1}=\{-\mathbb{1},0,\mathbb{1}\},$ so $e\boxplus e=e$. Thus the phase hyperfield is $\mathcal{A}_{0}$-idempotent, but is not $\circ$-distributive since $(a,b)e=(a,b)\cup\{\mathbb{0}\}\cup(-a,-b),$ whereas $(a,b)\boxplus(-a,-b)$ is the whole circle.
  4. (iv)
    (The weak phase hyperfield [1]; Akian-Gaubert’s modification of the phase hyperfield). The non-zero elements can still be represented by elements of the unit circle $S^{1}$. However, the hyperaddition differs – for $a,b\neq 0$, we now have
[…]
    where [a,b]{[a,b]} denotes the closed arc of less than $180$ degrees joining $a$ and $b$ (compare with the open arc $(a,b)$ in the phase hyperfield). Like the phase hyperfield, the weak phase hyperfield is not distributive.
  5. (v)
    One can further modify the weak phase hyperfield, to get a hypersemigroup which is not a hypergroup by defining $a\boxplus a=\emptyset.$
Example 3.7.
Let $\mathcal{H}$ be a set with an element stipulated $\mathbb{0}.$ In each of the following cases, $\mathcal{H}$ will be a hypersemigroup whose hyperpair $(\mathcal{A},\mathcal{A}_{0})$ will satisfy property N.
  1. (i)
    The following are uniquely negated, with $\mathcal{A}_{0}$ as in Remark 3.2(v).
    1. (a)
      Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\{\mathbb{0},s\}$, $\mathbb{0}\boxplus s=s\boxplus\mathbb{0}=s,$ and $s\boxplus s^{\prime}=\mathcal{H}$ for all nonzero $s\neq s^{\prime}$ in $\mathcal{H}.$ Here $\mathcal{A}=\mathcal{H}\cup\{\mathcal{H}\},$ and s=s.-s=s.
    2. (b)
      (As in [18]) When $|\mathcal{H}|\geq 3,$ define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\mathcal{H}\setminus s$ and $s\boxplus s^{\prime}=\{s,s^{\prime}\}$ for all nonzero $s\neq s^{\prime}$ in $\mathcal{H}.$ Here $\mathcal{A}=\mathcal{H}\cup_{s,s^{\prime}\in\mathcal{H}}\{s,s^{\prime}\}\cup_{s\in\mathcal{H}}(H\setminus\{s\})\cup\{\mathcal{H}\}.$ Again s=s.-s=s.
    3. (c)
      When $|\mathcal{H}|\geq 4,$ define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s^{\prime}=\mathcal{H}\setminus\{s,s^{\prime}\}$ for all nonzero $s,s^{\prime}\in\mathcal{H}.$ Here $\mathcal{A}=\mathcal{H}\cup_{s,s^{\prime}\in\mathcal{H}}\{s,s^{\prime}\}\cup_{s\in\mathcal{H}}(H\setminus\{s\})\cup_{s,s^{\prime}\in\mathcal{H}}(H\setminus\{s,s^{\prime}\})\cup\{\mathcal{H}\}.$ (Here $H\setminus\{s\}=(s^{\prime}\boxplus s^{\prime\prime})\boxplus s)$ where $s^{\prime},s^{\prime\prime}$ are distinct from $\{\mathbb{0},s\}.$)
  2. (ii)
    Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=s$ and $s\boxplus s^{\prime}=\mathcal{H}$ for all $s\neq s^{\prime}$ in $\mathcal{H}.$ This is idempotent, and is not uniquely negated.
  3. (iii)
    The following two examples are similar to those of [20, Proposition 3.1]. For all $a_{i}\in\mathcal{H},$
    1. (a)
      $a_{1}\boxplus a_{2}=\mathcal{H}$. This gives the admissible pair $(\mathcal{H}\cup\{\mathcal{H}\},\{\mathbb{0},\{\mathcal{H}\}\})$, which satisfies Property N.
    2. (b)
      $a_{1}\boxplus a_{2}=\emptyset$. $(\mathcal{H}\cup\{\emptyset\},\emptyset)$ is the minimal pair containing $\mathcal{H}$, but is not a hyperpair.
  4. (iv)
    $a_{1}\boxplus a_{2}=\{a_{1},a_{2}\}$. Hence $S_{1}\boxplus S_{2}=S_{1}\cup S_{2}$. $(\mathcal{H},\mathcal{H}_{0})$ fails Property N.
  5. (v)
    Now take $\mathcal{H}$ ordered, with $\mathbb{0}\in\mathcal{H}$ minimal. Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\mathcal{H}$ and $s\boxplus s^{\prime}=s^{\prime}$ for all $\mathbb{0}\neq s<s^{\prime}$ in $\mathcal{H}.$ Taking $\mathcal{A}=\mathcal{H}\cup\{\mathcal{H}\},$ and $\mathcal{A}_{0}=\{\mathcal{H}\}$, the hyperpair $(\mathcal{A},\mathcal{A}_{0})$ is admissible and $\mathcal{A}_{0}$-bipotent.
In each case above, $\mathcal{H}$ becomes a hyper-semiring when $\mathcal{H}$ is a cancellative multiplicative monoid.
3.2. Residue (quotient) hypermodules and hyperpairs
The following definition was inspired by Krasner [14], a bit more general than [25, Definition 3.1], but with the same verification as [25, Lemma 3.2].
Definition 3.8.
Suppose that $\mathcal{M}$ is a $C$-bimodule where the ring $C$ is viewed as a multiplicative monoid $\mathcal{T}$ as in Remark 2.3(i), and $G$ is a normal multiplicative subgroup of $\mathcal{T}$, in the sense that $bG=Gb$ for all $a\in\mathcal{M}$. Define the residue hypermodule $\mathcal{H}=\mathcal{M}/G$ over $\mathcal{T}/G$ to have multiplication induced by the cosets, and hyperaddition $\boxplus:\mathcal{H}\times\mathcal{H}\to\mathcal{P}(\mathcal{M})$ by
$b_{1}G\boxplus b_{2}G=\{cG:c\in b_{1}G+b_{2}G\}.$
When $\mathcal{M}=C$ is a field, the residue hypermodule $\mathcal{H}$ is a hyperfield (under multiplication of cosets), called the quotient hyperfield in the literature.
As defined in Remark 3.2(iv), $(\mathcal{M},\mathcal{M}_{0})$ is a pair, by [1, Proposition 7.21]. For the applications in the hyperfield literature one would take $\mathcal{M}$ to be a field with a multiplicative subgroup $\mathcal{T}$.
Remark 3.9.
In the residue hypermodule $\mathcal{H}=\mathcal{M}/G$,
  1. (i)
    $\mathbb{1}_{\mathcal{H}}=G.$
  2. (ii)
    $e=\mathbb{1}_{\mathcal{H}}\boxplus(-)\mathbb{1}_{\mathcal{H}}=\{g_{1}-g_{2}:g_{i}\in G\}.$
  3. (iii)
    According to Definition 2.9(iv), $Se=\{b_{1}g_{1}-b_{2}g_{2}:b_{i}\in S,g_{i}\in G\}$
  4. (iv)
    In particular, $ee=\{(g_{1}-g_{2})g_{3}\boxplus(g_{4}-g_{5})g_{6}:g_{i}\in G\}.$
  5. (v)
    $e\boxplus e=\{(g_{1}-g_{2})-(g_{4}-g_{5}):g_{i}\in G\}.$
$\mathcal{M}/G$ need not be $\circ$-distributive, since the phase hyperfield is a counterexample, cf. Examples 3.6(iii), but $\mathcal{M}/G$ does satisfy $ee=e\boxplus e$ by [25, Lemma 3.5].
Lemma 3.10.
Any surpassing relation $\preceq$ on $\mathcal{M}$ induces a surpassing relation on $\mathcal{M}/G$.
Proof.
As in Remark 3.4, define $b_{1}G\preceq b_{2}G$ when for each $g\in G$ there is $g^{\prime}\in G$ such that $b_{1}g\preceq b_{2}g$. We claim that if $b_{i}G\preceq b_{i}^{\prime}G$ for $i=1,2,$ then $b_{1}G\boxplus b_{2}G\preceq b_{1}^{\prime}G\boxplus b_{2}^{\prime}G.$
Indeed, if $b=b_{1}g_{1}+b_{2}g_{2}\in b_{1}G\boxplus b_{2}G,$ then taking $b_{i}^{\prime}g_{i}^{\prime}\succeq b_{i}g_{i}$ in $b_{i}^{\prime}G,$ we have
$b=b_{1}g_{1}+b_{2}g_{2}\preceq=b_{1}^{\prime}g_{1}^{\prime}+b_{2}^{\prime}g_{2}^{\prime}\in b_{1}^{\prime}G\boxplus b_{2}^{\prime}G.$
Example 3.11.
A huge assortment of examples of quotient hyperfields is given in [19, §2] [^9] In [25] more examples are obtained when addition and multiplication are reversed.. Here are some of them. We shall take $\mathcal{H}=\mathcal{M}/G,$ and its hyperpair $(\mathcal{A},\mathcal{A}_{0}),$ as in Remark 3.2(iv).
  1. (i)
    $G=\{\pm 1\}$. Then $\mathbb{0}\in\mathcal{H}\boxplus\mathcal{H},$ so $(\mathcal{A},\mathcal{A}_{0})$ has characteristic 2 and is multiplicatively idempotent.
  2. (ii)
    The Krasner hyperfield is $F/F^{\times}$, for any field $F$.
  3. (iii)
    The sign hyperfield is $\mathbb{R}/\mathbb{R}^{+}.$
  4. (iv)
    The phase hyperfield can be identified with the quotient hyperfield $\mathbb{C}/\mathbb{R}_{>0}$.
  5. (v)
    The weak phase hyperfield can be obtained by taking the quotient $\mathcal{F}/G$, where $\mathcal{F}=\mathbb{C}\{\{t^{\mathbb{R}}\}\}$, and $G$ is the group of (generalized) Puiseux series with positive real leading coefficient, where the leading coefficient is the coefficient $f_{\lambda}$ of the series $f=\sum_{\lambda\in\Lambda}f_{\lambda}t^{\lambda}$ such that $\lambda$ is the minimal element of $\{\lambda\in\Lambda:f_{\lambda}\neq 0\}$.
Lemma 3.12.
If $\mathcal{M}$ is free over $\mathcal{T}$ with a $G$-invariant base $\mathcal{B}$}, then $\bar{\mathcal{M}}:=\mathcal{M}/G$ is free over $\bar{\mathcal{B}}:=\{b_{i}G,b_{i}\in{\mathcal{B}}\}$,
Proof.
We may assume ${\mathcal{B}}=\{b_{i}g_{i}:i\in I\}.$ Clearly $\bar{\mathcal{B}}$ spans $\bar{\mathcal{M}}$. But if $\sum(a_{i}G)b_{i}G=\sum(a^{\prime}_{i}G)b_{i}G,$ then $\sum a_{i}b_{i}^{\prime\prime}g_{i}=\sum a^{\prime}_{i}b_{i}g_{i}^{\prime}g_{i}$ for suitable $g_{i},g_{i}^{\prime},g_{i}^{\prime\prime}\in G,$ implying $a_{i}g_{i}g^{\prime\prime}_{i}=a_{i}^{\prime}g_{i}^{\prime},$ and thus $a_{i}G=a_{i}^{\prime}G.$
4. Tensor products of bimodules
In [12] tensor products were treated from the point of view of semirings, following [13, 26]. There is a delicate issue here, concerning which are the morphisms in our category. One may start with the more straightforward instance of homomorphisms. Then it is rather easy to construct a tensor product satisfying the theory of [9, 21], whose theorems are applicable.
When we bring in weak morphisms on both sides, things become much more complicated, and our approach becomes more compatible with [4] and [20].
4.1. Construction of the tensor product
Example 4.1.
The free abelian semigroup $(\mathcal{S}(X),+)$ on a set $X$ is the set of formal sums without parentheses, defined by $X\subseteq\mathcal{S}(X)$ of length 1, and inductively, for $v,w\in\mathcal{S}(X)$ of lengths $m$ and $n$, $v+w=w+v\in\mathcal{S}(X)$ of length $m+n.$
We follow the exposition of the classical tensor product, as in [23, Chapter 18]. Throughout this section, $\mathcal{M}_{1}$ is a right $\mathcal{T}$-module and $\mathcal{M}_{2}$ is a left $\mathcal{T}$-module, for convenience [^10] More generally, for $\mathcal{M}_{1}$ a right $\mathcal{T}$-magma and $\mathcal{M}_{1}$ a left $\mathcal{T}$-magma, we can define the free magma $(\mathcal{F}(X),*)$ on a set $X$ is the set of formal expressions with parentheses, defined by $X\subseteq\mathcal{F}(X)$ of length 1, and inductively, for $v,w\in\mathcal{F}(X)$ of lengths $m$ and $n$, $(v*w)\in\mathcal{F}(X)$ of length $m+n,$ and the $\mathcal{T}$-tensor product magma $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ to be the magma $\mathcal{F}(\mathcal{M}_{1}\times\mathcal{M}_{2})/\Phi,$ where $\Phi$ is the congruence generated by $\big(\big(v_{1}*w_{1},x_{2}\big),\big(v_{1},x_{2})*(w_{1},x_{2}\big)\big),\ \big(\big(x_{1},v_{2}*w_{2}\big),(x_{1},v_{2})*(x_{1},w_{2})\big),$ $\big((x_{1}a,x_{2}),(x_{1},ax_{2})\big)$ $\forall x_{i},v_{i},w_{i}\in\mathcal{\mathcal{M}}_{i},$ $a\in\mathcal{T}$. The remainder of this subsection would go over, mutatis mutandis, but the notation is more involved..
Definition 4.2.
  1. (i)
    Define the $\mathcal{T}$-tensor product semigroup $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ to be the semigroup $\mathcal{S}(\mathcal{M}_{1}\times\mathcal{M}_{2})/\Phi,$ where $\Phi$ is the congruence generated by all
    $\bigg(\big(v_{1}+w_{1},x_{2}\big),\big(v_{1},x_{2})+(w_{1},x_{2}\big)\bigg),\quad\bigg(\big(x_{1},v_{2}+w_{2}\big),(x_{1},v_{2})+(x_{1},w_{2})\bigg),$ (4.1)
    $\bigg((x_{1}a,x_{2}),(x_{1},ax_{2}),\bigg)$ (4.2)
    $\forall x_{i},v_{i},w_{i}\in\mathcal{\mathcal{M}}_{i},$ $a\in\mathcal{T}$.
  2. (ii)
    When $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ have negation maps, we incorporate the negation map into the definition of tensor product, by enlarging $\Phi$ to include also $(\big((-)v_{1},v_{2}\big),\big(v_{1},(-)v_{2})\big)$ for all $v_{i}\in\mathcal{M}_{i}.$
  3. (iii)
    A simple tensor of $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is an element of the form $v_{1}\otimes v_{2}$ for $v_{i}\in\mathcal{M}_{i}.$
Note that this matches the situation that $\mathcal{T}$ is a semiring, in which case we would just forget its addition.
Remark 4.3.
  1. (i)
    The sum of two non-simple tensors could be simple, modulo associativity. For example, if $v_{3}=v_{1}+v_{2}$ then $(v_{3}\otimes v_{2})+(v_{2}\otimes v_{1})$ and $(v_{2}\otimes v_{2})+(v_{1}\otimes v_{3})$ are non-simple whereas
    $(v_{3}\otimes v_{2})+(((v_{2}\otimes v_{1})+(v_{2}\otimes v_{2}))+(v_{1}\otimes v_{3}))$ $=(v_{3}\otimes v_{2})+((v_{2}\otimes v_{3}))+(v_{1}\otimes v_{3}))$ (4.3)
    $=(v_{3}\otimes v_{2})+(v_{3}\otimes v_{3})=v_{3}\otimes(v_{2}+v_{3}).$
  2. (ii)
    If $a_{1}v_{2}=a_{2}w_{2},$ then $v_{1}a_{1}\otimes v_{2}+w_{1}a_{2}\otimes w_{2}=(v_{1}+w_{1})\otimes a_{2}w_{2}$. This process of “recombining” is straightforward when $\mathcal{T}$ is a group, since then it is applicable for any sum
    $v_{1}\otimes v_{2}+w_{1}\otimes w_{2}=(v_{1}a_{1}^{-1}+w_{1}a_{2}^{-1})\otimes a_{2}w_{2}.$
Remark 4.4.
As in [4], one could define the tensor product $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$ of a right $\mathcal{T}$-monoid $\mathcal{T}_{1}$ and a left $\mathcal{T}$-monoid $\mathcal{T}_{2}$ to be $(\mathcal{T}_{1}\times\mathcal{T}_{2})/\Phi$ where where $\Phi$ is the congruence generated by all $\bigg((x_{1}a,x_{2}),(x_{1},ax_{2})\bigg)$, $\forall x_{i}\in\mathcal{\mathcal{M}}_{i},$ $a\in\mathcal{T}$.
4.1.1. Tensor products of bimodule prepairs
In order to define the tensor product bimodule prepair, we need to define $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}$.
Lemma 4.5.
$((\mathcal{M}_{1}\otimes\mathcal{M}_{2}),(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule prepair.
Proof.
$((\mathcal{M}_{1}\otimes\mathcal{M}_{2}),(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0})$ is closed under left multiplication by $\mathcal{T}_{1}$ and right multiplication by $\mathcal{T}_{2}.$ $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}$ is closed under addition. ∎
Write $\mathcal{T}^{\operatorname{op}}$ for the opposite monoid. Then any $\mathcal{T}_{1},\mathcal{T}_{2}$ bimodule is a $\mathcal{T}_{1}\times\mathcal{T}_{2}^{\operatorname{op}}$ module, under the operation $(a_{1},a_{2})b=a_{1}ba_{2},$ by the obvious verification.
Lemma 4.6.
Suppose $\mathcal{T}_{i}\supseteq\mathcal{T}$ are commutative monoids. Then $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})$ is a $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$-module, under the action $(a_{1}\otimes a_{2})b=a_{1}ba_{2}.$
Proof.
We have the $\mathcal{T}_{1}\times\mathcal{T}_{2}$ action, which induces a $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$-action since $(a_{1}a\otimes a_{2})b=a_{1}aba_{2}=a_{1}baa_{2}=(a_{1}\otimes aa_{2})b.$
Corollary 4.7.
If $\mathcal{A}_{i}$ are $\mathcal{T}_{i}$-semialgebras with each $\mathcal{T}_{i}\subseteq\mathcal{T}$ and if $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ is a semialgebra as in Corollary 4.16(vi), then $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ is a $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$-semialgebra.
Definition 4.8.
Suppose $(\mathcal{M}_{1},{\mathcal{M}_{1}}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule prepair and $(\mathcal{M}_{2},{\mathcal{M}_{2}}_{0})$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule prepair. Define $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}:=({\mathcal{M}_{1}}_{0}\otimes_{\mathcal{T}}\mathcal{M}_{2})+(\mathcal{M}_{1}\otimes_{\mathcal{T}}{\mathcal{M}_{2}}_{0}).$
4.1.2. Pre-orders on the tensor product
Note that the multiplicative structure in the tensor product could be carried out along the lines of [4], but coping with addition makes the situation much more intricate.
Lemma 4.9.
If $\mathcal{M}_{i}$ each have a pre-order $\preceq$, then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ has the pre-order given by $x\preceq y$ if whenever we can write $x=\sum v_{i}\otimes w_{i}$ then we can write $y=\sum v^{\prime}_{i}\otimes w^{\prime}_{i}$ such that $v_{i}\preceq v^{\prime}_{i}$ and $w_{i}\preceq w_{i}^{\prime}$ for each $i.$
Proof.
Clearly this condition is transitive and passes to sums. ∎
Remark 4.10.
In general, for semialgebra pairs, one conceivably might have $\mathbb{1}\otimes\mathbb{1}\in(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}.$ In such a case the tensor product of surpassing relations cannot be a surpassing relation. But one must realize that even in the classical situation we could have $\mathbb{1}\otimes\mathbb{1}=\mathbb{0}.$
4.2. Properties of tensor products
Definition 4.11.
  1. (i)
    A 1-balanced map is a function $\psi:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{M}$ where $\mathcal{M}$ is a semigroup, satisfying
    $\psi(v_{1}a,v_{2})=\psi(v_{1},av_{2}),$ (4.4)
    for all $v_{i}\in\mathcal{M}_{i},$ $a\in\mathcal{T}.$
  2. (ii)
    A balanced map is a 1-balanced map satisfying
    $\psi(v_{1}+w_{1},x_{2})=\psi(v_{1},x_{2})+\psi(w_{1},x_{2}),$ (4.5)
    $\psi(x_{1},v_{2}+w_{2})=\psi(x_{1},v_{2})+\psi(x_{1},w_{2}),$ (4.6)
    for all $v_{i},w_{i},x_{i}\in\mathcal{M}_{i}.$
When $\mathcal{M}_{1},\mathcal{M}_{2}$ are given with negation maps, then a balanced map is required to satisfy $((-)v_{1})\otimes v_{2}=v_{1}\otimes(-v_{2}),$ which we define to be $(-)(v_{1}\otimes v_{2})$.
Theorem 4.12.
Suppose $\mathcal{M}_{1},\mathcal{N}_{1}$ are $(\mathcal{T}_{1},\mathcal{T})$-bimodules and $\mathcal{M}_{2},\mathcal{N}_{2}$ are $(\mathcal{T},\mathcal{T}_{2})$-bimodules. If $f_{1}:\mathcal{M}_{1}\to\mathcal{N}_{1}$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule homomorphism and $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule homomorphism, then there is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule homomorphism
$f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes_{\mathcal{T}}\mathcal{N}_{2}$
given by $(f_{1}\otimes f_{2})(v_{1}\otimes v_{2})=f_{1}(v_{1})\otimes f_{2}(v_{2}).$
Proof.
The balanced map $\Psi:(v_{1},v_{2})\mapsto f_{1}(v_{1})\otimes f_{2}(v_{2})$ sends $\Phi_{1}$ to $\Phi_{2},$ seen by applying $f_{1}$ and $f_{2}$ to the respective components. ∎
Corollary 4.13.
Assume throughout that $\mathcal{M}_{1}$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule and $\mathcal{M}_{2}$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule.
  1. (i)
    If $\mathcal{M}_{1}$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule, then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is a $\mathcal{T}_{1}$-module, with the $\mathcal{M}_{1}$-action $a(v_{1}\otimes v_{2})=av_{1}\,\otimes v_{2}$. Analogously for $\mathcal{M}_{2}.$
  2. (ii)
    If $\mathcal{M}_{1}$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule and $\mathcal{M}_{2}$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule, then
    1. (a)
      $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule, with the actions $a_{1}(v_{1}\otimes v_{2})a_{2}=a_{1}v_{1}\,\otimes\,v_{2}a_{2}$.
    2. (b)
      Any two bimodule homomorphisms $f_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}$ factor through a bimodule homomorphism $f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes\mathcal{N}_{2}.$ If $\mathcal{A}_{1},\mathcal{A}_{2}$ are weakly admissible with $\mathbb{1}\in\mathcal{T}$, then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ are weakly admissible over $\mathbb{1}\otimes\mathcal{T}=\mathcal{T}\otimes\mathbb{1}$.
  3. (iii)
    Any homomorphism $f:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}$ is determined by its action on simple tensors.
  4. (iv)
    If both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are $\mathcal{T}$-bimodules, then $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}\cong\mathcal{M}_{2}\otimes_{\mathcal{T}}\mathcal{M}_{1}$ via $(v_{1}\otimes v_{2})\mapsto v_{2}\otimes v_{1}.$
  5. (v)
    If both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are $\mathcal{T}$- semialgebras, then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is a semialgebra with multiplication given by $(v_{1}\otimes v_{2})(w_{1}\otimes w_{2})=(v_{1}w_{1}\otimes v_{2}w_{2}).$
Proof.
Repeated applications of Theorem 4.12.
(i) Take $f_{1}$ to be the identity, and $f_{2}$ to be right multiplication by $a,$ and symmetrically.
(ii) Take $f_{1}$ to be left multiplication by $a_{2}$, and $f_{2}$ to be right multiplication by $a_{2}$.
(iii) $f$ is determined by the balanced map $\psi:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}$ given by $\psi(v_{1},v_{2})=f(v_{1}\otimes v_{2}).$
(iv) Take the balanced map $(v_{1},v_{2})\mapsto v_{2}\otimes v_{1}$.
(v) First define multiplication by $w_{1}\otimes\mathbb{1}$ via the balanced map $(v_{1},v_{2})\mapsto(v_{1}\otimes w_{1}v_{2}),$ and then left multiplication by $w_{2}\otimes\mathbb{1}$. ∎
Remark 4.14.
  1. (i)
    If one takes $\mathcal{T}=\{\mathbb{1}\}$ as in Remark 3.2(iii), then Equation (4.2) is redundant.
  2. (ii)
    Associativity of addition in $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ follows from associativity of addition in $\mathcal{S}(\mathcal{M}_{1})\times\mathcal{S}(\mathcal{M}_{2}).$
Corollary 4.15.
The tensor product in this subsection satisfies the criteria of [21, Definitions 1,2].
Proof.
By Corollary 4.13. ∎
Corollary 4.16.
Suppose $\mathcal{M},\mathcal{M}_{i}$ are $\mathcal{T}_{1},\mathcal{T}$-bimodules and $\mathcal{N},\mathcal{N}_{j}$ are $\mathcal{T},\mathcal{T}_{2}$-bimodules.
  1. (i)
    $\mathcal{T}_{1}\otimes_{\mathcal{T}_{1}}\mathcal{M}\cong\mathcal{M}$ and $\mathcal{M}\otimes_{\mathcal{T}}{\mathcal{T}}\cong\mathcal{M}$,
  2. (ii)
    $(\oplus\mathcal{M}_{i})\otimes N\cong\oplus(\mathcal{M}_{i}\otimes N)$,
  3. (iii)
    $\mathcal{M}\otimes(\oplus\mathcal{N}_{i})\cong\oplus(M\otimes\mathcal{N}_{i})$,
  4. (iv)
    $(\oplus\mathcal{M}_{i})\otimes(\oplus\mathcal{N}_{j})\cong\oplus(\mathcal{M}_{i}\otimes\mathcal{N}_{j})$, as $\mathcal{T}_{1},\mathcal{T}_{2}$-bimodules.
  5. (v)
    If $\mathcal{N}$ is free over $\mathcal{T}$ with base $\{b_{i}:i\in I\}$ then every element of $\mathcal{M}\otimes\mathcal{N}$ can be written in the form $\sum_{i\in I}v_{i}\otimes b_{i}$, for uniquely determined $v_{i}\in\mathcal{M}$.
  6. (vi)
    If $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are semialgebras over $\mathcal{T}$ with $\mathcal{A}_{2}$ free, then $\mathcal{A}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{2}$ is a semialgebra. (Likewise if $\mathcal{A}_{1}$ is free.)
Proof.
(i) –(iv) These are results in [9, 21].
(v) The projection $\pi_{j}:\mathcal{N}\to\mathcal{N}$ onto the $j$ coordinate yields a projection $1\otimes\pi_{j}:\mathcal{M}\otimes\mathcal{N}\to\mathcal{M}\otimes\mathcal{N}$. If $\sum_{i\in I}v_{i}\otimes b_{i}=\sum_{i\in I}v^{\prime}_{i}\otimes b_{i}$ then applying $1\otimes\pi_{j}$ yields $v_{j}\otimes b_{j}=v^{\prime}_{j}\otimes b_{j}.$
(vi) Write tensors uniquely in terms of a base, by $\sum_{i\in I}v_{i}\otimes b_{i}$ as in (v), and then work with structure constants, writing $b_{i}b_{j}=\sum a_{ijk}b_{k}$, and putting $(\sum_{i\in I}v_{i}\otimes b_{i})(\sum_{j\in I}v_{j}\otimes b_{j})=(\sum_{i\in I}a_{ijk}v_{i}v_{j}\otimes b_{k})$. ∎
Remark 4.17.
We do not see how to remove the freeness restriction in (vi). We could follow [23, Theorem 18.21], to get $\mathcal{T}$-module isomorphisms
$(\mathcal{A}_{1}\otimes\mathcal{A}_{2})\otimes(\mathcal{A}_{1}\otimes\mathcal{A}_{2})\to(\mathcal{A}_{1}\otimes\mathcal{A}_{1})\otimes(\mathcal{A}_{2}\otimes\mathcal{A}_{2}),$
but now we encounter the difficulty that $\mu:\mathcal{A}_{1}\to\mathcal{A}_{1}$ might not induce a map $\bar{\mu}:\mathcal{A}_{1}\otimes\mathcal{A}_{1}\to\mathcal{A}_{1},$ since we don’t know that $(b_{1}+b_{2})\otimes(b_{1}+b_{2})=\sum_{i,j=1}^{2}b_{i}\otimes b_{j}$ and even if it does, what is $\bar{\mu}\otimes\bar{\mu}^{\prime}$?
Corollary 4.18.
Suppose $\mathcal{M}_{i}$ are $\mathcal{T}_{i},\mathcal{T}_{i+1}$-bimodules. Then $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})\otimes\mathcal{M}_{3}\cong\mathcal{M}_{1}\otimes(\mathcal{M}_{2}\otimes\mathcal{M}_{3})$ as $\mathcal{T}_{1},\mathcal{T}_{3}$-bimodules.
Proof.
The proof is from [21, Theorem 2]. ∎
Proposition 4.19.
If $\mathcal{M}_{i}=\mathcal{P}(\mathcal{H}_{i})$ where $\mathcal{M}_{1}$ is a right $\mathcal{T}$-module and $\mathcal{M}_{2}$ is a left $\mathcal{T}$-module. then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is $\subseteq$-distributive.
Proof.
We shall show that $S\otimes(S_{1}^{\prime}\boxplus S_{2}^{\prime})\subseteq SS_{1}^{\prime}\,\boxplus\,SS_{2}^{\prime}$ for all $S\subseteq\mathcal{H}_{1}$ and $S_{1}^{\prime},S_{2}^{\prime}\subseteq\mathcal{H}_{2}$. By definition, $S\otimes(S_{1}^{\prime}\boxplus S_{2}^{\prime})=\{a\otimes a^{\prime}:a\in S,\ a^{\prime}\in S_{1}^{\prime}\boxplus S_{2}^{\prime}\}=\{a(S_{1}^{\prime}\boxplus S_{2}^{\prime}):a\in S\}.$ Thus we need to show that $a\otimes(S_{1}^{\prime}\boxplus S_{2}^{\prime})\subseteq a\otimes S_{1}^{\prime}\,\boxplus\,a\otimes S_{2}^{\prime}$ for each $a\in S.$ But for each $a^{\prime}\in S_{1}^{\prime}\boxplus S_{2}^{\prime}$, if $a^{\prime}\in b_{1}^{\prime}\boxplus b_{2}^{\prime}$ for suitable nonzero $b_{i}^{\prime}\in S_{2}^{\prime}$, then $a\otimes a^{\prime}\in(a\otimes b_{1}^{\prime})\boxplus(a\otimes b_{2}^{\prime}),$ and we are done by induction on height. Hence we may assume that $a^{\prime}\in\mathcal{H}_{2},$ and the assertion is immediate. ∎
4.3. Tensor extensions
Definition 4.20.
Suppose $\mathcal{M}$ is a $\mathcal{T}$-module, and $\mathcal{T}^{\prime}$ is a monoid containing $\mathcal{T}$.
  1. (1)
    Define the $\mathcal{T}$-tensor extension $\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{M}$ to be the semigroup $\mathcal{S}(\mathcal{T}^{\prime}\times\mathcal{M})/\Phi,$ where $\Phi$ is the congruence generated by all
    $\bigg(\big(a^{\prime},v+w\big),\big(a^{\prime},v\big)+\big(a^{\prime},w\big)\bigg)\qquad\bigg((a^{\prime}a,w),(a^{\prime},aw)\bigg)\qquad\forall a,a^{\prime}\in\mathcal{T},\ v,w\in\mathcal{\mathcal{M}}.$ (4.7)
    When $\mathcal{M}$ has a negation map, we incorporate the negation map into the definition of tensor product, by enlarging $\Phi$ to include also $(\big((-)a,w),\big(a,(-)w)\big)$ for $a\in\mathcal{T},$ $w\in\mathcal{M}.$
Remark 4.21.
One sees easily that if $\mathcal{M}$ is left admissible (resp. left weakly admissible), then so is $\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{M}$.
Now we modify §4.2, starting with Definition 4.11.
Definition 4.22.
  1. (i)
    A balanced map is a function $\Psi:\mathcal{T}^{\prime}\times\mathcal{M}\to\mathcal{N}$ where $\mathcal{N}$ is a semigroup, satisfying
    $\Psi(a^{\prime},w_{1}+w_{2})=\Psi(a^{\prime},w_{1})+\Psi(a^{\prime},w_{2}),$ (4.8)
    $\Psi(v_{1}a,w_{1})=\Psi(v_{1},aw_{1}),$ (4.9)
    for all $a\in\mathcal{T},$ $a^{\prime}\in\mathcal{T}^{\prime},$ $v_{i},w_{i}\in\mathcal{M}_{i}.$
Lemma 4.23.
$\mathcal{M}^{\prime}:=\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{M}$ is a $\mathcal{T}^{\prime}$-module.
Proof.
The map $\mathcal{T}^{\prime}\times\mathcal{M}\to\mathcal{M}^{\prime}$ given by $(a,\sum a_{i}\otimes y_{i})\mapsto\sum aa_{i}\otimes y_{i}$ factors through the congruence $\Phi.$
As in Theorem 4.12, we have the key observation leading to the isomorphism theorems.
Theorem 4.24.
  1. (i)
    Suppose $\mathcal{T}_{1},\mathcal{T}_{2}$ are monoids containing $\mathcal{T}$, and $\mathcal{M}_{i}$ are $\mathcal{T}$-modules for $i=1,2$. If $f_{1}:\mathcal{T}_{1}\to\mathcal{T}_{2}$ is a monoid homomorphism fixing $\mathcal{T}$ and $f_{2}:\mathcal{M}_{1}\to\mathcal{M}_{2}$ is a module homomorphism, then there is a $\mathcal{T}$-module homomorphism
    $f_{1}\otimes f_{2}:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{T}_{2}\otimes_{\mathcal{T}}\mathcal{N}_{2}$
    given by $(f_{1}\otimes f_{2})(a\otimes w)=f_{1}(a)\otimes f_{2}(w).$
  2. (ii)
    If $\mathcal{M}$ is a $\mathcal{T}$-semialgebra and $\mathcal{T}_{1}$ is commutative, then $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}$ is a $\mathcal{T}_{1}$-semialgebra.
  3. (iii)
    Suppose $\mathcal{T}_{1},\mathcal{T}_{2}$ are monoids containing $\mathcal{T}$, and $\mathcal{A}_{1},\mathcal{A}_{2}$ are $\mathcal{T}$-semialgebras. If $f_{1}:\mathcal{T}_{1}\to\mathcal{T}_{2}$ is a monoid homomorphism and $f_{2}:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a semialgebra homomorphism, then there is a $\mathcal{T}_{1}$-semialgebra homomorphism
    $f_{1}\otimes f_{2}:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{1}\to\mathcal{T}_{2}\otimes_{\mathcal{T}}\mathcal{A}_{2}$
    given by $(f_{1}\otimes f_{2})(a\otimes b)=f_{1}(a)\otimes f_{2}(b),$ for $a\in\mathcal{T}_{1},\ b\in\mathcal{A}_{1}$
  4. (iv)
    If $\mathcal{T}_{1}$ is the disjoint union $\underset{\cdot}{\cup}\,c_{i}\mathcal{T}$ then every element of $\mathcal{T}_{1}\otimes\mathcal{M}_{2}$ can be written in the form $\sum_{i\in I}a_{i}\otimes b_{i}$, for uniquely determined $b_{i}\in\mathcal{M}_{2}$.
Proof.
(i) The balanced map $\Psi:(a,w)\mapsto f_{1}(a)\otimes f_{2}(w)$ sends the congruence $\Phi_{1}$ of Definition 4.2 to $\Phi_{2},$ seen by applying $f_{1}$ and $f_{2}$ to the respective components.
(ii) The projection $\pi_{j}:\mathcal{T}_{1}\to\mathcal{T}$ onto the coefficient of $c_{j}$ yields a projection $1\otimes\pi_{j}:\mathcal{T}_{1}\otimes\mathcal{M}_{2}\to\mathcal{M}_{2}$. If $\sum_{i\in I}a_{i}\otimes b_{i}=\sum_{i\in I}a_{i}\otimes b^{\prime}_{i}$ then applying $\pi_{j}\otimes 1$ yields $b_{j}=b^{\prime}_{j}.$
(iii) $f_{1}\otimes f_{2}$ respects multiplication, namely
$(f_{1}\otimes f_{2})((a_{1}\otimes w_{1})(a_{2}\otimes w_{2}))=((f_{1}\otimes f_{2})(a_{1}\otimes w_{1}))((f_{1}\otimes f_{2})(a_{2}\otimes w_{2})).$
(iv) Consider the balanced map $f_{a_{2};w_{2}}\mathcal{T}_{1}\times\mathcal{M}\to\mathcal{T}_{1}\otimes\mathcal{M}$ given by $(a,w)\mapsto(aa_{2})\otimes(aw_{2})$. This induces a map $\mathcal{T}_{1}\otimes\mathcal{M}\to\mathcal{T}_{1}\otimes\mathcal{M}$. ∎
The most important case is when $\mathcal{T}_{1}=\mathcal{T}$ and $f_{1}:\mathcal{T}\to\mathcal{T}_{2}$ is the inclusion map.
Corollary 4.25.
Suppose $\mathcal{T}_{2}$ is a monoid containing $\mathcal{T}$, and $\mathcal{M}_{1},\mathcal{M}_{2}$ are $\mathcal{T}$-modules. If $f:\mathcal{M}_{1}\to\mathcal{M}_{2}$ is a bimodule homomorphism, then there is a bimodule homomorphism
$1\otimes f:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{1}\to\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$
given by $(1\otimes f)(a\otimes w)=a\otimes f(w).$
Corollary 4.26.
Suppose that $\mathcal{A}$ is a $\mathcal{T}$-module.
  1. (i)
    If $\mathcal{T}\subseteq\mathcal{T}_{1}\subseteq\mathcal{T}_{2},$ then $\mathcal{T}_{2}\otimes(\mathcal{T}_{1}\otimes\mathcal{A})\cong\mathcal{T}_{2}\otimes\mathcal{A}.$
  2. (ii)
    $\mathcal{T}\otimes\mathcal{A}\cong\mathcal{A}$ as semialgebras.
  3. (iii)
    If $\mathcal{A}_{1}\to\mathcal{A}_{2}$ is an epimorphism of $\mathcal{T}$-semialgebras, then $\mathcal{T}_{1}\otimes\mathcal{A}_{1}\to\mathcal{T}_{1}\otimes\mathcal{A}_{2}$ is an epimorphism.
Proof.
As in [21, Theorem 3]. ∎
$4.$ Residue tensors viewed functorially
One nice feature of this approach is that the residue functor preserves tensor products. The notation for hyperaddition in residue hypermodules continues to be $\boxplus.$
Proposition 4.27.
Suppose that $\mathcal{T}_{1},\mathcal{T}_{2}$ are abelian monoids, $\mathcal{M}_{i}$ are $\mathcal{T}_{i}$-modules for $i=1,2,$ and $G$ is a common subgroup of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}.$ The tensor product of the residue hypermodules $\mathcal{H}_{1}=\mathcal{M}_{1}/G_{1}$ and $\mathcal{M}_{2}=\mathcal{M}_{2}/G_{2}$ is isomorphic to $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})/(G_{1}\otimes G_{2})$.
Proof.
The natural map $\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to(\mathcal{M}_{1}/G_{1})\otimes(\mathcal{M}_{2}/G_{2})$ given by $y_{1}\otimes y_{2}\mapsto y_{1}G_{1}\otimes y_{2}G_{2}$ preserves the defining congruence, thereby inducing a map of the cosets, which is 1:1 and onto. ∎
5. Tensor products of morphisms which need not be homomorphisms
Now we get to the sticky part. The proof of Theorem 4.12 fails for morphisms which are not homomorphisms, and thus its usefulness for hypersemigroups is limited. Our goal nevertheless is to salvage whatever information that we can. We do not even have an obvious functor for residue hypersemigroups in general.
Remark 5.1.
Suppose $f_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}$ are homomorphisms of $\mathcal{T}_{i}$-modules, for $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ abelian, and $G$ is a common subgroup of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}.$ Then writing $\overline{\mathcal{M}_{i}}=\mathcal{M}_{i}/G$, $\overline{\mathcal{N}_{i}}=\mathcal{N}_{i}/G,$ and $\bar{f}_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}/G$ for the weak morphism induced by $f_{i}$, we would want a balanced map $\Psi:=\bar{f}_{1}\times\bar{f}_{2}:{\mathcal{M}_{1}}\times{\mathcal{M}_{2}}\to\overline{\mathcal{N}_{1}}\otimes\overline{\mathcal{N}_{2}}$ given by
$(y_{1}\times y_{2})\mapsto f_{1}(y_{1})Gf_{2}(y_{2})=\{f_{1}(y_{1})gf_{2}(y_{2}):g\in G\},$
to induce a weak morphism $\overline{\mathcal{M}_{1}}\otimes\overline{\mathcal{M}_{2}}\to\overline{\mathcal{N}_{1}}\otimes\overline{\mathcal{N}_{2}}$.
But $\Psi(v_{1}+w_{1},x_{2})=f_{1}(v_{1}+w_{1})Gf_{2}(x_{2})$ whereas $\Psi(v_{1},x_{2})\boxplus\Psi(w_{1},x_{2})=f_{1}(v_{1})Gf_{2}(x_{2})\boxplus f_{1}(w_{1})Gf_{2}(x_{2}),$ which could be larger. Thus, the tensor product of homomorphisms modulo the residue map is too naive in general (but see Corollary 5.9 below).
Remark 5.2.
In [20], to expedite a categorical approach for hypermagmas, the tensor product of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ is defined as satisfying
$(v_{1}\otimes v_{2})\boxplus(v_{1}\otimes v_{2})=(v_{1}\boxplus v_{1})\otimes v_{2}\ \cup\ v_{1}\otimes(v_{2}\boxplus v_{2}),$ (5.1)
$(v_{1}\otimes v_{2})\boxplus(v_{1}^{\prime}\otimes v_{2})=(v_{1}\boxplus v_{1}^{\prime})\otimes v_{2},$ (5.2)
$(v_{1}\otimes v_{2})\boxplus(v_{1}\otimes v_{2}^{\prime})=v_{1}\otimes(v_{2}\boxplus v_{2}^{\prime}),$ (5.3)
$(v_{1}\otimes v_{2})\boxplus(v_{1}^{\prime}\otimes v_{2}^{\prime})=\emptyset,$ (5.4)
for $v_{1}\neq v_{1}^{\prime}\in\mathcal{M}_{1}$ and $v_{2}\neq v_{2}^{\prime}\in\mathcal{M}_{2}$. As they point out, this is the smallest hypermagma satisfying $(v_{1}\boxplus v_{1}^{\prime})\otimes v_{2}\subseteq(v_{1}\otimes v_{2})\boxplus(v_{1}^{\prime}\otimes v_{2})$ and $v_{1}\otimes(v_{2}\boxplus v_{2}^{\prime})\subseteq(v_{1}\otimes v_{2})\boxplus(v_{1}\otimes v_{2}^{\prime})$ for all $v_{1},v_{1}^{\prime}\in\mathcal{M}_{1}$ and $v_{2},v_{2}^{\prime}\in\mathcal{M}_{2}$, and has excellent functorial properties, but there is considerable collapsing, and associativity of addition even fails in the bipotent situation:
Example 5.3.
If $v_{1}\boxplus v_{2}=v_{1}$ and $w_{1}\boxplus w_{2}=w_{1}$ then $((v_{1}\otimes w_{1})\boxplus(v_{1}\otimes w_{2}))\boxplus((v_{2}\otimes w_{1})\boxplus(v_{2}\otimes w_{2}))=(v_{1}\otimes w_{1})\boxplus(v_{2}\otimes w_{1})=(v_{1}\otimes w_{1})$ whereas
$(v_{1}\otimes w_{1})\boxplus((v_{1}\otimes w_{2})\boxplus(v_{2}\otimes w_{1}))\boxplus(v_{2}\otimes w_{2}))=(v_{1}\otimes w_{1})\boxplus(\emptyset\boxplus(v_{2}\otimes w_{2}))=\emptyset.$
Thus, although [20] is appropriate for $\preceq$-morphisms in the category of hypermagmas, we feel that the more traditional approach may be needed in the category of hypermodules.
5.1. Defining $f_{1}\otimes f_{2}$ via simple tensors
Given morphisms $f_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}$ for $i=1,2,$ we search for a well-defined morphism $f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes\mathcal{N}_{2}$.
Unfortunately, there does not seem to be a single approach that works for everything.
Example 5.4.
  1. (i)
    One might be tempted to define $\tilde{f}=f_{1}\otimes f_{2}$ by $\tilde{f}(v\otimes w)=f_{1}(v)\otimes f_{2}(w)$ and $f_{1}\otimes f_{2}$ to be zero on non-simple tensors. However, this would lead to the situation where $\tilde{f}(\lambda_{1}\otimes\lambda_{2}+\lambda_{2}\otimes\lambda_{1})=\tilde{f}(\lambda_{1}\otimes\lambda_{1}+\lambda_{2}\otimes\lambda_{2})=\mathbb{0}$ whereas $\tilde{f}((\lambda_{1}+\lambda_{2})\otimes(\lambda_{1}+\lambda_{2}))=f(\lambda_{1}+\lambda_{2})\otimes g(\lambda_{1}+\lambda_{2}).$
  2. (ii)
    Even worse, the decomposition of a tensor into the sum of two simple tensors need not be unique. For example
    $\lambda_{1}\otimes\lambda_{2}+(\lambda_{1}+\lambda_{2})\otimes\lambda_{3}=\lambda_{1}\otimes(\lambda_{2}+\lambda_{3})+\lambda_{2}\otimes\lambda_{3}.$
Here is the piece that we can salvage in general.
Proposition 5.5.
Suppose $\mathcal{M}_{1},\mathcal{N}_{1}$ are $\mathcal{T}$-modules and $\mathcal{M}_{2},\mathcal{N}_{2}$ are left $\mathcal{T}$-modules. If $f_{1}:\mathcal{M}_{1}\to\mathcal{N}_{1}$ and $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$ are multiplicative maps, then there is a 1-balanced map, which we denote as $f_{1}\times_{\circ}f_{2}:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes\mathcal{N}_{2}$, given by $(b_{1},b_{2})\mapsto f_{1}(b_{1})\otimes f_{2}(b_{2}).$
Proof.
$(f_{1}\times_{\circ}f_{2})(b_{1}a,b_{2})=f_{1}(b_{1})a\otimes f_{2}(b_{2})=f_{1}(b_{1})\otimes af_{2}(b_{2})=(f_{1}\times_{\circ}f_{2})(b_{1},ab_{2}).$
Remark 5.6.
If $\mathcal{M}_{1},\mathcal{N}_{1}$ are $(\mathcal{T}_{1},\mathcal{T})$-bimodules and $\mathcal{M}_{2},\mathcal{N}_{2}$ are $(\mathcal{T},\mathcal{T}_{2})$-bimodules, then $\mathcal{M}_{1}\times\mathcal{M}_{2}$ is a $\mathcal{T}_{1},\mathcal{T}_{2}$-bimodule, where the $\mathcal{T}_{1}$ action is on the left on $\mathcal{M}_{1},$ and the $\mathcal{T}_{2}$ action is on the left on $\mathcal{M}_{2}.$ Then $f_{1}\times_{\circ}f_{2}$ of Proposition 5.5 preserves these left and right actions.
Given a $(\mathcal{T}_{1},\mathcal{T})$-bimodule prepair $({\mathcal{M}_{1}},{\mathcal{M}_{1}}_{0})$, a $\mathcal{T},\mathcal{T}_{2}$-bimodule prepair $({\mathcal{M}_{2}},{\mathcal{M}_{2}}_{0})$, and a $\mathcal{T}_{1},\mathcal{T}_{2}$-bimodule prepair $(\mathcal{N},\mathcal{N}_{0})$, define $\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{N})$ to be the set of $\mathcal{T}_{1},\mathcal{T}_{2}$ multiplicative maps $f:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}$ (as in Remark 5.6) satisfying the property that if $b_{j},b_{i,j}\in\mathcal{M}_{i}$ with $\sum_{j}b_{i,j}\in{\mathcal{M}_{i}}_{0},$ then
$f(\sum_{j}b_{1,j},b_{2}),f(b_{1},\sum_{j}b_{2,j})\in\mathcal{N}_{0}.$
Lemma 5.7.
If $f_{i}:(\mathcal{M}_{i},{\mathcal{M}_{i}}_{0})\to(\mathcal{N}_{i},{\mathcal{N}_{i}}_{0})$ are weak morphisms, then
$f_{1}\times_{\circ}f_{2}\in\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{N}_{1}\otimes\mathcal{N}_{2}).$
Proof.
$(f_{1}\times_{\circ}f_{2})(\sum_{j}b_{1,j},b_{2})=f_{1}(\sum_{j}b_{1,j})\otimes f_{2}(b_{2})\in{(\mathcal{N}_{1}\otimes\mathcal{N}_{2}})_{0}.$ Likewise for the other side. ∎
Nevertheless, note that $\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{N}_{1}\otimes\mathcal{N}_{2})$ need not induce weak morphisms in the tensor product, since there may be some weird sum in $(\mathcal{M}_{1}\times\mathcal{M}_{2})_{0}$.
5.2. The tensor product of morphisms when $\mathcal{M}_{2}$ is free
Proposition 5.5 raises the issue of how to define $(f\otimes g)(v\otimes w)$, hoping to extend it to all of $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$.
5.2.1. The tensor product of a weak morphism and a homomorphism when $\mathcal{M}_{2}$ is free
When $\mathcal{M}_{2}$ is free as a $\mathcal{T}$-module, we can still define the tensor product of a weak morphism $f_{1}$ and a homomorphism $f_{2}$. Namely, we write tensors uniquely in terms of a base, by taking an arbitrary element $\sum_{i\in I}v_{i}\otimes b_{i}$ as in Corollary 4.16(v), and put $(f_{1}\otimes f_{2})(\sum_{i\in I}v_{i}\otimes b_{i})=f_{1}(v_{i})\otimes f_{2}(b_{i})$. We formulate this as:
Theorem 5.8.
Notation as in Theorem 4.12, suppose $\mathcal{M}_{2}$ is free, with base $\{b_{i}:i\in I\}$. Then for any weak $(\mathcal{T}_{1},\mathcal{T})$-bimodule morphism (resp.  $\preceq$-morphism) $f_{1}:\mathcal{M}_{1}\to\mathcal{N}_{1}$ and any $(\mathcal{T},\mathcal{T}_{2})$-bimodule homomorphism $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$, there is a weak $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule morphism (resp.  $\preceq$-morphism)
$f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{N}_{1}\otimes_{\mathcal{T}}\mathcal{N}_{2}$
given by $(f_{1}\otimes f_{2})(\sum_{j}v_{j}\otimes b_{j})=\sum_{j}f_{1}(v_{j})\otimes f_{2}(b_{j}).$
Proof.
We do it for $\preceq$-morphisms. The proof for weak morphisms is parallel. Define the map $f_{1}\times_{\circ}f_{2}$ as in Proposition 5.5. Then $f_{1}\otimes f_{2}$ is well-defined when we write vectors of $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ uniquely in the form $\sum v_{j}\otimes b_{j}$, and
$(f_{1}\otimes f_{2})$ $(\sum v_{j}\otimes b_{j}+\sum v_{j}^{\prime}\otimes b_{j})=(f_{1}\otimes f_{2})\left((\sum v_{j}+\sum v_{j}^{\prime})\otimes b_{j}\right)$ (5.5)
$=f_{1}(\sum v_{j}+\sum v_{j}^{\prime})\otimes f_{2}(b_{j})\preceq\sum_{j}f_{1}(v_{j})\otimes f_{2}(b_{j})+\sum_{j}f_{1}(v_{j}^{\prime})\otimes f_{2}(b_{j})$
$=(f_{1}\otimes f_{2})(\sum v_{j}\otimes b_{j})+(f_{1}\otimes f_{2})(\sum v_{j}^{\prime}\otimes b_{j}).$
Corollary 5.9.
We do have functoriality in the residue case under the hypothesis of Lemma 3.12.
Likewise, at times we can extend $\preceq$-morphisms and weak morphisms to tensor extensions.
Theorem 5.10.
Suppose $\mathcal{T}_{1},\mathcal{T}_{2}$ are monoids containing $\mathcal{T}$ where $\mathcal{T}_{1}$ is the disjoint union $\underset{\cdot}{\cup}\,c_{i}\mathcal{T}$, and $\mathcal{M}_{1},\mathcal{M}_{2}$ are $\mathcal{T}$-modules. If $f_{1}:\mathcal{T}_{1}\to\mathcal{T}_{2}$ is a monoid homomorphism and $f_{2}:\mathcal{M}_{1}\to\mathcal{M}_{2}$ is a weak morphism, then there is a weak $\mathcal{T}_{1}$-morphism
$f_{1}\otimes f_{2}:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{1}\to\mathcal{T}_{2}\otimes_{\mathcal{T}}\mathcal{A}_{2}$
given by $(f_{1}\otimes f_{2})(a\otimes b)=f_{1}(a)\otimes f_{2}(b),$ for $a\in\mathcal{T}_{1},\ b\in\mathcal{M}_{1}$
Proof.
$f_{1}\otimes f_{2}$ is well defined because of the hypothesis on $\mathcal{T}_{1},$ so the proof is as in Theorem 4.24(iii). ∎
5.2.2. The tensor product of two weak morphisms or two $\preceq$-morphisms when $\mathcal{M}_{2}$ is free
When dealing with two weak morphisms or $\preceq$-morphisms, we modify the idea of [20], introducing the empty set and partial maps.
Proposition 5.11.
If $\mathcal{M}_{2}$ is free and $f_{1}$ and $f_{2}$ both are weak morphisms or $\preceq$-morphisms, we can define
$(f_{1}\otimes f_{2})(\sum_{i=1}^{m}v_{i}\otimes b_{i})=\begin{cases}f_{1}(v_{1})\otimes f_{2}(b_{1}),\qquad t=1,\\ \emptyset,\qquad t>1.\end{cases}$ (5.6)
Proof.
We do it for $\preceq$. The only instance where $f_{1}\otimes f_{2}$ is defined on a sum is for $v_{1}\otimes b+v_{1}^{\prime}\otimes b=(v_{1}+v_{1}^{\prime})\otimes b,$ where
$(f_{1}\otimes f_{2})((v_{1}+v_{1}^{\prime})\otimes b)=f_{1}(v_{1}+v_{1}^{\prime})\otimes f_{2}(b)\preceq f_{1}(v_{1})\otimes f_{2}(b)+f_{1}(v_{1}^{\prime})\otimes f_{2}(b)=(f_{1}\otimes f_{2})(v_{1}\otimes b)+(f_{1}\otimes f_{2})(v_{1}^{\prime}\otimes b).$
5.3. $\preceq$-Tensor products of $\preceq$-morphisms
Another way to obtain single-valued functions is by introducing an extra condition on the surpassing relation $\preceq$ (which holds for hyperrings, for example). This tensor product might not satisfy the adjoint isomorphism below, but is intriguing.
Remark 5.12.
Recall from [1, Theorem 5.36] that any hypermodule must satisfy the condition that any set of subsets of $\mathcal{P}(\mathcal{H})$ has an infinum, namely the intersection.
Theorem 5.13.
Suppose $(\mathcal{N},\preceq)$ is a semilattice closed under infina of arbitrary sets. Then given maps $f_{i}:\mathcal{M}_{i}\to\mathcal{N}$, define $f:=f_{1}\otimes_{\preceq}f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}$ by
$f(\sum v_{i}\otimes w_{i})=\bigwedge\left\{\sum f_{1}(v_{i}^{\prime})\otimes f_{2}(w_{i}^{\prime}):\sum v_{i}^{\prime}\otimes w_{i}^{\prime}=\sum v_{i}\otimes w_{i}\right\}.$ (5.7)
Then $f$ is a $\preceq$-morphism.
Proof.
We need to show that $f(\sum v_{i,1}\otimes w_{i,1}+\sum v_{i,2}\otimes w_{i,2})\preceq f(\sum v_{i,1}\otimes w_{i,1})+f(\sum v_{i,2}\otimes w_{i,2})$. But if $\sum v_{i,1}^{\prime}\otimes w_{i,1}^{\prime}=\sum v_{i,1}\otimes w_{i,1}$ and $\sum v_{i,2}^{\prime}\otimes w_{i,2}^{\prime}=\sum v_{i,2}\otimes w_{i,2}$ then
$\sum v_{i,1}^{\prime}\otimes w_{i,1}^{\prime}+\sum v_{i,2}^{\prime}\otimes w_{i,2}^{\prime}=\sum v_{i,1}\otimes w_{i,1}+\sum v_{i,2}\otimes w_{i,2},$
so
$f\left(\sum v_{i,1}\otimes w_{i,1}+\sum v_{i,2}\otimes w_{i,2}\right)$ $\preceq\bigwedge\left(\sum f_{1}(v_{i,1}^{\prime})\otimes f_{2}(w_{i,1}^{\prime})\right)+\bigwedge\left(\sum f_{1}(v_{i,2}^{\prime})\otimes f_{2}\left(w_{i,2}^{\prime}\right)\right)$ (5.8)
$=f\left(\sum v_{i,1}\otimes w_{i,1}\right)+f\left(\sum v_{i,2}\otimes w_{i,2}\right),$
as desired. ∎
Theorem 5.13 can be applied to hypermodules in view of Remark 3.4, but at cost of introducing the empty set (the $\infty$ element of $\mathcal{P}(\mathcal{H}),$ cf. footnote 3 of Definition 3.1). In other words, $f$ could be undefined on tensors which can be rewritten in inequivalent ways.
5.3.1. Digression: Hypertensor products of maps
An intriguing (but problematic) way is to introduce a hyperstructure.
Remark 5.14.
Given maps $f_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}$, $i=1,2,$ define $f:=f_{1}\otimes_{\supseteq}f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to{\mathcal{P}}^{\sharp}(\mathcal{N}_{1}\otimes\mathcal{N}_{2})$ by $f(\sum v_{i}\otimes w_{i})=\{\sum f_{1}(v_{i}^{\prime})\otimes f_{2}(w_{i}^{\prime}):\sum v_{i}^{\prime}\otimes w_{i}^{\prime}=\sum v_{i}\otimes w_{i}\}.$
It is easy to see that if $f_{1},f_{2}$ are $\supseteq$-morphisms on hyperpairs then $f_{1}\otimes_{\supseteq}f_{2}$ is a $\supseteq$-morphism. The difficulty with this observation is that for the hyperpairs of greatest interest (for quotient hyperfields) the natural morphisms are $\subseteq$-morphisms.
5.4. Tensor extensions of $\preceq$-morphisms and weak morphisms of pre-semialgebra pairs
One can push this theory further, taking tensor extensions of pre-semialgebras as in Definition 4.20.
Theorem 5.15.
Suppose that $\mathcal{A}$ is a weakly admissible $\mathcal{T}$-semialgebra, $\mathcal{T}^{\prime}$ is a monoid containing $\mathcal{T}$, and $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a weak pre-semialgebra morphism (resp.  $\preceq$-morphism).
  1. (i)
    If $\mathcal{A}_{1}$ is free over $\mathcal{T}$ with base $\{b_{i}:i\in I\},$ then $f$ extends naturally to a weak pre-semialgebra morphism (resp.  $\preceq$-morphism) $\tilde{f}:{\mathcal{T}^{\prime}}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{N}_{2}$ given by $\tilde{f}(\sum_{i}a^{\prime}_{i}\otimes b_{i})=\sum_{i}a^{\prime}_{i}\otimes f(b_{i}).$
  2. (ii)
    If $\mathcal{T}^{\prime}$ is the disjoint union $\underset{\cdot}{\cup}\,c_{i}\mathcal{T}$ for $c_{i}\in\mathcal{T}^{\prime},$ then $f$ extends naturally to a weak pre-semialgebra morphism (resp.  $\preceq$-morphism) $\tilde{f}:{\mathcal{T}^{\prime}}\otimes_{\mathcal{T}}\mathcal{M}_{2}\to\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{N}_{2}$ given by $\tilde{f}(c_{i}a\otimes y)=c_{i}\otimes f(ay).$
Proof.
We do it for $\preceq$-morphisms.
$\text{(i)}\quad\tilde{f}\left(\sum_{i}(a_{i}\otimes b_{i})+\sum_{i}(a_{i}^{\prime}\otimes b_{i})\right)=\tilde{f}\bigg(\sum_{i}(a_{i}+a_{i}^{\prime})$ $b_{i}\bigg)$ (5.9)
$\preceq\sum_{i}a_{i}\otimes f(b_{i})+\sum_{i}a_{i}^{\prime}\otimes f(b_{i})=\tilde{f}(\sum_{i}a_{i}$ $\otimes b_{i})+\tilde{f}(\sum_{i}a_{i}^{\prime}\otimes b_{i}),\quad\forall a_{i},a_{i}^{\prime}\in\mathcal{T}^{\prime}.$
(ii) Using Theorem 4.24, we have
$\qquad\tilde{f}(c_{i}a\otimes y+$ $c_{i}a^{\prime}\otimes y^{\prime})=\tilde{f}(c_{i}(a\otimes y+a^{\prime}\otimes y^{\prime}))=c_{i}\otimes f(ay+a^{\prime}y^{\prime})$ (5.10)
$\preceq c_{i}a\otimes f(y)+c_{i}a^{\prime}\otimes f(y^{\prime})=\tilde{f}(c_{i}a\otimes y)+\tilde{f}(c_{i}a^{\prime}\otimes y^{\prime}).$
Example 5.16.
Any $\preceq$-morphism (resp. weak morphism) $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ of semiring pairs extends to $\preceq$-morphism (resp. weak morphism) $f:(\mathcal{A}[\lambda],\mathcal{A}_{0}[\lambda])\to(\mathcal{A}^{\prime}[\lambda],\mathcal{A}^{\prime}_{0}[\lambda])$ of the polynomial semiring pairs, taking $\mathcal{T}^{\prime}=\cup_{i\in\mathbb{N}}\mathcal{T}\lambda^{i}.$
5.5. The adjoint correspondence
Remark 5.17.
Suppose that $\mathcal{M}_{i}$ are $(\mathcal{T}_{i},\mathcal{T}_{i+1})$-bimodules. The celebrated adjoint isomorphism theorem says
$\operatorname{Hom}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})\approx\operatorname{Hom}(\mathcal{M}_{2},\operatorname{Hom}(\mathcal{M}_{1},\mathcal{M}_{3})).$
The usual proof for modules over rings, say in [22, Proposition 2.10.9] matches values on simple tensors, and then defines the homomorphism $f=f_{1}\otimes f_{2}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{N}$ from its action on simple tensors. This still works for homomorphisms of modules over semirings.
But what about different sorts of morphisms, i.e., $\preceq$-morphisms or weak morphisms? First of all, we have seen for $\preceq$-morphisms that $f\otimes g$ need not be well-defined on simple tensors. Even if it is well-defined, extending it to all tensors is a challenge, which we take on in this section. First we lower our sights in general.
Theorem 5.18.
$\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{M}_{3})\approx\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3})).$
Proof.
As in the proof of [22, Proposition 2.10.9]. Given $f\in\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{M}_{3})$ and $w\in\mathcal{M}_{2},$ we define $f_{w}:\mathcal{M}_{1}\to\mathcal{M}_{3}$ by $v\mapsto f(v\otimes w).$ Then for $\sum v_{i}\in{\mathcal{M}_{1}}_{0}$,
$f_{w}(v_{1}+v_{2})=f\left(\left(\sum v_{i}\right)\otimes w\right)=f\left(\sum v_{i}\otimes w\right)\in{\mathcal{M}_{3}}_{0},$
and likewise for the other side, so $f_{w}\in\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}).$ We get $\Phi\in\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}))$ by declaring $\Phi(f)(w)=f_{w}.$
Conversely, given $g:\mathcal{M}_{2}\to\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}),$ one defines the map $\psi_{g}:(v,w)\mapsto g(w)(v)$.
These correspondences clearly are inverses. ∎
However, the maps thus obtained in $\operatorname{WMor}(\mathcal{M}_{1}\times\mathcal{M}_{2},\mathcal{M}_{3})$ need not be balanced, so might not produce maps of tensor products. In case $\mathcal{M}_{2}$ is free, with base $\mathcal{B}$ over $\mathcal{T}_{2},$ we utilize Proposition 5.11. Define ${\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\mathcal{N})$ to be the multiplicative maps which are 0 except on multiples of base elements.
Theorem 5.19.
${\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3}))$ is a section of ${\operatorname{Mor}_{\preceq}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ when $\mathcal{M}_{2}$ is free over $\mathcal{T}_{2}.$
Proof.
We adapt the proof of Theorem 5.18. Given $f\in{\operatorname{Mor}_{\preceq}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ and $w=ab$ for $b\in{\mathcal{B}},$ we define $f_{w}:\mathcal{M}_{1}\to\mathcal{M}_{3}$ by $v\mapsto f(v\otimes w).$ Then $f_{w}(v+v^{\prime})=f((v+v^{\prime})\otimes w)\preceq f(v\otimes w)+f(v^{\prime}\otimes w).$ We define $\Phi\in{\operatorname{Mor}_{\preceq}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ by $\Phi(f)(w)=f_{w}.$
In the other direction, given $g\in{\operatorname{Mor}_{\preceq}}_{\mathcal{B}}(\mathcal{M}_{2},\operatorname{Mor}_{\operatorname{wk}}(\mathcal{M}_{1},\mathcal{M}_{3})),$ one defines $\psi_{g}:\mathcal{M}_{1}\otimes\mathcal{M}_{2}\to\mathcal{M}_{3}$ by
$\psi_{g}(\sum_{j=1}^{t}v_{j}\otimes b_{j})=\begin{cases}g(b_{1})v_{1},\qquad t=1,\\ 0,\qquad\qquad t>1.\end{cases}$
$\psi_{g}$ is a $\preceq$-morphism since $\psi_{g}((\sum_{j=1}^{t}v_{j}\otimes b_{j}))+(\sum_{j=1}^{t^{\prime}}v^{\prime}_{j}\otimes b^{\prime}_{j})=0$ unless $t=t^{\prime}=1$ and $b_{1}^{\prime}=b_{1},$ in which case $\psi_{g}((v_{1}\otimes b_{1}))+(v^{\prime}_{1}\otimes b_{1}))=\psi_{g}((v_{1}+v^{\prime}_{1})\otimes b_{1})=g(b_{1})(v_{1}+v^{\prime}_{1})\preceq g(b_{1})(v_{1})+g(b_{1})(v_{1^{\prime}})$.
Clearly $\Phi(\psi_{g}(w))(v)=0$ unless $w=ab$ for $a\in\mathcal{T}_{2},$ $b\in{\mathcal{B}},$ in which case
$\Phi(\psi_{g}(w))(v)={\psi_{g}}_{w}(v)=ag(b)(v),$
implying $\Phi\psi_{g}=g$. ∎
The analogous result holds for $\mathcal{M}_{1}$ free. The following general observation specializes to the usual adjoint isomorphism in module theory.
Lemma 5.20.
$\mathcal{M}_{i}$ are $(\mathcal{T}_{i},\mathcal{T}_{i+1})$-bimodules. Define ${\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ to be the $\preceq$-morphisms. Then there are canonical functors
$\Phi:{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})\to{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{2}}(\mathcal{M}_{2},{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1},\mathcal{M}_{3})),$
and
$\Psi:{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{2}}(\mathcal{M}_{2},{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1},\mathcal{M}_{3}))\to{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3}),$
given in the proof.
Proof.
As in the proof of [22, Proposition 2.10.9]. Given $f\in{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{M}_{3})$ and an $w\in\mathcal{M}_{2},$ we define $f_{w}:\mathcal{M}_{1}\to\mathcal{M}_{3}$ by $v\mapsto f(v\otimes w).$ Then
$f_{w}(v_{1}+v_{2})=f((v_{1}+v_{2})\otimes w)=f(v_{1}\otimes w+v_{2}\otimes w)\preceq f(v_{1}\otimes w)+f(v_{2}\otimes w),$
so $f\in{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1},\mathcal{M}_{3}).$ We get $\Phi\in{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{2}}(\mathcal{M}_{2},{\operatorname{Mor}_{\preceq}}_{\mathcal{T}_{1}}(\mathcal{M}_{1},\mathcal{M}_{3}))$ by declaring $\Phi(f)(w)=f_{w}.$
Conversely, given $g:\mathcal{M}_{2}\to{\operatorname{Mor}_{\preceq}}(\mathcal{M}_{1},\mathcal{M}_{3}),$ one can define the map $\psi_{g}:(v,w)\mapsto g(w)(v)$, which yields a right $\preceq$-defined morphism $\bar{\psi}_{g}$ using Theorem 5.13, namely
$\bar{\psi}_{g}(\sum_{i}v_{i}\otimes w_{i})=\bigwedge\left\{\sum g(w^{\prime}_{i})(v^{\prime}_{i}):\sum v_{i}\otimes w_{i}=\sum v^{\prime}_{i}\otimes w^{\prime}_{i}\right\}.$